### v2

# A stochastic reconstruction theorem on rectangular increments with an application to a mixed hyperbolic SPDE

Abstract
We extend the stochastic reconstruction theorem to a setting where the underlying family of distributions satisfies some natural conditions involving rectangular increments. This allows us to prove the well-posedness of a new class of mixed stochastic partial differential equations of hyperbolic type which combines standard Walsh stochastic integration and Young products.
Keywords: Reconstruction theorem, Multiparameter stochastic integration, Hyperbolic SPDEs.
MSC 2020: 60L30, 60H15, 46F10.
1 Introduction
1.1 Stochastic reconstruction and hyperbolic SPDEs
Since its introduction in [Hai14], the reconstruction theorem has not only proved to be a foundational result in the theory of regularity structures but also a powerful result in the theory of distributions, as explained in [CZ20]. Loosely speaking, this theorem identifies optimal conditions on a family of distributions $F=\{F_{\mathbf{x}}\colon\mathbf{x}\in\mathbb{R}^{d}\}$, called germ, under which there exists a unique distribution $\mathcal{R}(F)$ which provides a “good approximation” of $F_{\mathbf{x}}$ around each point $\mathbf{x}\in\mathbb{R}^{d}$. The simplest set of conditions to formulate these properties involves the property of coherence [^1] The usual definition of coherence involves $\gamma\in\mathbb{R}$ but in the context of this article, we focus on the case of unique reconstruction, which requires $\gamma>0$. : for some $\gamma>0$ and $\alpha\leq 0$ one has a growth condition of the form
$\left|(F_{\mathbf{x}}-F_{\mathbf{y}})(\psi_{\mathbf{x}}^{\lambda})\right|\lesssim\lambda^{\alpha}(|\mathbf{x}-\mathbf{y}|+\lambda)^{\gamma-\alpha}$ (1.1)
where $\lambda\in(0,1]$, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, $\psi_{\mathbf{x}}^{\lambda}(\cdot)=\lambda^{-d}\psi(\lambda^{-1}(\cdot-\mathbf{x}))$ is a rescaled and recentered version of a test function $\psi$ and $\lesssim$ denotes inequality up to a constant. Further analytic conditions were studied in [BL23, ZK22].
In this paper, we are interested in a generalization of this result where the underlying family $F$ is composed of random distributions depending on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ which are properly adapted to a multiparameter filtration =$\mathcal{F}=\{\mathcal{F}_{\mathbf{x}}\colon\mathbf{x}\in\mathbb{R}^{d}\}$ (see Definition 2.15). This setting is not novel in the literature. A previous result [Ker23] showed we can modify the conditions (1.1) into
$\begin{split}\left\|(F_{\mathbf{x}}-F_{\mathbf{y}})(\psi_{\mathbf{x}}^{\lambda})\right\|_{m}&\lesssim\lambda^{\alpha}(|\mathbf{x}-\mathbf{y}|+\lambda)^{\gamma-\alpha}\,,\\ \left\|\mathbb{E}^{i}_{\mathbf{x}}(F_{\mathbf{x}}-F_{\mathbf{y}})(\psi_{\mathbf{x}}^{\lambda})\right\|_{m}&\lesssim\lambda^{\alpha}(\left|\mathbf{x}-\mathbf{y}\right|+\lambda)^{\gamma-\alpha+\delta}\quad\text{for any}\;i\leq d\,,\end{split}$ (1.2)
where $\delta>0$, $\|X\|_{m}=\mathbb{E}(|X|^{m})^{1/m}$ and $\mathbb{E}^{i}_{\mathbf{x}}(X)=\mathbb{E}[X|\mathcal{F}^{i}_{\mathbf{x}}]$ is the conditional expectation with respect to $\mathcal{F}^{i}_{\mathbf{x}}$, the filtration generated by the $i$-th component of $\mathbf{x}$. Remarkably, the uniqueness result holds even in the case $20>\gamma>-d/$ 2 provided that $\gamma+\delta>0$. Heuristically, we leverage the gain of regularity obtained by conditioning the germ by allowing a lower index of general regularity $\gamma$. This type of argument is not new in the rough analysis literature but it follows the same idea of the stochastic sewing lemma [Le20], from which [Ker23] was inspired. Moreover, stochastic sewing arises as a special case of the stochastic reconstruction when $d=1$, see [Ker23, Section 5]. From a stochastic analysis perspective, stochastic reconstruction provides a new analytic framework to study Walsh integration [Wal86] with respect to a white noise $\xi$ over +$\mathbb{R}^{d}_{+}:=[0,+\infty)^{d}$ as a Young product, see [CZ20]. Recalling the existence of a measurable modification of $\xi$ in the space of distribution, for any adapted process $Y$ satisfying some weak regularity assumption one can check that the germ $(Y\cdot\xi)_{\mathbf{x}}(\psi):=Y_{\mathbf{x}}\xi(\psi)$ satisfies (1.2) and one obtains
$\int_{[0,+\infty)^{d}}\psi(\mathbf{x})Y_{\mathbf{x}}\xi(d\mathbf{x})=\mathcal{R}(Y\cdot\xi)(\psi)\,,$ (1.3)
where the left-hand side is a standard Walsh integral, see [Ker23, Theorem 53].
However, looking at the several applications coming from stochastic sewing in the recent literature, very few of them were extended in higher dimensions via stochastic reconstruction. One key example to understand this discrepancy is the impossibility of multiplying a multiparameter martingale with a deterministic distribution. Consider e.g. a $d$-dimensional Brownian sheet $B$ and a deterministic distribution $\zeta$ and take the germ $(B\cdot\zeta)_{\mathbf{x}}(\psi):=B_{\mathbf{x}}\zeta(\psi)$. In that case, we see that the quantity
$\mathbb{E}^{i}_{\mathbf{x}}[((B\cdot\zeta)_{\mathbf{y}}-(B\cdot\zeta)_{\mathbf{x}})(\psi_{\mathbf{x}}^{\lambda})]$ (1.4)
has a better regularity (it just vanishes) only when $d=1$. In higher dimensions we do not have any improvement, locking the possibility of using the stochastic reconstruction theorem.
The first major novelty of this paper is the introduction of new conditions on random germs so that it is still possible to reconstruct stochastic integrals, and they also allow a germ like $B\cdot\zeta$ to be reconstructed. Indeed in the simple case $d=2$ we can consider the rectangular increment of the previous germ $B\cdot\zeta$
$\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}(B\cdot\zeta):=(B\cdot\zeta)_{(y_{1},y_{2})}-(B\cdot\zeta)_{(x_{1},y_{2})}-(B\cdot\zeta)_{(y_{1},x_{2})}+(B\cdot\zeta)_{(x_{1},x_{2})}\,.$
A simple computation shows that for $i\in\{1,2\}$
$\mathbb{E}^{i}_{\mathbf{x}}(\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}(B\cdot\zeta)(\psi_{(x_{1},x_{2})}^{\lambda}))=0\,.$
And we recover the regularising effect of conditioning. By simply noticing that this cancellation also holds even when we consider the germ $(Y\cdot\xi)$ a version of stochastic reconstruction on rectangular increment could be a more powerful tool to extend the properties of stochastic sewing in several dimensions. This idea does fit also with the historical multiparameter extension of stochastic calculus, see [Imk85, Kho02, Nor95] where the combinatorics of rectangular increments play an important role.
To state a proper result we will need to apply conditional expectation in different directions using the notation $\mathbb{E}^{\eta}_{\mathbf{x}}:=\prod_{i\in\eta}\mathbb{E}^{i}_{\mathbf{x}}$ for any subset $\eta\subset\{1\,,\ldots\,,d\}$. Note that this is only well-defined if the filtrations $\mathcal{F}^{i}_{\mathbf{x}}$ fulfill a commuting property, which is standard for multiparameter martingale theory, see for example [Kho02, Section 3.4]. Combining the conditioning operators and the rectangular increments of the form $\left|\square^{\theta}_{\mathbf{x},\mathbf{y}}F(\psi)\right|$ for each $\theta\subset\{1,\dots,d\}$, see Section 2, we can formulate in a simplified form the first main result of the paper. For a rigorous statement see Theorem 3.12.
Theorem 1.1.
Let $F$ be a stochastic germ satisfying the rectangular coherence property: for some $m\geq 2$, $\alpha<0$, $\delta>0$ and $\gamma>-\frac{1}{2}$ such that $\gamma+\delta>0$ one has the growth condition
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square^{\theta}_{\mathbf{x},\mathbf{y}}F(\psi_{\mathbf{y}}^{{\bm{\lambda}}})\right\|_{m}\lesssim\prod_{i=1}^{d}\lambda_{i}^{\alpha}\prod_{i\in\theta}(|x_{i}-y_{i}|+\lambda_{i})^{\gamma-\alpha}\prod_{j\in\eta}(|x_{j}-y_{j}|+\lambda_{i})^{\delta}\,,$
where ${\bm{\lambda}}=(\lambda_{1}\,,\ldots\,,\lambda_{d})\in(0,1]^{d}$, $\mathbf{x}=(x_{1}\,,\ldots,x_{d})$ $\mathbf{y}=(y_{1}\,,\ldots,y_{d})$ are in compact set, and $\psi_{\mathbf{y}}^{{\bm{\lambda}}}$ is a rescaled test function like in (2.7). Then there exist unique family of distributions $\mathcal{R}^{\theta}_{\mathbf{x}}(F)$ for each $\theta\subset\{1,\dots,d\}$ , such that $\mathcal{R}_{\mathbf{x}}^{\theta}(F)$ does not depend on $x_{i}$ for $i\in\theta$ and for any non empty subset one has the growth $\eta\subset\theta\neq\emptyset$,
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\sum_{\hat{\theta}\subset\theta}(-1)^{\sharp\hat{\theta}}\mathcal{R}^{\hat{\theta}}_{\mathbf{x}}(F)(\psi^{\bm{\lambda}}_{\mathbf{z}})\right\|_{m}\lesssim\prod_{i\in\theta}\lambda_{i}^{\gamma}\prod_{j\in\eta}\lambda_{j}^{\delta}\prod_{k\in\theta^{c}}\lambda_{k}^{\alpha}\,,$ (1.5)
for any $\mathbf{x},\mathbf{z}$ such that $x_{i}=z_{i}$ for $i\in\theta$ and ${\bm{\lambda}}\in(0,1]^{d}$. We call $\mathcal{R}^{\{1,\ldots,d\}}_{\mathbf{x}}(F)$ the multiparameter reconstruction of the stochastic germ $F$ and we denote it by $\mathcal{R}(F)$.
Remark 1.2.
Unlike the statements in [CZ20, Ker23], in this statement we allow one to freeze certain variables $x_{i},i\notin\theta\subset\{1,\dots,d\}$ and only apply the reconstruction techniques for directions $i\in\theta$. As a result, we do not get a single distribution $\mathcal{R}(F)$, but a family of partial reconstructions $\mathcal{R}^{\theta}_{\mathbf{x}}(F)$. Let us illustrate these on example (1.3) in $d=2$. By the same arguments as [Ker23, Theorem 53],
$\mathcal{R}^{\emptyset}_{\mathbf{x}}(Y\cdot\xi)(\psi)$ $=\int_{[0,+\infty)^{2}}\psi(\mathbf{z})Y_{\mathbf{x}}\xi(d\mathbf{z})$
$\mathcal{R}^{\{1\}}_{\mathbf{x}}(Y\cdot\xi)(\psi)$ $=\int_{[0,+\infty)^{2}}\psi(\mathbf{z})Y_{(z_{1},x_{2})}\xi(d\mathbf{z})$
$\mathcal{R}^{\{2\}}_{\mathbf{x}}(Y\cdot\xi)(\psi)$ $=\int_{[0,+\infty)^{2}}\psi(\mathbf{z})Y_{(x_{1},z_{2})}\xi(d\mathbf{z})$
$\mathcal{R}(Y\cdot\xi)(\psi)=\mathcal{R}^{\{1,2\}}_{\mathbf{x}}(Y\cdot\xi)(\psi)$ $=\int_{[0,+\infty)^{2}}\psi(\mathbf{z})Y_{\mathbf{z}}\xi(d\mathbf{z})\,,$
are the Walsh integrals, where $\theta$ decides over which variables we integrate $Y$. For $i\notin\theta$, $Y$ gets evaluated at the “frozen” point $x_{i}$. This also links (1.5) to rectangular increments, as in the above example
$\sum_{\hat{\theta}\subset\theta}(-1)^{\sharp{\hat{\theta}}}\mathcal{R}^{\hat{\theta}}_{\mathbf{x}}(Y\cdot\xi)(\psi)=\int_{[0,+\infty)^{2}}\psi(\mathbf{z})\square^{\theta}_{\mathbf{x},\mathbf{z}}Y\xi(d\mathbf{z})$
is the integral over the rectangular increments of $Y$ for all $\theta\subset\{1,2\}$.
Remark 1.3.
It should be noted that the proof of the stochastic reconstruction theorem is closer to the original proof of Martin Hairer than the later improvements [CZ20, BL23, ZK22]. This is because Martin Hairer’s proof is based on wavelet approximations, which approximate the target distribution $f$ with the sum
$\mathcal{R}(F)(\psi)\approx\sum_{\mathbf{x}\in\Delta_{\mathbf{n}}}F_{\mathbf{x}}(\phi^{\mathbf{n}}_{\mathbf{y}})\left\langle\phi^{\mathbf{n}}_{\mathbf{y}},\psi\right\rangle\,,$ (1.6)
where $\phi^{\mathbf{n}}_{\mathbf{y}}$ is a Daubechies-wavelet and $\Delta_{\mathbf{n}}$ is an increasingly fine mesh over $\mathbb{R}^{d}$, see Section 2. This particular structure allows to treat the convergence of these sums in the same way as [Le20] by using a proper estimate on sums via a generalized BDG inequality, see Lemma 2.14. Moreover, wavelets also play an important role when we want to integrate the reconstruction to obtain a proper random field (see Proposition 4.7).
Coming back again to dimension one, the possibility of multiplying a Brownian motion with a deterministic distribution via stochastic sewing was one of the inspiring principles to develop in [FHL21] a complete theory of rough stochastic differential equations of the form
$dY_{t}=b_{t}(Y_{t})+\sigma_{t}(Y_{t})dW_{t}+f_{t}(Y_{t})d\mathbf{X}_{t}$ (1.7)
[…]
$\begin{split}\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}u(\mathbf{x})&=\sigma(u(\mathbf{x}))\xi+f(\mathbf{x})u(\mathbf{x})\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}Z(\mathbf{x})\quad\mathbf{x}\in\mathbb{R}_{+}^{d}\,\\ u(\mathbf{x})&=v(\mathbf{x})\qquad\quad\mathbf{x}\in\partial\mathbb{R}_{+}^{d}=\bigcup_{i=1}^{d}\{\mathbf{x}\in\mathbb{R}_{+}^{d}\colon x_{i}=0\}\end{split}$ (1.8)
where $\xi$ is a white noise over $\mathbb{R}_{+}^{d}$ and $Z\colon\mathbb{R}_{+}^{d}\to\mathbb{R}$ is a deterministic Hölder field in $C^{\beta}$, see Section 3.1. Moreover, $\sigma\colon\mathbb{R}\to\mathbb{R}$ is a generic non-linearity and $f\colon\Omega\times\mathbb{R}_{+}^{d}\to\mathbb{R}$ is a given random field. The reconstruction theorem in Theorem 1.1 allows us to provide a general solution theory for (1.8).
Formally speaking, equation (1.8) represents a simplified version of (1.7) in several variables but at the same time it is connected with new and old problems both in PDE and SPDE theory: firstly, when $\sigma=0$ there is no randomness, the equation (1.8) is a special case of a differential equation driven by a Hölder field on hyper-cubes, see [Har21] or a special rotated wave equation when $d=2$, see [QST07]. The choice of this specific non-linearity plays an important role in the theory of signature Kernels, see e.g. [SCF+21, LL24]. Concerning the opposite case when $f=0$ the underlying SPDE is classical and it has been intensively studied in dimension $2$, see e.g. [Cai72, NY89b, NY89a, MSS94]. To our knowledge, equations such as (1.8) do not seem to have been addressed in the literature yet, and from a technical point of view, both classical deterministic techniques and techniques based on standard stochastic integration fail to establish well-posedness of this problem.
To give a rigorous meaning to equation (1.8), we recall the identification of stochastic integration with the reconstruction in (1.3) and we will describe the products in the right-hand side as reconstruction of germs of the simple form recalled before. Then we can write the equation into its mild solution form after applying an integration map
$f\to\int_{{\bm{0}}}^{\mathbf{x}}f(d\mathbf{z})\,,\quad\int_{{\bm{0}}}^{\mathbf{x}}=\int_{0}^{x_{1}}\ldots\int_{0}^{x_{d}}$
taking the primitive of a distribution in $d$ different variables, and we get to the equation
$u(\mathbf{x})=\mathcal{I}(v)(\mathbf{x})+\int_{{\bm{0}}}^{\mathbf{x}}\sigma(u(\mathbf{y}))\xi(d\mathbf{y})+\int_{\bm{0}}^{\mathbf{x}}f(\mathbf{y})u(\mathbf{y})\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}Z(d\mathbf{y})\,,$ (1.9)
for some explicit boundary term $\mathcal{I}(v)(\mathbf{x})$, see equation (4.1). Our second main result will be the identification of a class of random field $u$, the space $C^{\alpha,\delta}L_{m}$, see Definition 3.1 where the equation (1.9) is globally well-posed in the so-called Young regime, i.e. when the trajectories of $Z$ are slightly more regular than the trajectories of Brownian sheet. See Theorem 4.14 for a rigorous statement.
Theorem 1.4.
Let $m$ be in $[2,+\infty)$ and $Z\in C^{\beta}$ with $\beta>1/2$. Let $\sigma$ be a Lipschitz function and consider the parameters $\delta$ and $\alpha$ such that $\delta=\beta-1/2$ and $\delta\leq\alpha<1/2$, $\alpha+\beta+\delta>1$. Then for every $v\colon\partial\mathbb{R}_{+}^{d}\to\mathbb{R}$ such that $\mathcal{I}(v)\in C^{\alpha,\delta}L_{m}$ and any $f\in C^{\alpha,\delta}L_{\infty}$ there exists a unique solution $u\in C^{\alpha,\delta}L_{m}$ to (1.9) starting from $v$ over any finite time interval.
Recently, a new stochastic extension of the multiparameter sewing lemma was proposed in [BHK23] to study the effect of regularisation by noise phenomena on a similar SPDE. While some techniques developed in that article are fundamental in the proof of the reconstruction theorem, we think that direct use of that result to solve (1.8), even if possible, would require a higher complexity in the study of the spaces $C^{\alpha,\delta}L_{m}$. In particular, with stochastic reconstruction, one has the advantage of not dealing with multiparameter delta operators, which become cumbersome in high dimensions. In our context, we treated this complexity by developing a specific integration theory in Section 4. Furthermore, a formulation of this SPDE using a reconstruction theorem could be more helpful in extending techniques inspired by regularity structures to more general hyperbolic SPDEs. Further details between our version of stochastic reconstruction and the stochastic multiparameter sewing lemma are studied in Section 5.
1.2 Outline of the paper
We outline the paper by summarising the content of its sections. In Section 2, we start by introducing in full detail the main tools to prove the reconstruction theorem on rectangular increments. We recall the main properties of rectangular increments of maps as explained in [Har21] and extend them to germs. We then recall the main properties of wavelets [Dau88] that we will adapt to define the approximating sequence in (1.6). We also recall some basic results on multiparameter filtrations and the main stochastic estimate in Lemma (2.12).
In Section 3, we then pass to properly define the main spaces of random distributions, random fields, and random germs to properly state Theorem 1.1 and prove it. All definitions will also allow different Hölder parameter and regularisation regularities ${\bm{\alpha}}=(\alpha_{1}\,,\ldots\alpha_{d})$, ${\bm{\delta}}=(\delta_{1}\,,\ldots\delta_{d})$ to keep track of the regularisation effect on any components.
In Section 4, we finally introduce the main operations to define the SPDE (1.9) as a fixed point of the space $C^{\alpha,\delta}L_{m}$. We will consider in particular the composition (see Proposition 4.1), the pointwise product (see Proposition 4.2) and the pointwise product between a random field and a distribution (see Propositions 4.3 and 4.4) which allows us to define all the non-linearities at the level of random distributions. We then combine this operation with a usual integration map, which extends a deterministic result obtained in [Bra19] in several dimensions. Theorem 4.14 then checks that all these operations provide a proper fixed point over the spaces introduced in the previous section.
In Section 5, we show that the multiparameter stochastic sewing lemma from [BHK23] is a special case of our reconstruction theorem 3.12. This is shown in Proposition 5.4. It is based on the same construction, which connects sewing and reconstruction (see e.g. [BZ22]), as well as stochastic sewing and stochastic reconstruction (see [Ker23]), in dimension $d=1$.
1.3 Open extensions arising from our work
We conclude the introduction by outlining two natural extensions raised by the results we established.
In our current formulation, we can only solve (1.9) for non-linearities of the form $f(\mathbf{x})u(\mathbf{x})$. One natural extension of the problem would be to consider a generic non-linearity $g(u)$ in (1.8). However, it seems that the structure of the spaces $C^{\alpha,\delta}L_{m}$ is not well suited to take generic compositions $g(u)$ for $u\in C^{\alpha,\delta}L_{m}$ into account. That is, even for smooth functions $g$, $g(u)$ might not belong to the same space as $u$. This stands in stark contrast to the deterministic case: In $d=2$, it is well-known that if $g\in C^{2}$ then $g(u)\in C^{\bm{\alpha}}$ for any $u\in C^{\bm{\alpha}}$ and ${\bm{\alpha}}=(\alpha_{1},\alpha_{2})\in(0,1)^{2}$ (see e.g. [QST07, Lemma 3.1]). Moreover, for any Hölder function $\sigma\in C^{\eta}$ for some $\eta\in(1,2)$, we still have that $\sigma(u)\in C^{\tilde{\alpha}}$ for a some $\tilde{\alpha}<\alpha$ (see [BHR22, Lemma 15]). Both of these results rely on the application of the classical Taylor formula to the rectangular increment
$\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}g(u)=g(u(y_{1},y_{2}))-g(u(x_{1},y_{2}))-g(u(y_{1},x_{2}))+g(u(x_{1},x_{2}))\,,$
which behaves like the product $\square^{\{1\}}_{\mathbf{x},\mathbf{y}}u\square^{\{2\}}_{\mathbf{x},\mathbf{y}}u$ under some knowledge on the second derivative of $g$. However, in the stochastic case, we need to control $\|\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}g(u)\|_{m}$, where the same argument provides a control of this quantity in terms of $\|\square^{\{1\}}_{\mathbf{x},\mathbf{y}}u\square^{\{2\}}_{\mathbf{x},\mathbf{y}}u\|_{m}$, which might explode since $\square^{\{1\}}_{\mathbf{x},\mathbf{y}}u\square^{\{2\}}_{\mathbf{x},\mathbf{y}}u$ only has finite moments up to $m/2$.
In [FHL21], the authors encountered a similar problem when $d=1$ which they managed to solve with the use of $L_{m,n}$ norms. Using our notations, this would be equivalent to controlling the quantities
$\left\|\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}g(u)\right\|_{m,n}^{\eta}:=\left\|(\mathbb{E}_{\mathbf{x}}^{\eta}|\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}g(u)|^{m})^{\frac{1}{m}}\right\|_{n}\,\quad\emptyset\neq\eta\subset\{1,2\}\,.$
The presence of a conditional expectation inside the norm with an extra parameter $n$ were two fundamental tools to obtain a well-posed composition map in [FHL21, Lemma 3.13]. Here, the lack of integrability is absorbed in the parameter $n$ (which becomes smaller). We expect that a similar result holds even in the rectangular setting modulo a proper extension of the space. Unfortunately, this space is not suited to establish a proper reconstruction theorem in dimension $d>1$. While one might generalize Theorem 3.12 to an $L_{m,n}$ germ $F$, the resulting estimate (1.5) coming from rectangular increments does not give us a direct estimate on $\mathcal{R}(F)(\psi_{\mathbf{x}}^{\lambda})$, but the best we can hope for is to have bounds on
$\left\|\mathbb{E}_{\mathbf{x}}^{\eta}\sum_{\hat{\theta}\subset\theta}(-1)^{\sharp{\hat{\theta}}}\mathcal{R}^{\hat{\theta}}_{\mathbf{x}}(F)(\psi^{\lambda}_{\mathbf{z}})\right\|_{m,n}^{\theta\setminus\eta}\,.$
In dimension one, this estimate together with the assumption that $\left\|F_{\mathbf{x}}\right\|_{n}$ is bounded uniformly suffices to ensure that
$\left\|\mathcal{R}(F)(\psi_{\mathbf{x}}^{\lambda})\right\|_{m,n}^{\{1\}}<\infty\,.$
However, already in dimension two the same assumption on $F_{\mathbf{x}}$ only gives that
$\left\|\mathcal{R}^{\{1\}}_{\mathbf{x}}(F)(\psi_{\mathbf{x}}^{\lambda})\right\|_{m,n}^{\{1\}},\quad\left\|\mathcal{R}^{\{2\}}_{\mathbf{x}}(F)(\psi_{\mathbf{x}}^{\lambda})\right\|_{m,n}^{\{2\}}<\infty\,,$
but makes no statement about
$\left\|\mathcal{R}^{\{2\}}_{\mathbf{x}}(F)(\psi_{\mathbf{x}}^{\lambda})\right\|_{m,n}^{\{1\}},\quad\left\|\mathcal{R}^{\{1\}}_{\mathbf{x}}\,(F)(\psi_{\mathbf{x}}^{\lambda})\right\|_{m,n}^{\{2\}}\,.$
Therefore we cannot conclude that $\left\|\mathcal{R}(\psi_{\mathbf{x}}^{\lambda})\right\|_{m,n}^{\{i\}}$ is finite for $i=1$ or $i=2$. An extension to understand this problem could unlock the possibility of solving a more general class of SPDEs than (1.8).
Another natural extension arising in this work would be to consider a deterministic perturbation $Z$ in (1.8) with lower Hölder regularity $\beta\leq 1/2$ exactly as studied in [FHL21]. While it seems like a naive question, a possible solution to this question would naturally involve a rough path theory of differential equations driven by a Hölder field on hyper-cubes. A consistent solution theory for such an equation has been an open problem in the theory of rough paths since the introduction of rough sheets [CG14, TC15] and it does seem out of reach with the actual technology. However, we think that the presence of this article may arouse new interest in the issue and become a further motivation to analyze this problem again.
Acknowledgements
CB is supported by the ERC starting grant “LoRDeT”; he also acknowledges the support of TU Berlin where he was employed when this project was started. HK is supported by the Research Unit FOR2402 in TU Berlin and the International Research Training Group 2544 of the Berlin Mathematical School. Both authors would like to thank the organizers of the conference “Structural Aspects of Signatures and Rough Paths” held at the Centre for Advanced Study of Oslo, where a preliminary version of these results was presented.
Notation
Throughout the paper, we will adopt the convention of writing multi-indices in $\mathbf{n}\in\mathbb{N}^{d}$ and points in $\mathbf{x}\in\mathbb{R}^{d}$ with bold characters, with the usual notations $\mathbf{x}=(x_{1}\,,\ldots\,,x_{d})$ and $\mathbf{n}=(n_{1}\,,\ldots\,,n_{d})$ and the convention =$\mathbb{N}=\{0,1,2,\ldots\}$. We also use the shorthand notation $[d]=\{1\,,\ldots\,,d\}$ for some integer $d\geq 1$, $\mathbf{0}$ for the zero vector and $\mathbf{1}$ if all the components are equal to $1$. We write $\mathbf{x}\leq\mathbf{y}$ for $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$ if $x_{i}\leq y_{i}$ for all $i\in[d]$. We extend <,><,>< , > and $\geq$ analogously. Generic subsets $\theta\subset[d]$ will also play an important role at the level of multi-indices. We denote by $\mathbf{1}_{\theta}$ the multi-index whose components are equal to $1$ according to the set $\theta$. By definition $\mathbf{0}=\mathbf{1}_{\emptyset}$, $\mathbf{1}=\mathbf{1}_{[d]}$ and in these two special cases we will adopt the notation of the left-hand side. More generally, for any $\mathbf{x}\in\mathbb{R}^{d}$ we will denote by $\mathbf{x}_{\theta}$ the vector whose components are equal to $\mathbf{x}$ according to $\theta$ and they are 00 otherwise. Furthermore, we set $\sharp\theta$ to be the cardinality of $\theta\subset[d]$.
Most classical operations on real numbers extend trivially to the components of a vector. For any non-empty indexset $\theta\subset[d]$ , we set in particular
$(\mathbf{y}^{\mathbf{x}})_{\theta}=\prod_{i\in\theta}y_{i}^{x_{i}}\,,$
$\max(\mathbf{x},\mathbf{y})=\mathbf{x}\vee\mathbf{y}=(x_{1}\vee y_{1},\dots,x_{d}\vee y_{d})\,,\quad\min(\mathbf{x},\mathbf{y})=\mathbf{x}\wedge\mathbf{y}=(x_{1}\wedge y_{1},\dots,x_{d}\wedge y_{d})\,,$
$\mathbf{x}*\mathbf{y}=(x_{1}y_{1},\dots,x_{d}y_{d})\,,$
for $\mathbf{y},\mathbf{x}\in{\mathbb{R}}^{d}$ whenever each component in the product is well-defined. Note that by this notation, $(\mathbf{y}^{\mathbf{x}})_{\theta}=\mathbf{y}^{\mathbf{x}_{\theta}}$. When we have $\theta=[d]$ we will omit the set in the operation. This expression also extends to the case $\theta=\emptyset$ by imposing $(\mathbf{y}^{\mathbf{x}})_{\emptyset}=1$. Similarly, we even allow mixed expressions among scalars and vectors like $\alpha^{\mathbf{y}}=\alpha^{y_{1}}\cdots\alpha^{y_{d}}$ without any ambiguities. The absolute value of a vector in our setting is also a vector which is given by $\left|\mathbf{x}\right|_{\text{c}}=(\left|x_{1}\right|\,,\ldots,\left|x_{n}\right|)$. This notation extends to any subset $\theta\subset[d]$ the product
$\left|\mathbf{x}-\mathbf{y}\right|^{{\bm{\alpha}}}_{\text{c},\theta}=\prod_{i\in\theta}\left|x_{i}-y_{i}\right|^{\alpha_{i}}\,.$
in a synthetic way.
Partial derivatives of a differentiable function $\psi\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$ are denoted by
$\partial^{\mathbf{n}}\psi=(\partial^{n_{1}}_{x_{1}}\ldots\partial^{n_{d}}_{x_{d}})\psi\,.$
We denote by $C^{r}_{c}$, $r\in\mathbb{N}$ the space of $r$-times continuously differentiable functions $\psi\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$ with compact support. We will refer to them as test functions. We equip $C^{r}_{c}$ with the norm
$\left\|\psi\right\|_{C^{r}_{c}}:=\sum_{\left|\mathbf{n}\right|\leq r}\left\|\partial^{\mathbf{n}}\phi\right\|_{\infty}\,,$
where $\left|\mathbf{n}\right|=n_{1}+\cdots+n_{d}$ and $\left\|\cdot\right\|_{\infty}$ is the sup norm. For a compact set $K\subset\mathbb{R}^{d}$, we will denote by $C^{r}(K)$ the subset of $C^{r}_{c}$ containing functions $\psi$ whose support $\mbox{supp}(\psi)$ is contained in $K$ and by $B^{r}(K)$ the subset of functions of $C^{r}(K)$ such that $\left\|\psi\right\|_{C^{r}}=1$.
Throughout the text, we will use the standard notation $x\lesssim y$ to denote the existence of a constant $C>0$, independent of the variables in $x$ and $y$, such that $x\leq Cy$.
2 Preliminaries
We recall now the main analytic and probabilistic tools to formulate the stochastic reconstruction theorem in the context of rectangular increments.
2.1 Rectangular increments of germs
We begin by recalling the rectangular increments of a generic function $f\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$ and of a generic germ of distributions. This part follows [BHK23, Har21] with some slight differences in the notation. For more classical introductions to the topic, especially concerning multiparameter martingale theory, we recommend [Kho02, CW75].
For any given $i\in[d]$ and $\mathbf{x},\mathbf{y}\in{\mathbb{R}}^{d}$ we set the projection $\pi^{i}_{\mathbf{x}}\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ on the $i$-th coordinate of $\mathbf{x}$ by
$\pi^{i}_{\mathbf{x}}\mathbf{y}:=(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{d})\,.$
The operators $\pi^{i}_{\mathbf{x}}$ commute among themselves, allowing us to extend these projections on any subset $\theta\subset[d]$ and $\mathbf{x}\in{\mathbb{R}}^{d}$ as follows
$\pi^{\theta}_{\mathbf{x}}:=\prod_{i\in\theta}\pi^{i}_{\mathbf{x}}\,,$ (2.1)
where we adopt the shorthand notation $\pi^{i}_{\mathbf{x}}\pi^{j}_{\mathbf{x}}=\pi^{i}_{\mathbf{x}}\circ\pi^{j}_{\mathbf{x}}$ and the convention $\pi^{\emptyset}_{\mathbf{x}}=\text{Id}$ the identity function on ${\mathbb{R}}^{d}$. We apply these projections to any function $f\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$ by setting $\pi^{\theta}_{\mathbf{x}}f(\mathbf{y}):=f(\pi^{\theta}_{\mathbf{x}}\mathbf{y})$.
For any given $\mathbf{x},\mathbf{y}\in{\mathbb{R}}^{d}$ and subset $\theta\subset[d]$, the rectangular increment of a function $f\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$ is defined by the expression
$\square_{\mathbf{x},\mathbf{y}}^{\theta}f:=$ $\prod_{i\in\theta}(\text{Id}-\pi^{i}_{\mathbf{x}})f(\pi^{\theta^{c}}_{\mathbf{x}}\mathbf{y})=\sum_{\theta^{\prime}\subset\theta}(-1)^{\sharp(\theta\setminus\theta^{\prime})}f(\pi^{\theta^{\prime}}_{\mathbf{y}}\mathbf{x})=\prod_{i\in\theta}(\pi^{i}_{\mathbf{y}}-\mbox{Id})f(\mathbf{x})\,,$ (2.2)
where we use the same convention of writing composition of operators as product and we set $\theta^{c}=[d]\setminus\theta$. The second expression in (2.2) is a sum over all subset of $\theta^{\prime}\subset\theta$. For $\theta=\emptyset$ we have $\square_{\mathbf{x},\mathbf{y}}^{\emptyset}f=f(\mathbf{x})$. The equality of these two expressions above follows from straight-forward combinatorial considerations, see [Har21, Definition 1].
For instance, when $d=2$ we have the expressions
$\square_{(x_{1},x_{2}),(y_{1},y_{2})}^{\{1\}}f=f(y_{1},x_{2})-f(x_{1},x_{2})\,,\quad\square_{(x_{1},x_{2}),(y_{1},y_{2})}^{\{2\}}f=f(x_{1},y_{2})-f(x_{1},x_{2})\,,$
$\square_{(x_{1},x_{2}),(y_{1},y_{2})}^{\{1,2\}}f=f(y_{1},y_{2})-f(y_{1},x_{2})-f(x_{1},y_{2})+f(x_{1},x_{2})\,.$
Using this expression, we can synthetically show some lengthy identities.
Lemma 2.1.
Let $\mathbf{x},\mathbf{y}\in{\mathbb{R}}^{d}$ and $f,g\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$. We then have the following properties.
  1. 1.
    For any couple of disjoint sets $\theta_{1},\theta_{2}\subset[d]$,
    $\square_{\mathbf{x},\mathbf{y}}^{\theta_{1}}(\square_{\cdot,\mathbf{y}}^{\theta_{2}}f)=\square^{\theta_{1}\cup\theta_{2}}_{\mathbf{x},\mathbf{y}}f\,.$ (2.3)
  2. 2.
    For every $\theta\subset[d]$, we have
    $\square_{\mathbf{x},\mathbf{y}}^{\theta}(f\cdot g)=\sum_{\theta=\theta_{1}\cup\theta_{2}}\square^{\theta_{1}}_{\mathbf{x},\mathbf{y}}f\cdot\square_{\mathbf{x},\mathbf{y}}^{\theta_{2}}g\,,$ (2.4)
    where the sum runs over all $\theta_{1},\theta_{2}\subset[d]$ (not necessarily disjoint!), such that $\theta=\theta_{1}\cup\theta_{2}$.
  3. 3.
    For any couple of subsets $\eta\subset\theta\subset[d]$
    $\square_{\mathbf{x},\mathbf{y}}^{\theta}f=\sum_{\theta^{\prime}\subset\theta\setminus\eta}(-1)^{\#(\theta\setminus(\eta\cup\theta^{\prime}))}\square^{\eta}_{\pi^{\theta^{\prime}}_{\mathbf{y}}\mathbf{x},\mathbf{y}}f\,.$ (2.5)
Proof.
Composing the two definitions of rectangular increments we have
$\square_{\mathbf{x},\mathbf{y}}^{\theta_{1}}(\square_{\cdot,\mathbf{y}}^{\theta_{2}}f)=\prod_{i\in\theta_{1}}(\text{Id}-\pi^{i}_{\mathbf{x}})\prod_{i\in\theta_{2}}(\text{Id}-\pi^{i}_{\pi^{\theta_{1}^{c}}_{\mathbf{x}}\mathbf{y}})f(\pi^{\theta_{2}^{c}}_{\pi^{\theta_{1}^{c}}_{\mathbf{x}}\mathbf{y}}\mathbf{y})\,.$
From the identities
$\pi^{\theta_{2}^{c}}_{\pi^{\theta_{1}^{c}}_{\mathbf{x}}\mathbf{y}}\mathbf{y}=\pi^{\theta_{2}^{c}\cap\theta_{1}^{c}}_{\mathbf{x}}\mathbf{y}\,,\quad\pi^{i}_{\pi^{\theta_{1}^{c}}_{\mathbf{x}}\mathbf{y}}=\pi^{i}_{\mathbf{x}}\,\quad\text{for any $i\in\theta_{2}$}$
we deduce the desired equality. For the second part, note that for each $i\in[d]$, direct computation shows
$\square^{\{i\}}_{\mathbf{x},\mathbf{y}}(f\cdot g)=(\square^{\{i\}}_{\mathbf{x},\mathbf{y}}f)g(\mathbf{x})+f(\mathbf{x})(\square^{\{i\}}_{\mathbf{x},\mathbf{y}}g)+(\square^{\{i\}}_{\mathbf{x},\mathbf{y}}f)(\square^{\{i\}}_{\mathbf{x},\mathbf{y}}g)\,.$
Iterating this with the use of (2.3) leads to (2.4).
The third part follows automatically from (2.3) by choosing $\theta_{2}=\eta$, $\theta_{1}=\theta\setminus\eta$ and writing out the sum definition of $\square_{\mathbf{x},\mathbf{y}}^{\theta\setminus\eta}$. ∎
The definition of rectangular increments (2.2) ensures that $f(\mathbf{x})$ is always contained in $\square^{\theta}_{\mathbf{x},\mathbf{y}}f$, but $f(\mathbf{y})$ might not be. In the context of the coherence property (see Definition 3.8), it is more natural to work with rectangular increments containing $f(\mathbf{y})$. The following lemma allows us to do so.
Lemma 2.2.
Fix $\mathbf{x},\mathbf{y}\in{\mathbb{R}}^{d}$ and a function $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$. For any couple of disjoint sets $\theta,\eta\subset[d]$ one has
$\square_{\mathbf{x},\mathbf{y}}^{\theta}f=\sum_{\eta=\eta_{1}\sqcup\eta_{2}}(-1)^{\sharp{\eta_{2}}}\square^{\theta\cup\eta_{2}}_{\pi^{\eta_{1}}_{\mathbf{y}}\mathbf{x},\mathbf{y}}f\,,$ (2.6)
where we sum over all possible ways to represent $\eta$ as the disjoint union of $\eta_{1}$ and $\eta_{2}$.
Proof.
We think of $\square^{\theta}_{\mathbf{x},\mathbf{y}}f$ as a function of $\mathbf{x}$ for fixed $\mathbf{y}\in\mathbb{R}^{d}$. Then direct computation shows
$\square_{\mathbf{x},\mathbf{y}}^{\theta}f$ $=\prod_{j\in\eta}(\pi_{\mathbf{y}}^{j}-(\pi_{\mathbf{y}}^{j}-\text{Id}))\square_{\mathbf{x},\mathbf{y}}^{\theta}f$
$=\sum_{\eta=\eta_{1}\sqcup\eta_{2}}(-1)^{\sharp{\eta_{2}}}\pi_{\mathbf{y}}^{\eta_{1}}\prod_{i\in\eta_{2}}(\pi^{i}_{\mathbf{y}}-\text{Id})\square_{\mathbf{x},\mathbf{y}}^{\theta}f$
$=\sum_{\eta=\eta_{1}\sqcup\eta_{2}}(-1)^{\sharp{\eta_{2}}}\square^{\theta\cup\eta_{2}}_{\pi^{\eta_{1}}_{\mathbf{y}}\mathbf{x},\mathbf{y}}f\,.$
Example 2.3.
If we work in dimension $d=2$, Lemma 2.2 simplifies to the following three identities applied to the sets $\theta=\emptyset,\{1\}$ and $\eta=\{1,2\}$, $\{2\}$.
$f(\mathbf{x})$ $=\square_{\mathbf{x},\mathbf{y}}^{\emptyset}f=\pi^{1}_{\mathbf{y}}f(\mathbf{x})-\square^{\{1\}}_{\mathbf{x},\mathbf{y}}f\,,$
$f(\mathbf{x})$ $=\square_{\mathbf{x},\mathbf{y}}^{\emptyset}f=f(\mathbf{y})-\square^{\{1\}}_{\pi^{2}_{\mathbf{y}}\mathbf{x},\mathbf{y}}f-\square^{\{2\}}_{\pi^{1}_{\mathbf{y}}\mathbf{x},\mathbf{y}}f+\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}f\,,$
$\square^{\{1\}}_{\mathbf{x},\mathbf{y}}f$ $=\square^{\{1\}}_{\pi_{\mathbf{y}}^{2}\mathbf{x},\mathbf{y}}f-\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}f\,.$
It especially follows, that all the rectangular increments on the right-hand side contain $f(\mathbf{y})$, as long as $\theta\cup\eta=[d]$. This observation extends to the general case $d\geq 2$.
The notion of rectangular increment does not only apply to real-valued functions $f\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$ but it can be easily extended to $f$ taking values in a generic vector space. In our case, the desired vector space will be a proper set of distributions on ${\mathbb{R}}^{d}$ and the resulting functions will take the name of germs. Let us first introduce our space of test functions:
Definition 2.4.
A linear map $T:C_{c}^{r}\to\mathbb{R}$ is called a local distribution of order $r\in\mathbb{N}$ if for all compact sets $K\subset{\mathbb{R}}^{d}$ there exists a constant $C(K)>0$ such that
$\left|T(\psi)\right|\leq C(K)\left\|\psi\right\|_{C^{r}_{c}}\,,$
for all $\psi\in C^{r}_{c}(K)$. We denote by $\mathcal{D}_{r}^{\prime}$ the set of local distributions of order $r\in\mathbb{N}$. For some index $r$ we call any map $F\colon{\mathbb{R}}^{d}\to\mathcal{D}_{r}^{\prime}$ such that for all test functions $\psi$ the function $\mathbf{x}\to F_{\mathbf{x}}(\psi)$ is a measurable germ.
Remark 2.5.
The original definition of germ, see [CZ20, Definition 4.1] is given for general distributions without mentioning the order. This different choice will be motivated by a slightly different analytic setting we will use to prove the reconstruction theorem. As illustrated in the definition, we will consistently write in subscript the evaluation of a germ on a point $\mathbf{x}$. We can equivalently represent a germ $F$ as a family of distributions $F=(F_{\mathbf{x}})_{\mathbf{x}\in{\mathbb{R}}^{d}}$.
For any given germ $F=(F_{\mathbf{x}})_{\mathbf{x}\in{\mathbb{R}}^{d}}$, we can transfer (2.2) by fixing a test function $\psi$ and applying the definition to the function $f(\mathbf{x})=F_{\mathbf{x}}(\psi)$.
Let us introduce the notion of scaling in our setting. Given a vector $\mathbf{x}\in{\mathbb{R}}^{d}$ we want to rescale a test function $\psi\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$ with respect to a generic vector ${\bm{\lambda}}\in(0,1]^{d}$ with the definition
$\psi_{\mathbf{x}}^{\bm{\lambda}}(\mathbf{y}):={\bm{\lambda}}^{-{\bm{1}}}\psi((y_{1}-x_{1})\lambda_{1}^{-1}\,,\ldots\,,(y_{d}-x_{d})\lambda_{d}^{-1})\,.$ (2.7)
For $\mathbf{x}=\mathbf{0}$ we write $\psi^{\bm{\lambda}}(\mathbf{y})=\psi_{\mathbf{0}}^{\bm{\lambda}}(\mathbf{y})$. For any given $T\in\mathcal{D}_{r}^{\prime}$ the map ${\bm{\lambda}}\to T(\psi_{\mathbf{x}}^{\bm{\lambda}}(\mathbf{y}))$ describes the behavior of $T$ close to $\mathbf{x}$ at different scales.
2.2 Wavelet bases
To prove the reconstruction theorem in the rectangular increment setting we will have to use all the regularising effects of wavelet theory, going back to [Dau88]. First, we will recall some general facts about wavelet basis, mainly following [Hai14, Ker23, Mey93].
Multidimensional wavelets can be constructed from one-dimensional wavelets, which are fundamental tools to describe the Hilbert space $L^{2}({\mathbb{R}}^{d})$ with its standard scalar product
[…]
While the scaling introduced in the last section preserves the $L^{1}(\mathbb{R}^{d})$-norm this does not hold true for the $L^{2}(\mathbb{R}^{d})$ norm. For wavelets $\phi$ we will instead use the $L^{2}$-scaling: Given any $d$-dimensional wavelet $\phi$ or $\hat{\phi}^{\zeta}$ (see (2.10) for the definition), $\mathbf{x}\in\mathbb{R}^{d}$ and any multi-index $\mathbf{n}\in\mathbb{N}^{d}$, we set
$\phi^{\mathbf{n}}_{\mathbf{x}}(\mathbf{y}):=2^{\frac{1}{2}\mathbf{n}}\phi(2^{n_{1}}(y_{1}-x_{1}),\ldots,2^{n_{d}}(y_{d}-x_{d}))\,,$
where we define $\hat{\phi}^{\zeta,\mathbf{n}}_{\mathbf{x}}$ analogously. Using the notation $\Delta_{n}:=\{2^{-n}k\leavevmode\nobreak\ \colon\leavevmode\nobreak\ k\in\mathbb{Z}\}$ the one-dimensional grid of mesh size $2^{-n}$ for all $n\in\mathbb{N}$, we recall basic results of wavelet theory which we will use as starting point of our analysis, see [Mey93, Chapter 3] for the proof.
Theorem 2.6.
For any $r>0$ there exist two functions $\varphi$, $\hat{\varphi}\in C^{r}_{c}(\mathbb{R})$ such that
  1. 1.
    $\varphi$ is self replicable: i.e. there exists finitely many $a_{k}\in\mathbb{R},a_{k}\neq 0$, $k\in\Delta_{1}$ such that
    $\varphi=\sum_{k\in\Delta_{1}}a_{k}\varphi^{1}_{k}\,.$ (2.8)
  2. 2.
    For all integer $0\leq m\leq r$,
[…]
  3. 3.
    For all $n\in\mathbb{N}$, the set $\{\varphi^{n}_{x}\,,\hat{\varphi}^{m}_{y}\,\colon\,m\geq n\,,x\in\Delta_{n}\,,y\in\Delta_{m}\}$ is a complete orthonormal system of $L^{2}({\mathbb{R}})$.
  4. 4.
    For all $n\in\mathbb{N}$, the $\text{span}\{\hat{\varphi}_{x}^{n}\colon x\in\Delta_{n}\}$ is the orthogonal complement of $\text{span}\{\varphi_{x}^{n}\leavevmode\nobreak\ \colon x\in\Delta_{n}\}$ as a subspace of $\text{span}\{\varphi_{x}^{n+1}\colon x\in\Delta_{n+1}\}$. More precisely, any linear combination
    $\psi=\sum_{x\in\Delta_{n+1}}c_{x}\varphi^{n+1}_{x}$
    where $c_{x}=0$ for all but finitely many $x$ holds can be written as a finite linear combination
    $\psi=\sum_{x\in\Delta_{n}}b_{x}\varphi^{n}_{x}+\sum_{x\in\Delta_{n}}\hat{b}_{x}\hat{\varphi}_{x}^{n}\,.$
  5. 5.
    For all $\psi\in C_{c}^{r}$ one has the convergence
    $\psi$ $=\lim_{m\to\infty}\sum_{x\in\Delta_{m}}\varphi_{x}^{m}\left\langle\varphi_{x}^{m},\psi\right\rangle=\sum_{x\in\Delta_{n}}\varphi_{x}^{n}\left\langle\varphi_{x}^{n},\psi\right\rangle+\sum_{m\geq n}\sum_{x\in\Delta_{m}}\hat{\varphi}_{x}^{m}\left\langle\hat{\varphi}_{x}^{m},\psi\right\rangle\,,$ (2.9)
    where the limits hold in $L^{2}(\mathbb{R})$
We now pass to the multi-dimensional setting. For any integer $d\geq 1$ and multi-index $\mathbf{n}\in\mathbb{N}^{d}$ we set the $\mathbf{n}$-th grid to be
$\Delta_{\mathbf{n}}$ :=$:=\left\{\sum_{i=1}^{d}2^{-n_{i}}k_{i}e_{i}\leavevmode\nobreak\ \colon\leavevmode\nobreak\ \mathbf{k}\in\mathbb{Z}^{d}\right\}\subset\mathbb{R}^{d}\,,$
where $e_{i}$ is the canonical basis of $\mathbb{R}^{d}$. Similarly, for any subset $\zeta\subset[d]$ we can also project the $\mathbf{n}$-th grid to
$\Delta^{\zeta}_{\mathbf{n}}:=\pi^{\zeta}_{\mathbf{0}}(\Delta_{\mathbf{n}})\,.$
We will use the letter $\zeta$ to usually describe subsets regarding wavelets and will reserve $\theta,\eta$ as much as possible for the rectangular increments of a germ. The subset notation factorizes easily at the level of wavelets
Definition 2.7.
For any $r>0$ and $\zeta\subset[d]$, we define the $d$-dimensional wavelets $\phi$, $\hat{\phi}\in C^{r}_{c}(\mathbb{R}^{d})$ by setting
$\begin{split}\phi(\mathbf{x})=\hat{\phi}^{\emptyset}(\mathbf{x})&:=\prod_{i=1}^{d}\varphi(x_{i})\,,\quad\hat{\phi}^{\zeta}(\mathbf{x}):=\prod_{i\notin\zeta}\varphi(x_{i})\prod_{i\in\zeta}\hat{\varphi}(x_{i})\,,\end{split}$ (2.10)
where the functions $\varphi$, $\hat{\varphi}$ are given by Theorem (2.6).
As in [Ker23], we can replace the one-dimensional wavelets $\varphi,\hat{\varphi}$ with shifted wavelets $\varphi_{x},\hat{\varphi}_{x}$ for any $x\in\mathbb{R}$ without losing any properties of the wavelets. Thus we can assume that the wavelets $\phi,\hat{\phi}$ have a support in $[C,R]^{d}$ for some constants $0<C<R$ and that $a_{k}\neq 0$ only holds for $k\geq 0$.
We set $V_{\mathbf{n}}:=\text{span}\{\phi_{\mathbf{x}}^{\mathbf{n}}\colon\mathbf{x}\in\Delta_{\mathbf{n}}\}$ and $\hat{V}_{\mathbf{n}}^{\zeta}:=\text{span}\{\hat{\phi}^{\zeta,\mathbf{n}}_{\mathbf{x}}\colon x\in\Delta_{\mathbf{n}}\}$  and we denote by $P_{\mathbf{n}},\hat{P}_{\mathbf{n}}^{\zeta}$ the projections into the respective spaces. Iterating (2.9), one has formally
$\psi$ $=\lim_{\mathbf{n}\to\infty}P_{\mathbf{n}}\psi=\sum_{\zeta\subset[d]}\sum_{\mathbf{k}\in\pi_{\mathbf{0}}^{[d]\setminus\zeta}(\mathbb{N}^{d})}\hat{P}^{\zeta}_{\mathbf{n}+\mathbf{k}}\psi$
and it is easy to see that these equalities do hold for all $\psi\in C_{c}^{r}(\mathbb{R}^{d})$ as limits in $L^{2}(\mathbb{R}^{d})$. The limit $\mathbf{n}\to\infty$ stands for a limit along any sequence $(\mathbf{n}^{(k)})_{k\in\mathbb{N}^{d}}$ such that $n^{(k)}_{i}$ goes to $\infty$ for each $i\in[d]$. By [Ker23, Lemma 62], the above limits also hold in $C_{c}^{q}$ for $q<r$ as long as $r$ is big enough, as well as
$f(\psi)$ $=\lim_{\mathbf{n}\to\infty}f(P_{\mathbf{n}}(\psi))=\sum_{\zeta\subset[d]}\sum_{\mathbf{k}\in\pi_{\mathbf{0}}^{[d]\setminus\zeta}(\mathbb{N}^{d})}f(\hat{P}^{\zeta}_{\mathbf{n}+\mathbf{k}}\psi)$
holds for all (random) distributions $f$. This convergence also comes with quantitative bounds. To calculate those, we use the fact that our wavelets eliminate all functions, which are polynomials of degree $\leq r$ in at least one index $i\in\zeta$. More precisely, for any non- empty indexset $\zeta\subset[d]$, $i\in\zeta$ and level $m\leq r$, it holds that $\left\langle x_{i}^{m}\pi_{\mathbf{0}}^{i}(\psi),\hat{\phi}^{\zeta}\right\rangle=0$ for any test function $\psi\in C_{c}^{r}(\mathbb{R}^{d})$. We use this to show the following estimate:
Lemma 2.8.
Let $r>0$ and let $\psi\in C_{c}^{rd}$. Then:
$\left|\left\langle\hat{\phi}^{\zeta,\mathbf{n}}_{\mathbf{y}},\psi^{{\bm{\lambda}}}_{\mathbf{z}}\right\rangle\right|\lesssim 2^{-r\mathbf{n}_{\zeta}-\frac{1}{2}\mathbf{n}}{\bm{\lambda}}^{-r\mathbf{1_{\zeta}}-{\bm{1}}}\left\|\psi\right\|_{C_{c}^{rd}}$
uniformly in $\mathbf{n}\in\mathbb{N}^{d}$, $\mathbf{y}\in\Delta_{\mathbf{n}}$, $\mathbf{z}\in\mathbb{R}^{d}$, $\zeta\subset[d]$ non empty and ${\bm{0}}<{\bm{\lambda}}\leq{\bm{1}}$.
Proof.
For $i\in[d]$, we denote by $T^{i}_{\mathbf{z}}$ the Taylor-polynomial of level $r-1$ in the $i$-th coordinate, i.e.
$T^{i}_{\mathbf{z}}\psi(\mathbf{y})=\sum_{k=0}^{r-1}\frac{1}{k!}\frac{\partial^{k}}{\partial x_{i}^{k}}\psi(\pi_{z_{i}}^{i}\mathbf{y})(y_{i}-z_{i})^{k}\,.$
Extending multiplicatively $T^{i}_{\mathbf{z}}$ we denote by $T^{\theta}_{\mathbf{z}}:=\prod_{i\in\theta}T^{i}_{\mathbf{z}}$. Then it holds from the usual Taylor formula with integral remainder that
$\prod_{i\in\theta}(\mbox{Id}-T^{i}_{\mathbf{z}})(\psi)(\mathbf{y})=R^{\theta}(\mathbf{z},\mathbf{y})$
with the remainder function $R^{\theta}$ given by
$R^{\theta}(\mathbf{z},\mathbf{y})=\int_{\mathbf{z}_{\theta}}^{\mathbf{y}_{\theta}}\left(\frac{1}{r!}\right)^{\sharp\theta}\partial^{r\mathbf{1}_{\theta}}\psi(\pi^{\theta}_{\mathbf{t}_{\theta}}\mathbf{y})\prod_{i\in\theta}(y_{i}-t_{i})^{r-1}d\mathbf{t}_{\theta}\,.$
By hypothesis on $\psi$ it follows that
$\left|R^{\theta}(\mathbf{z},\mathbf{y})\right|\lesssim\left\|\psi\right\|_{C_{c}^{rd}}\left|\mathbf{y}-\mathbf{z}\right|^{r}_{\text{c},\theta}\,.$
Since $\hat{\phi}^{\zeta}$ integrate to zero when tested against the polynomials $T^{\eta}_{\mathbf{z}}\psi$ for all $\emptyset\neq\eta\subset\zeta$, we have
$\left\langle\hat{\phi}^{\zeta},\psi\right\rangle=\left\langle\hat{\phi}^{\zeta},\prod_{i\in\zeta}(\mbox{Id}-T^{i}_{\mathbf{z}})\psi\right\rangle\,.$
Considering now the rescaling, we set ${\bm{\gamma}}:=(2^{-n_{1}},\dots,2^{-n_{d}})$ and ${\bm{\mu}}:=(\lambda_{1}^{-1},\dots,\lambda_{d}^{-1})$ and by applying a classical change of variable we calculate
[…]
Remark 2.9.
Note that the convergence speed is faster than one would expect: If one starts with any function $\phi$, one would expect the rates $2^{-\mathbf{n}/2}{\bm{\lambda}}^{-{\bm{1}}}$ simply from the localizations of $\phi^{\mathbf{n}}_{\mathbf{y}}$ and $\psi^{\bm{\lambda}}_{\mathbf{z}}$. However, the fact that $\phi^{\zeta}$ is orthogonal to polynomials of degree $\leq r$ in some $i\in\zeta$ allows us to get the additional factor $2^{-(r\mathbf{n})_{\zeta}}{\bm{\lambda}}^{-(r\mathbf{1})_{\zeta}}$. This is a more precise version of [Ker23, Lemma 61] (which itself is derived from [Hai14, Proposition 3.20]), where they only used the weaker condition, that all $\phi^{\zeta}$ for $\zeta\neq\emptyset$ eliminate multivariate polynomials of degree $\leq r$.
2.3 Commuting filtration
We now recall the probabilistic assumptions to formulate our results. Throughout this paper, we fix a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Given $m\geq 1$ the set of $m$- integrable random variables $X\colon\Omega\to{\mathbb{R}}$ will be denoted by $L_{m}(\Omega)$ and its norm by $\|\cdot\|_{m}$. We will keep this discussion rather short, as it has been done in length in [BHK23]. One can find more classical references in [Imk85, Wal86, Kho02].
We equip our probability space with a filtration associated with the rectangular increment setting.
Definition 2.10.
A family of sigma-algebras $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in{\mathbb{R}}^{d}},\mathcal{F}_{\mathbf{t}}\subset\mathcal{F}$ for all $\mathbf{t}\in\mathbb{R}^{d}$, is called a multiparameter filtration, if for all $\mathbf{s}\leq\mathbf{t}$, we have $\mathcal{F}_{\mathbf{s}}\subset\mathcal{F}_{\mathbf{t}}$.
Remark 2.11.
Alternatively, it is possible to consider filtrations indexed by $\mathbf{t}\in K$ with $K\subset{\mathbb{R}}^{d}$ a compact set. We will use that when considering SPDEs and all definitions will extend trivially.
Definition 2.12.
For any given multiparameter filtration $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in{\mathbb{R}}^{d}}$ and subset $\theta\subset[d]$, we set
$\mathcal{F}^{\theta}_{\mathbf{t}}:=\bigvee_{\mathbf{s}\in{\mathbb{R}}^{d}}\mathcal{F}_{\pi^{\theta}_{\mathbf{t}}\mathbf{s}}\,,$
where $\bigvee$ represents the sigma-algebra generated by the union. When $\theta=\{i\}$ we call $\mathcal{F}^{i}:=\mathcal{F}^{\{i\}}$ the $i$-th marginal filtration of $(\mathcal{F}_{\mathbf{t}})$.
We denote by $\mathbb{E}^{\theta}_{\mathbf{t}}$ the conditional expectation $\mathbb{E}[\cdot|\mathcal{F}_{\mathbf{t}}^{\theta}]$. A natural condition to impose on a filtration $\mathcal{F}$ to compute $\mathbb{E}[\cdot|\mathcal{F}_{\mathbf{t}}^{\theta}]$ efficiently is the notion of commutativity. I.e. we assume that for all $i,j\in[d]$ and $\mathbf{s},\mathbf{t}\in{\mathbb{R}}^{d}$ one has the a.s. identity
$\mathbb{E}^{i}_{\mathbf{t}}\mathbb{E}^{j}_{\mathbf{s}}=\mathbb{E}^{j}_{\mathbf{s}}\mathbb{E}^{i}_{\mathbf{t}}\,.$ (2.11)
We call a filtration $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in\mathbb{R}^{d}}$ fulfilling (2.11) a commuting filtration. It should be noted that in the literature, several equivalent properties to (2.11) are used. Walsh introduced in [CW75, Section 1], the condition that for all $\mathbf{x}\in\mathbb{R}^{d}$, $\mathcal{F}^{i}_{\mathbf{x}},i=1,\dots,d$ are conditionally independent on $\mathcal{F}_{\mathbf{x}}$. [Kho02, Section 3.4] prefers to assume that for all bounded random variables $X:\Omega\to\mathbb{R}$ which are $\mathcal{F}_{\mathbf{t}}$ measurable for some $\mathbf{t}\in\mathbb{R}^{d}$, we have
$\mathbb{E}^{[d]}_{\mathbf{s}}X=\mathbb{E}^{[d]}_{\mathbf{s}\wedge\mathbf{t}}X\,,$
for all $\mathbf{s}\in{\mathbb{R}}^{d}$. In both cases, the authors show that their condition is equivalent to (2.11). We recall two basic results of multiparameter filtrations.
Proposition 2.13.
Let $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in{\mathbb{R}}^{d}}$ be a multiparameter filtration. We then have the following properties
  • i)
    If $(\mathcal{F}_{\mathbf{t}})$ is complete, then so is $(\mathcal{F}_{\mathbf{t}}^{\theta})$ for all subsets $\theta\subset[d]$.
  • ii)
    If $\mathcal{F}$ is commuting, then for all subsets $\theta\subset[d]$
    $\mathbb{E}^{\theta}_{\mathbf{t}}=\prod_{i\in\theta}\mathbb{E}^{i}_{\mathbf{t}}.$
Proof.
i) immediately follows from $\mathcal{F}_{\mathbf{t}}\subset\mathcal{F}_{\mathbf{t}}^{\theta}$ for all $\mathbf{t}\in\mathbb{R}^{d},\theta\subset[d]$. ii) is shown in [Kho02] for $\theta=[d]$ in Theorem 3.4.1. The case $\theta\subset[d]$ follows with the same proof. ∎
The stochastic reconstruction theorem, as the stochastic sewing lemma, is based on BDG-type inequalities which in $d=1$ can be found in [Le20], Equation (2.5). The multiparameter version of this inequality has been found in [BHK23, Lemma 25].
Lemma 2.14 (extended BDG inequality).
Let $i\in\{1,\dots,d\}$ and $I_{i}\subset\mathbb{N}$ be a finite set with minimal element $y_{i}$. Define for each indexset $\theta\subset[d]$
$I_{\theta}=\prod_{i\in\theta}I_{i}\times\prod_{j\notin\theta}\{y_{j}\}:=\{\mathbf{x}\in\mathbb{N}^{d}\leavevmode\nobreak\ \colon\leavevmode\nobreak\ x_{i}\in I_{i}\text{ if $i\in\theta$, }x_{i}=y_{i}\text{ otherwise}\}\,.$
Further, let $(\mathcal{F}_{\mathbf{n}})_{\mathbf{n}\in\mathbb{N}^{d}}$ be a commuting $d$-parameter filtration and let $Z_{\mathbf{k}}\in L_{m}(\Omega)$ for each $\mathbf{k}\in I_{[d]}$. We assume that $Z_{\mathbf{k}}$ is $\mathcal{F}_{\mathbf{k}+\bm{1}}$ measurable. Assume that there exist constants $0<b_{i,k_{i}},a_{i,k_{i}}$, $i\in[d],k_{i}\in I_{i}$, as well as a $c>0$, such that for each $\theta\subset[d]$:
$\left\|\mathbb{E}^{\theta}_{\mathbf{k}}Z_{\mathbf{k}}\right\|_{m}\leq c\,b_{\mathbf{k},\eta}\,a_{\mathbf{k},\eta^{c}}\,,$ (2.12)
where $b_{\mathbf{k},\eta}:=\prod_{i\in\eta}b_{i,k_{i}}$ and $a_{\mathbf{k},\eta^{c}}=\prod_{i\in\eta^{c}}a_{i,k_{i}}$. Then, it follows that for each $\theta,\eta\subset[d]$:
$\left\|\mathbb{E}^{\eta}_{\mathbf{y}}\sum_{\mathbf{k}\in I_{\theta}}Z_{\mathbf{k}}\right\|_{m}\lesssim c\sum_{\theta\setminus\eta=\theta_{1}\sqcup\theta_{2}}\left(\sum_{\mathbf{k}\in I_{\theta\setminus\theta_{2}}}b_{\mathbf{k},\eta\cup\theta_{1}}\right)\left(\sum_{\mathbf{l}\in I_{\theta_{2}}}a_{\mathbf{l},(\eta\cup\theta_{1})^{c}}^{2}\right)^{\frac{1}{2}}\,,$ (2.13)
where we sum over all disjoint $\theta_{1},\theta_{2}$, such that $\theta\setminus\eta=\theta_{1}\cup\theta_{2}$. It especially follows
$\left\|\sum_{\mathbf{k}\in I_{\theta}}Z_{\mathbf{k}}\right\|_{m}\lesssim c\sum_{\theta=\theta_{1}\sqcup\theta_{2}}\left(\sum_{\mathbf{k}\in I_{\theta_{1}}}b_{\mathbf{k},\theta_{1}}\right)\left(\sum_{\mathbf{l}\in I_{\theta_{2}}}a_{\mathbf{l},\theta_{2}}^{2}\right)^{\frac{1}{2}}\,.$
[…]
Definition 2.15.
We say that a random germ $(F_{\mathbf{x}})$ is adapted to a multiparameter filtration $(\mathcal{F}_{\mathbf{x}})$, if for all $\mathbf{x}\in\mathbb{R}^{d}$ and test function $\psi\in C_{c}^{r}$ with support in $(-\infty,\mathbf{y}]$, $F_{\mathbf{x}}(\psi)$ is $\mathcal{F}_{\mathbf{x}\vee\mathbf{y}}$-measurable.
3 The stochastic reconstruction theorem
We now pass to formulating the main theorem of the paper. To make it rigorous we will have to introduce the proper spaces of functions and distributions where the reconstruction map acts and describe their main properties. We fix a time horizon $T>0$ throughout this section.
3.1 Hölder spaces for random distributions
We start by recalling the deterministic multiparameter Hölder spaces, see [QST07, Har21] for further references.
Given ${\bm{\alpha}}\in[0,1]^{d}$ we set $C^{\bm{\alpha}}$ to be the space of all maps $Y\colon[0,T]^{d}\to\mathbb{R}$, such that
$\left\|Y\right\|_{C^{\bm{\alpha}}}:=\sum_{\theta\subset[d]}\left\|\square^{\theta}Y\right\|_{{\bm{\alpha}},\theta}<\infty\,,\quad\left\|\square^{\theta}Y\right\|_{{\bm{\alpha}},\theta}:=\sup_{\mathbf{x},\mathbf{y}\in[0,T]^{d},\,\mathbf{x}\neq\mathbf{y}}\frac{|\square^{\theta}_{\mathbf{x},\mathbf{y}}Y|}{|\mathbf{x}-\mathbf{y}|^{{\bm{\alpha}}}_{\text{c},\theta}}\,.$ (3.1)
This definition can be extended naturally to any ${\bm{\alpha}}\geq{\bm{0}}$ but in our case, we will always consider ${\bm{\alpha}}\leq{\bm{1}}$. On the other hand, we will also consider multi-index Hölder spaces where the components of ${\bm{\alpha}}$ are negative. In this case we set $r$ as the smallest possible integer such that $r\geq\sum_{i=1}^{d}\left|\alpha_{i}\right|$ and we look at continuous linear maps over $f:C_{c}^{r}([0,T]^{d})\to{\mathbb{R}}$. More precisely, we set $C^{{\bm{\alpha}}}([0,T]^{d})$ as the set of all $f\in\mathcal{D}^{r}([0,T]^{d})$ such that
$\left\|f\right\|_{{\bm{\alpha}}}=\sup_{\mathbf{x},\psi,{\bm{\lambda}}\in(0,1]}\frac{\left|f(\psi_{\mathbf{x}}^{{\bm{\lambda}}})\right|}{{\bm{\lambda}}^{{\bm{\alpha}}}}<\infty\,,$
where the supremum runs over all $\mathbf{x}\in[0,T]^{d},\psi\in B^{r}([0,T]^{d})$ and ${\bm{0}}<{\bm{\lambda}}\leq{\bm{1}}$ such that $\psi_{\mathbf{x}}^{{\bm{\lambda}}}$ has support in $[0,T]^{d}$.
To extend these spaces when there is a filtered probability space $(\Omega,\mathcal{F},P)$ with a commuting $d$-parameter filtration $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in[0,T]^{d}}$, we need to condition the random distributions over a careful choice of filtrations. More specifically, we want to quantify the seminorm $\square^{\theta}_{\mathbf{x},\mathbf{y}}Y$ in (3.1) with a $L_{m}$ norm and additionally we want the conditional expectation $\mathbb{E}^{\eta}_{\mathbf{x}}$ to add another regularity factor.
Definition 3.1.
Let ${\bm{\delta}}\geq{\bm{0}}$, ${\bm{\alpha}}\in[0,1]^{d}$ and $m\geq 1$. We then set $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ to be the space of random fields $Y:[0,T]^{d}\times\Omega\to\mathbb{R}$ such that for all $\mathbf{t}\in[0,T]^{d}$, $Y_{\mathbf{t}}$ is $\mathcal{F}_{\mathbf{t}}$-measurable and for all $\eta\subset\theta\subset[d]$ one has
$\left\|\square Y\right\|_{\theta,\eta,m}$ $:=\sup_{\mathbf{x},\mathbf{y}\in[0,T]^{d},\;x_{i}\leq y_{i}\text{for }i\in\theta}\frac{\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square^{\theta}_{\mathbf{x},\mathbf{y}}Y\right\|_{m}}{\left|\mathbf{x}-\mathbf{y}\right|^{{\bm{\alpha}}}_{\text{c},\theta}\left|\mathbf{x}-\mathbf{y}\right|^{{\bm{\delta}}}_{\text{c},\eta}}<\infty\,.$ (3.2)
In most cases, ${\bm{\alpha}},{\bm{\delta}}$, and $T$ are clear from the context. If they are not, we will write $\left\|\square Y\right\|_{{\bm{\alpha}},{\bm{\delta}},\theta,\eta,m}$ and $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,T]^{d})$. Recall that for $\theta=\emptyset$, the above quantities simplify to
$\left\|Y\right\|_{\emptyset,\emptyset,m}=\sup_{\mathbf{x}\in[0,T]^{d}}\left\|Y_{\mathbf{x}}\right\|_{m}=:\left\|Y\right\|_{m,\infty}\,.$
The space $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ is trivially a Banach space with respect to the norm
$\left\|Y\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}$ $:=\left\|Y\right\|_{m,\infty}+\left\llbracket Y\right\rrbracket_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\,,$
$\left\llbracket Y\right\rrbracket_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}$ $:=\sum_{\eta\subset\theta\subset[d],\theta\neq\emptyset}\left\|\square Y\right\|_{\theta,\eta,m}\,.$
Note that $\left\llbracket Y\right\rrbracket$ is not a norm on $Y$ but only a semi-norm over $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$. For ${\bm{\delta}}={\bm{0}}$, we write $C^{{\bm{\alpha}}}L_{m}:=C^{{\bm{\alpha}},{\bm{0}}}L_{m}$. Further observe that strong martingales as described by [Imk85], [Wal86] fulfill (3.2) uniformly for all ${\bm{\delta}}>{\bm{0}}$. In this case, we set
$\left\|Y\right\|_{C^{{\bm{\alpha}},\infty}L_{m}}=\sup_{{\bm{\delta}}>{\bm{0}}}\left\|Y\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}$
and write $Y\in C^{{\bm{\alpha}},\infty}L_{m}$ if $\left\|Y\right\|_{C^{{\bm{\alpha}},\infty}L_{m}}<\infty$.
We now pass on describing the embedding properties of these spaces. Concerning the index ${\bm{\alpha}}$ the spaces $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ behave exactly as the classical Hölder spaces. That is, for ${\bm{0}}<{\bm{\alpha}}<{\bm{\beta}}<\mathbf{1}$ one has $C^{{\bm{\beta}},{\bm{\delta}}}L_{m}\subset C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$. Moreover, it follows trivially from the definition (3.2) that for any $\eta\subset\theta\subset[d],\theta\neq\emptyset$ we have
$\left\|\square Y\right\|_{{\bm{\alpha}},{\bm{\delta}},\theta,\eta,m}\leq(T\mathbf{1})^{{\bm{\beta}}-{\bm{\alpha}}}_{\theta}\left\|\square Y\right\|_{{\bm{\beta}},{\bm{\delta}},\theta,\eta,m}\,.$ (3.3)
This estimate gets even stronger if $Y$ vanishes on ${\bm{0}}$ or alternatively on the whole boundary $\partial_{{\bm{0}}}[0,T]^{d}:=\bigcup_{i=1}^{d}\{\mathbf{x}\in[0,T]^{d}\colon x_{i}=0\}$.
Proposition 3.2.
For any ${\bm{0}}<{\bm{\alpha}}<{\bm{\beta}}<\mathbf{1}$ we have $C^{{\bm{\beta}},{\bm{\delta}}}L_{m}([0,T]^{d})\subset C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,T]^{d})$. Moreover, we have the following estimates:
  • for any $Y\in C^{{\bm{\beta}},{\bm{\delta}}}L_{m}$ such that $Y({\bm{0}})=0$
    $\left\|Y\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\lesssim\left(\sum_{\emptyset\neq\theta\subset[d]}(T{\bm{1}})^{{\bm{\beta}}}_{\theta}+\sum_{\emptyset\neq\theta\subset[d]}(T\mathbf{1})^{({\bm{\beta}}-{\bm{\alpha}})}_{\theta}\right)\left\|Y\right\|_{C^{{\bm{\beta}},{\bm{\delta}}}L_{m}}\,.$ (3.4)
  • for any $Y\in C^{{\bm{\beta}},{\bm{\delta}}}L_{m}$ such that $Y(\mathbf{x})=0$ for any $\mathbf{x}\in\partial_{{\bm{0}}}[0,T]^{d}$
    $\left\|Y\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\lesssim\left(\sum_{\theta\subset[d]}(T\mathbf{1})^{{\bm{\beta}}-{\bm{\alpha}}_{\theta}}\right)\left\|Y\right\|_{C^{{\bm{\beta}},{\bm{\delta}}}L_{m}}\,.$ (3.5)
Proof.
Let us first show (3.4). Applying Lemma 2.2, we can write
$Y(\mathbf{x})=Y({\bm{0}})+\sum_{\emptyset\neq\theta\subset[d]}(-1)^{\sharp(\theta^{c})}\square^{\theta}_{\pi_{\mathbf{x}}^{\theta^{c}}{\bm{0}},\mathbf{x}}Y\,.$
Thus from the condition $Y({\bm{0}})=0$ and the usual triangular inequality one immediately has
$\left\|Y\right\|_{m,\infty}\leq\sum_{\emptyset\neq\theta\subset[d]}(T{\bm{1}})^{{\bm{\beta}}}_{\theta}\left\|\square Y\right\|_{{\bm{\beta}},{\bm{\delta}},\theta,\emptyset,m}\,.$
The final inequality (3.4) then follows from the above together with (3.3).
We pass to (3.5). If $Y(\mathbf{x})=0$ for all $\mathbf{x}\in B$, we get that
$Y(\mathbf{x})=\square_{{\bm{0}},\mathbf{x}}^{[d]}Y$
which immediately gives us
$\left\|Y\right\|_{m,\infty}\leq(T{\bm{1}})^{{\bm{\beta}}}\left\|\square Y\right\|_{{\bm{\beta}},{\bm{\delta}},[d],\emptyset,m}\,.$
Furthermore, from the same hypothesis we have for all $\mathbf{x},\mathbf{y}\in[0,T]^{d}$ and $\theta\subset[d]$
$\square_{\mathbf{x},\mathbf{y}}^{\theta}Y=\square^{[d]}_{\pi^{\theta}_{\mathbf{x}}{\bm{0}},\mathbf{y}}Y\,,$
which immediately gives one for any $\eta\subset\theta\subset[d]$ the inequality
$\left\|\square Y\right\|_{{\bm{\alpha}},{\bm{\delta}},\theta,\eta,m}\lesssim(T{\bm{1}})^{{\bm{\beta}}-{\bm{\alpha}}_{\theta}}\left\|\square Y\right\|_{{\bm{\beta}},{\bm{\delta}},[d],\eta,m}\,.$
Summing both sides over $\eta\subset\theta\subset[d],\theta\neq\emptyset$ we derive (3.5). ∎
Using the same reasoning, we can also describe the embedding properties of $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ with respect to the parameter ${\bm{\delta}}$.
Proposition 3.3.
Let ${\bm{\alpha}}\in[0,1]^{d}$ and let $0\leq{\bm{\delta}}\leq{\bm{\gamma}}$. Then we have $C^{{\bm{\alpha}},{\bm{\gamma}}}L_{m}([0,T]^{d})\subset C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,T]^{d})$ with the estimates:
$\left\|Y\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\lesssim\left(\sum_{\eta\subset[d]}T_{\eta}^{{\bm{\gamma}}-{\bm{\delta}}}\right)\left\|Y\right\|_{C^{{\bm{\alpha}},{\bm{\gamma}}}L_{m}}\,.$
where the constant in $\lesssim$ only depends on $d$.
For negative ${\bm{\alpha}}<{\bm{0}}$, we no longer have an equivalent of $\square^{\theta}Y$ for $\theta\subsetneq[d]$. Thus, we can forget one of the indexsets in Definition 3.1 and naturally adapt the deterministic definition into a stochastic setting.
Definition 3.4.
Let ${\bm{\alpha}}<{\bm{0}}$, $1\leq m$ and ${\bm{\delta}}\geq{\bm{0}}$. We define $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ as the space of all random distributions $f\in\mathcal{D}^{r}$ with $r$ chosen as in the deterministic setting, satisfying the following properties:
  • For any test function $\psi$ with support in $(-\infty,\mathbf{x}]$, $f(\psi)$ is $\mathcal{F}_{\mathbf{x}}$-measurable.
  • for all $\psi\in B^{r}([1/4,3/4]^{d})$ and ${\bm{\lambda}}\in(0,1]^{d}$, $\mathbf{x}\in[0,T]^{d}$ such that $[\mathbf{x},\mathbf{x}+{\bm{\lambda}}]\subset[0,T]^{d}$ we have uniformly over $\eta\subset[d]$:
    $\begin{split}\left\|f(\psi^{{\bm{\lambda}}}_{\mathbf{x}})\right\|_{m}&\lesssim{\bm{\lambda}}^{{\bm{\alpha}}}\\ \left\|\mathbb{E}^{\eta}_{\mathbf{x}}f(\psi^{{\bm{\lambda}}}_{\mathbf{x}})\right\|_{m}&\lesssim{\bm{\lambda}}^{{\bm{\alpha}}}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}\,,\end{split}$ (3.6)
$C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ becomes a Banach space by using the norm
$\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}:=\sup_{\psi,\mathbf{x},{\bm{\lambda}},\eta\subset[d]}\frac{\left\|\mathbb{E}^{\eta}_{\mathbf{x}}f(\psi^{{\bm{\lambda}}}_{\mathbf{x}})\right\|_{m}}{{\bm{\lambda}}^{{\bm{\alpha}}}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}}\,,$
where the supremum runs over the $\psi,\mathbf{x},{\bm{\lambda}}$ as described above. As in the positive case, we use the notation $C^{{\bm{\alpha}}}L_{m}:=C^{{\bm{\alpha}},{\bm{0}}}L_{m}$ and write $f\in C^{{\bm{\alpha}},\infty}L_{m}$ if
$\left\|f\right\|_{C^{{\bm{\alpha}},\infty}L_{m}}=\sup_{{\bm{\delta}}>{\bm{0}}}\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}<\infty\,.$
Remark 3.5.
Note that wavelet approximations $P_{\mathbf{n}}\psi$ enlarge the support of $\psi$. Thus, by using test functions with support in $[1/4,3/4]^{d}$ we can ensure that $P_{\mathbf{n}}(\psi_{\mathbf{x}}^{\bm{\lambda}})$ has support in $[\mathbf{x},\mathbf{x}+{\bm{\lambda}}]$ for any $\mathbf{x}\in\mathbb{R}^{d},{\bm{\lambda}}\in(0,1]^{d}$ and large enough $\mathbf{n}\in\mathbb{N}^{d}$. This is especially important to calculate the Hölder regularity of the reconstruction, see Corollary 3.18. The constants $1/4,3/4$ are chosen arbitrarily, any $10<C<R<$ 1 would work the same.
Let us give an easy but useful result regarding the sum of distributions in different $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ spaces.
Lemma 3.6.
Let $f\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ and $g\in C^{\tilde{\bm{\alpha}},\tilde{\bm{\delta}}}L_{m}$ for ${\bm{\delta}},\tilde{\bm{\delta}}\geq 0$ and ${\bm{\alpha}},\tilde{\bm{\alpha}}<0$. Then we have $f+g\in C^{{\bm{\beta}},{\bm{\gamma}}}L_{m}$ where
${\bm{\beta}}={\bm{\alpha}}\wedge\tilde{\bm{\alpha}}\,,\quad{\bm{\gamma}}=({\bm{\alpha}}+{\bm{\delta}})\wedge(\tilde{\bm{\alpha}}+\tilde{\bm{\delta}})-{\bm{\beta}}\,.$
Furthermore, we have the estimate
$\left\|f+g\right\|_{C^{{\bm{\beta}},{\bm{\gamma}}}L_{m}}\leq\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}+\left\|g\right\|_{C^{\tilde{\bm{\alpha}},\tilde{\bm{\delta}}}L_{m}}\,.$ (3.7)
Proof.
Simple calculation gives
$\left\|\mathbb{E}_{\mathbf{x}}^{\eta}(f+g)(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{m}$ $\leq\left\|\mathbb{E}_{\mathbf{x}}^{\eta}f(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{m}+\left\|\mathbb{E}_{\mathbf{x}}^{\eta}g(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{m}$
$\leq(\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}+\left\|g\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}){\bm{\lambda}}^{{\bm{\beta}}}{\bm{\lambda}}_{\eta}^{\gamma}\,,$
where ${\bm{\lambda}},\mathbf{x},\psi,\eta$ are chosen as in (3.6). One checks that $\left\|(f+g)(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{m}$ follows a similar bound. Straightforward calculation then gives (3.7). ∎
Remark 3.7.
It is straightforward to see that a similar lemma holds for positive ${\bm{\alpha}},\tilde{\bm{\alpha}}$.
The most simple examples of elements of $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ spaces are Brownian sheet and white noise. A Brownian sheet is a centered Gaussian process $B_{\mathbf{t}}$, $\mathbf{t}\in\mathbb{R}_{+}^{d}$ characterized by the covariance structure
$\mathbb{E}(B_{\mathbf{t}}B_{\mathbf{s}})=\prod_{i=1}^{d}t_{i}\wedge s_{i},\qquad\mathbf{s},\mathbf{t}\in\mathbb{R}^{d}_{+}\,,$
see for example [Kho02, CW75]. It especially fulfills for any $m\geq 1$
$\left\|\square^{\theta}_{\mathbf{s},\mathbf{t}}B\right\|_{m}\leq c(m)\left|\mathbf{t}-\mathbf{s}\right|_{\text{c},\theta}^{\frac{1}{2}{\bm{1}}}.$
for some constant $c$ depending on $m$ and all $\mathbf{s}\leq\mathbf{t},\theta\subset[d]$. Its conditional expectation fulfills for any $\emptyset\neq\eta\subset\theta$
$\mathbb{E}^{\eta}_{\mathbf{s}}\square^{\theta}_{\mathbf{s},\mathbf{t}}B=0\,.$
This implies $B\in C^{\frac{1}{2}{\bm{1}},\infty}L_{m}$.
Its distributional derivative is given by white noise. More generally, we define white noise as the centered Gaussian process $\xi(\psi)$, $\psi\in L_{2}(\mathbb{R}_{+}^{d})$ with covariance structure
[…]
see [Nua95] for reference. Using the Gaussianity of $\xi$, one can calculate
$\left\|\xi(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{L_{m}}=c(m)\left\|\xi(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{L_{2}}=c(m){\bm{\lambda}}^{-\frac{1}{2}{\bm{1}}}\left\|\psi\right\|_{L_{2}(\mathbb{R}^{d})}\,,$
for any $\psi\in L_{2}(\mathbb{R}^{d}),\mathbf{x}\in\mathbb{R}^{d}_{+}$ and ${\bm{\lambda}}\in(0,1]^{d}$, such that $\psi_{\mathbf{x}}^{\bm{\lambda}}$ has support in +$\mathbb{R}^{d}_{+}$ and some universal constants $c(m)$ depending only on $m$. It follows together with
$\mathbb{E}^{\eta}_{\mathbf{x}}\xi(\psi_{\mathbf{x}}^{\bm{\lambda}})=0\,,$
for $\eta\subset[d]$ and $\mbox{supp}(\psi)\subset[0,1]^{d}$ that $\xi\in C^{-\frac{1}{2}{\bm{1}},\infty}L_{m}$.
3.2 Coherent stochastic germs and main statement
We now pass to the main statement of the article. Similar to [CZ20, Ker23], we first introduce the notion of coherence, which is the minimal assumption under which reconstruction is possible. For any ${\bm{\alpha}}<{\bm{0}}$, we choose $r>0$ to be the smallest integer, such that $r+\alpha_{i}>0$ holds for all $i\in[d]$. Throughout this section, we assume $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in[0,T]^{d}}$ to be a commuting filtration.
Definition 3.8 (stochastic coherence).
Let ${\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}\in{\mathbb{R}}^{d}$ with ${\bm{\delta}}\geq{\bm{0}}$ and ${\bm{\gamma}}\geq{\bm{\alpha}}$. An $(\mathcal{F}_{\mathbf{x}})_{\mathbf{x}\in[0,T]^{d}}$-adapted germ $(F_{\mathbf{x}})$ is called stochastic $({\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}})$-coherent, if for all $\eta\subset\theta\subset[d]$, $\psi\in B^{r}([0,1]^{\theta}\times[-1,1]^{\theta^{c}})$ and $\mathbf{x},\mathbf{y}\in[0,T]^{d}$ such that $\mathbf{x}_{\theta}\leq\mathbf{y}_{\theta}$ and $\mbox{supp}(\psi_{\mathbf{x}}^{\bm{\lambda}})\in[0,T]^{d}$, it holds that
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square^{\theta}_{\mathbf{x},\mathbf{y}}F(\psi_{\mathbf{y}}^{{\bm{\lambda}}})\right\|_{m}$ $\lesssim{\bm{\lambda}}^{{\bm{\alpha}}}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\gamma}}-{\bm{\alpha}}}_{\theta}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\delta}}}_{\eta}\,.$ (3.8)
Remark 3.9.
Let us explain why we prefer to consider test functions with support in $[0,1]^{\theta}\times[-1,1]^{\theta^{c}}$. While one can work with test functions with generic support in $[0,T]^{d}$, it is more convenient to work with a normalized support in $[-1,1]^{d}$ and explore more general cases by recentering and rescaling $\psi$ to get $\psi_{\mathbf{x}}^{\bm{\lambda}}$, see for example [CZ20] or [Hai14] for similar settings. Since our theory is stochastic, we make the further distinction of whether a test function only “sees the future” in the direction $i\in[d]$, i.e. has support in $[0,1]$, or has generic support in $[-1,1]$. So the condition $\mbox{supp}(\psi)\subset[0,1]^{\theta}\times[-1,1]^{\theta^{c}}$ allows us to get the correct measurability conditions later on.
Remark 3.10.
As it turns out, this is equivalent to working with test functions $\psi$ with support in $[0,1]^{d}$. Since $\square^{\theta}_{\mathbf{x},\mathbf{y}}F$ does not depend on $\mathbf{y}_{\theta^{c}}$, it holds for any $\mathbf{z}\in\mathbb{R}^{d}$, that
$\square_{\mathbf{x},\mathbf{y}}^{\theta}F(\psi_{\mathbf{y}})=\square_{\mathbf{x},\mathbf{y}+\mathbf{z}_{\theta^{c}}}^{\theta}F(\hat{\psi}_{\mathbf{y}+\mathbf{z}_{\theta^{c}}})\,,$
where $\hat{\psi}=\psi_{-\mathbf{z}_{\theta^{c}}}$. This allows one to switch from test functions with support in $[0,1]^{d}$ to test functions with support in $[0,1]^{\theta}\times[-\frac{1}{2},\frac{1}{2}]^{\theta^{c}}$ at no cost. We prefer to remember that the support of $\psi$ is allowed “to see the past” for indexes $i\in\theta^{c}$, so we explicitly allow supports in $[0,1]^{\theta}\times[-1,1]^{\theta^{c}}$.
We will refer to these elements as stochastic $({\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}})$-coherent germs and we denote them with the set $G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}L_{m}$. We set the norm for this space to be
$\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}L_{m}}:=\sup_{\mathbf{x},\mathbf{y},\theta,\eta,{\bm{\lambda}},\psi}\frac{\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square^{\theta}_{\mathbf{x},\mathbf{y}}F(\psi_{\mathbf{y}}^{{\bm{\lambda}}})\right\|_{m}}{{\bm{\lambda}}^{{\bm{\alpha}}}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\gamma}}_{\theta}-{\bm{\alpha}}_{\theta}+{\bm{\delta}}_{\eta}}_{\theta}}\,,$ (3.9)
where the supremum runs over $\mathbf{x},\mathbf{y},\eta,\theta,{\bm{\lambda}}$ and $\psi$ as in the above definition. Note that this implies for $\eta=\emptyset=\theta$ that $\left\|F_{\mathbf{x}}(\psi_{\mathbf{y}}^{\bm{\lambda}})\right\|_{m}\lesssim{\bm{\lambda}}^{{\bm{\alpha}}}\,,$ giving us that $F_{\mathbf{x}}$ belongs to $C^{{\bm{\alpha}},{\bm{0}}}L_{m}([0,T]^{d})$. This is only reasonable for ${\bm{\alpha}}<{\bm{0}}$, so we will assume from this point onwards that ${\bm{\alpha}}<{\bm{0}}$. Further note that the assumption ${\bm{\gamma}}\geq{\bm{\alpha}}$ comes for free, as we can always lower ${\bm{\alpha}}$ at the cost of a constant and a higher $r$.
Since the right-most point of $\square_{\mathbf{x},\mathbf{y}}^{\theta}$ is not $\mathbf{y}$ but $\pi^{\theta}_{\mathbf{y}}\mathbf{x}$, it would be reasonable to use $\psi_{\pi^{\theta}_{\mathbf{y}}\mathbf{x}}^{\bm{\lambda}}$ as the localized test function. That is, one might want to replace property (3.8) with
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square^{\theta}_{\mathbf{x},\mathbf{y}}F(\psi_{\pi^{\theta}_{\mathbf{y}}\mathbf{x}}^{{\bm{\lambda}}})\right\|_{m}\lesssim{\bm{\lambda}}^{{\bm{\alpha}}}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\gamma}}-{\bm{\alpha}}}_{\theta}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\delta}}}_{\eta}\,.$
The following result shows that this is equivalent to our definition:
Lemma 3.11.
The following conditions are equivalent:
  1. i)
    $F$ is a stochastic $({\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}})$-coherent germ.
  2. ii)
    $F$ fulfills for each $\eta\subset\theta\subset[d]$ and $\mathbf{x},\mathbf{y}$ with $\mathbf{x}_{\theta}\leq\mathbf{y}_{\theta}$
    $\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square_{\mathbf{x},\mathbf{y}}^{\theta}F(\psi_{\pi^{\theta}_{\mathbf{y}}\mathbf{x}}^{{\bm{\lambda}}})\right\|_{m}\lesssim{\bm{\lambda}}^{{\bm{\alpha}}}(\left|x-y\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\gamma}}-{\bm{\alpha}}}_{\theta}(\left|x-y\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\delta}}}_{\eta}\,.$
  3. iii)
    For all $\mathbf{k}=\mathbf{k}_{\theta^{c}}\in[0,T]^{d}$ and for $\eta,\theta,\mathbf{x},\mathbf{y},\psi,{\bm{\lambda}}$ as in the coherence property, we have
    $\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square_{\mathbf{x},\mathbf{y}}^{\theta}F(\psi_{\mathbf{y}+\mathbf{k}_{\theta^{c}}}^{{\bm{\lambda}}})\right\|_{m}\lesssim{\bm{\lambda}}^{{\bm{\alpha}}}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\gamma}}-{\bm{\alpha}}}_{\theta}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\delta}}}_{\eta}\,,$
    where the constant in $\lesssim$ does not depend on $\mathbf{k}$.
Proof.
$i)$ and $iii)$ are obviously equivalent, since $\square^{\theta}_{x,y}=\square^{\theta}_{x,y+{k_{\theta}^{c}}}$. $i)\Rightarrow ii)$ follows from the simple identity $\square_{\mathbf{x},\mathbf{y}}^{\theta}=\square_{\mathbf{x},\pi_{\mathbf{y}}^{\theta}\mathbf{x}}^{\theta}$. On the other hand, $ii)\Rightarrow i)$ is shown by using Lemma 2.2 to calculate
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square_{\mathbf{x},\mathbf{y}}^{\theta}F(\psi_{\mathbf{y}}^{{\bm{\lambda}}})\right\|_{m}$ $\leq\sum_{\theta^{c}=\theta_{1}\sqcup\theta_{2}}\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square^{\theta\cup\theta_{2}}_{\pi_{\mathbf{y}}^{\theta_{1}}\mathbf{x},\mathbf{y}}F(\psi_{\mathbf{y}}^{{\bm{\lambda}}})\right\|_{m}\lesssim{\bm{\lambda}}^{{\bm{\alpha}}}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\gamma}}-{\bm{\alpha}}}_{\theta}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\delta}}}_{\eta}\,.$
Using these notions, we can now state precisely the stochastic reconstruction theorem.
Theorem 3.12 (Stochastic reconstruction theorem on rectangular increments).
Let $(F_{\mathbf{x}})$ be a stochastic $({\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}})$-coherent germ for ${\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}\in\mathbb{R}^{d},{\bm{\delta}}\geq 0,{\bm{\gamma}}\geq{\bm{\alpha}},{\bm{\alpha}}<{\bm{0}}$ and let $m\geq 2$. Assume that ${\bm{\gamma}}+{\bm{\delta}}>{\bm{0}}$, and that ${\bm{\gamma}}>-\frac{1}{2}{\bm{1}}$. Then there exists a unique family of random distributions $\mathcal{R}_{\mathbf{x}}^{\eta}(F)\in\mathcal{D}^{rd}((0,T)^{d})$ with $\mathbf{x}\in[0,T]^{d}$ and $\eta\subset[d]$ characterized by the following properties:
  1. 1.
    $\mathcal{R}^{\emptyset}_{\mathbf{x}}(F)=F_{\mathbf{x}}$.
  2. 2.
    $\mathcal{R}^{\theta}_{\mathbf{x}}(F)$ only depends on $x_{i}$ for $i\in\theta^{c}$, i.e. $\mathcal{R}^{\theta}_{\mathbf{x}}(F)=\mathcal{R}^{\theta}_{\pi_{\bm{0}}^{\theta}\mathbf{x}}(F)$.
  3. 3.
    For all $\theta\subset[d]$, $\mathbf{x}\in[0,T]^{d}$ and $\psi$ with support in $[0,y_{1}]\times\dots\times[0,y_{d}]$, $\mathcal{R}^{\theta}_{\mathbf{x}}(F)(\psi)$ is $\mathcal{F}_{\mathbf{x}\vee\mathbf{y}}$-measurable.
  4. 4.
    For all $\psi$ with support in $[1/4,3/4]^{\theta}\times[-3/4,3/4]^{\theta^{c}}$ and $\left\|\psi\right\|_{C_{c}^{rd}}=1$, any subset $\eta\subset\theta\neq\emptyset$ one has the estimate
    $\left\|E^{\eta}_{\mathbf{x}}\sum_{\hat{\theta}\subset\theta}(-1)^{\sharp{\hat{\theta}}}\mathcal{R}^{\hat{\theta}}_{\mathbf{x}}(F)(\psi^{{\bm{\lambda}}}_{\mathbf{z}})\right\|_{m}\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}L_{m}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\theta}{\bm{\lambda}}^{\bm{\delta}}_{\eta}{\bm{\lambda}}^{{\bm{\alpha}}}_{\theta^{c}}\,,$ (3.10)
    uniformly over $\mathbf{z}\in[0,T]^{d}$, ${\bm{\lambda}}\in(0,1]^{d}$ such that $[\mathbf{z}-{\bm{\lambda}}_{\eta^{c}},\mathbf{z}+{\bm{\lambda}}]\subset[0,T]^{d}$ and $\mathbf{x}\in[0,T]^{d}$ with $\mathbf{x}_{\theta}=\mathbf{z}_{\theta}$.
We call $\mathcal{R}^{[d]}_{\mathbf{x}}(F)$ the multiparameter reconstruction of the stochastic germ $F$ which we denote by $\mathcal{R}(F)$.
Remark 3.13.
Note that we switch to test functions with $rd$ derivatives, as we use Lemma 2.8 in the proof of the reconstruction theorem.
Remark 3.14.
While the coherence property works for test functions with support in $[0,1]^{\theta}\times[-1,1]^{\theta^{c}}$, we have to reduce this set by some fixed distance in (3.10). This is because for any test function $\psi$ with support in some compact set $K$, its wavelet approximation
$\sum_{\mathbf{x}\in\Delta_{\mathbf{n}}}\phi^{\mathbf{n}}_{\mathbf{x}}\left\langle\phi^{\mathbf{n}}_{\mathbf{x}},\psi\right\rangle$
has support in the $\epsilon_{\mathbf{n}}$-fattening of $K$ given by
$\{\mathbf{y}\in\mathbb{R}^{d}\colon\text{dist}(\mathbf{y},K)\leq\epsilon_{\mathbf{n}}\}\,.$
for some $\epsilon_{\mathbf{n}}$ which approaches 00 as $\mathbf{n}\to\infty$. Thus, by choosing $\psi$ with a slightly smaller support, we can apply the coherence property to its wavelet approximation for sufficiently large $\mathbf{n}$, see also Remark 3.5.
3.3 Proof of the stochastic reconstruction theorem
We now pass to the proof of Theorem 3.12. To achieve that we recall a preliminary result whose proof can be found in [BHK23, Lemma 19].
Lemma 3.15.
Let $(h_{\mathbf{n}})_{\mathbf{n}\in\mathbb{N}^{d}}$ be a multiparameter sequence in a Banach space $\mathcal{H}$. We define the operators $I^{i}h_{\mathbf{n}}=h_{\mathbf{n}+{\bm{1}}_{\{i\}}}$. Assume that there is a constant $C>0$ and ${\bm{\alpha}}=(\alpha_{1},\dots,\alpha_{d})>{\bm{0}}$ such that for all $\mathbf{n}\in\mathbb{N}^{d}$ and $\theta\subset[d]$
$\left\|\prod_{i\in\theta}(\mbox{Id}-I^{i})h_{\mathbf{n}_{\theta}}\right\|\leq C2^{-{\bm{\alpha}}*\mathbf{n}}_{\theta}\,.$
Then for all $\mathbf{n}\leq\mathbf{k}$ and $\theta\subset[d]$ we have
$\left\|h_{\mathbf{n}_{\theta}}-h_{\mathbf{k}_{\theta}}\right\|\lesssim C\sum_{i\in\theta}2^{-n_{i}\alpha_{i}}\,.$
It especially follows that $(h_{\mathbf{n}})$ is a Cauchy sequence.
We fix wavelets $\phi,\hat{\phi}^{\chi},\chi\subset[d]$ as described in Section 2.2. Knowing the above result, to construct the family of distributions $\mathcal{R}_{\mathbf{x}}^{\eta}(F)$ in Theorem 3.12, it is natural to consider the following multiparameter approximations
$\mathcal{R}^{\eta,\mathbf{n}}_{\mathbf{x}}(\psi):=\sum_{\mathbf{y}\in\Delta_{\mathbf{n}}}F_{\pi^{\eta}_{\mathbf{y}}\mathbf{x}}(\phi^{\mathbf{n}}_{\mathbf{y}})\left\langle\phi^{\mathbf{n}}_{\mathbf{y}},\psi\right\rangle\,.$
(we will drop the dependency on $F$ for the sake of clarity) and to show proper bounds of the quantity
$\left\|\prod_{i\in\theta}(\mbox{Id}-I^{i})\mathcal{R}^{\eta,\mathbf{n}}_{\mathbf{x}}(\psi)\right\|_{m}\,,$ (3.11)
where $I^{i}\mathcal{R}^{\eta,\mathbf{n}}_{\mathbf{x}}=\mathcal{R}^{\eta,\mathbf{n}+{\bm{1}}_{\{i\}}}_{\mathbf{x}}$. Once we have done so, we can conclude that $\mathcal{R}^{\eta,\mathbf{n}}_{\mathbf{x}}$ converges to some limit. Before we show the inequality, let us introduce some more notation: For any $\kappa\subset[d]$, we set
$g^{\kappa,\mathbf{n}}_{\mathbf{x}}(\psi):=\sum_{\eta\subset\kappa}(-1)^{\sharp\eta}\mathcal{R}^{\eta,\mathbf{n}}_{\mathbf{x}}(\psi)=(-1)^{\sharp\kappa}\sum_{\mathbf{y}\in\Delta_{\mathbf{n}}}\square^{\kappa}_{\mathbf{x},\mathbf{y}}F(\phi_{\mathbf{y}}^{\mathbf{n}})\left\langle\phi_{\mathbf{y}}^{\mathbf{n}},\psi\right\rangle\,,$
and set $I^{i}g^{\kappa,\mathbf{n}}_{\mathbf{x}}:=g^{\kappa,\mathbf{n}+{\bm{1}}_{\{i\}}}_{\mathbf{x}}$, as before. Further, recall that the one-dimensional wavelets $\varphi,\hat{\varphi}$ fulfill the property
$\varphi_{z}^{n}=\sum_{k\in\Delta_{n+1}}a_{k}^{n}\varphi^{n+1}_{z+k}\,.$
Comparing this to (2.8) one has $a_{k}^{n}=a_{2^{n}k}$. We similarly introduce the notation $b^{n}_{k}:=\left\langle\varphi_{z+k}^{n+1},\hat{\varphi}_{z}^{n}\right\rangle$. Extending the properties of $1$ dimensional wavelets to $d$-dimensional wavelets one has
$\left\langle\phi^{\mathbf{n}+{\bm{1}}_{\chi}+{\bm{1}}_{\theta}}_{\mathbf{z}+\mathbf{k}},\hat{\phi}^{\chi,\mathbf{n}}_{\mathbf{z}}\right\rangle=a_{\mathbf{k},\theta}^{\mathbf{n}}b_{\mathbf{k},\chi}^{\mathbf{n}}\delta_{{\bm{0}},\mathbf{k}}^{(\chi\cup\theta)^{c}}\,,$ (3.12)
for all $\chi,\theta\subset[d]$ such that $\chi\cap\theta=\emptyset$, $\mathbf{z}\in\Delta_{\mathbf{n}},\mathbf{k}\in\Delta_{\mathbf{n}+{\bm{1}}_{\chi}+{\bm{1}}_{\theta}}$, where we use the notations
$a_{\mathbf{k},\theta}^{\mathbf{n}}:=\prod_{i\in\theta}a_{k_{i}}^{n_{i}}\,,\quad b_{\mathbf{k},\chi}^{\mathbf{n}}:=\prod_{i\in\chi}b_{k_{i}}^{n_{i}}\,,\quad\delta_{{\bm{0}},\mathbf{k}}^{(\chi\cup\theta)^{c}}=\prod_{i\in(\chi\cup\theta)^{c}}\delta_{0,k_{i}}\,,$
where $\delta_{0,k_{i}}=1$ if $k_{i}=0$ and 00, otherwise. Instead of (3.11), we will instead show a bound on
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\prod_{i\in\theta}(\mbox{Id}-I^{i})g^{\kappa,\mathbf{n}}_{\mathbf{x}}(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{m}\,.$
which has the advantage that it can also be used to show (3.10). To get there, we split the underlying test function $\psi$ according to the functions $\hat{\phi}^{\chi}$
$\prod_{i\in\theta}(\mbox{Id}-I^{i})g^{\kappa,\mathbf{n}}_{\mathbf{x}}(\psi)=\sum_{\chi\subset\theta}\prod_{i\in\theta}(\mbox{Id}-I^{i})g^{\kappa,\mathbf{n}}_{\mathbf{x}}(\hat{P}^{\chi,\mathbf{n}}\psi)\,,$ (3.13)
where we used that $g^{\kappa,\mathbf{n}}_{\mathbf{x}}$ belongs to $V^{\mathbf{n}}$. A first lemma bounds the right-hand side of (3.13).
Lemma 3.16.
Under the assumptions of the reconstruction theorem, it holds that for all $\mathbf{n}\in\mathbb{N}^{d}$, $\chi\subset\theta\subset[d]$, $\mathbf{x},\mathbf{z}\in\Delta_{\mathbf{n}}\cap[0,T]^{d}$, such that $\mathbf{x}_{\kappa\setminus(\theta\setminus\chi)}\leq\mathbf{z}_{\kappa\setminus(\theta\setminus\chi)}$ one has
$\left\|\mathbb{E}^{\eta}_{\pi_{\mathbf{z}}^{\theta\setminus\chi}\mathbf{x}}\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}}(\hat{\phi}^{\chi,\mathbf{n}}_{\mathbf{z}})\right\|_{m}\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}2^{-\mathbf{n}*(\frac{1}{2}{\bm{1}}+{\bm{\alpha}})}$ $(\pi_{0}^{\theta\setminus\chi}\left|\mathbf{x}-\mathbf{z}\right|_{\text{c}}+2^{-\mathbf{n}})^{({\bm{\gamma}}-{\bm{\alpha}})}_{\kappa}$
$\times$ $(\pi_{0}^{\theta\setminus\chi}\left|\mathbf{x}-\mathbf{z}\right|_{\text{c}}+2^{-\mathbf{n}})^{{\bm{\delta}}}_{\eta}\,,$
where we assume $\mbox{supp}(\hat{\phi}_{\mathbf{z}}^{\chi,\mathbf{n}})\subset[0,T]^{d}$ and $\eta\subset\kappa\subset[d]$.
Proof.
By definition of $g^{\kappa,\mathbf{n}}_{\mathbf{x}}(\psi)$ we have the identity
$\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}}(\hat{\phi}^{\chi,\mathbf{n}}_{\mathbf{z}})=\sum_{\eta\subset\kappa}\prod_{i\in\theta}(\mbox{Id}-I^{i})(-1)^{\sharp\eta}\mathcal{R}^{\eta,\mathbf{n}}_{\mathbf{x}}(\hat{\phi}^{\chi,\mathbf{n}}_{\mathbf{z}})\,.$ (3.14)
By construction of our wavelets, we have $\left\langle\phi^{\mathbf{m}}_{\mathbf{x}},\hat{\phi}_{\mathbf{y}}^{\chi,\mathbf{n}}\right\rangle=0$ if there exists an $i\in\chi$ such that $m_{i}\leq n_{i}$, especially implying $\mathcal{R}_{\mathbf{x}}^{\kappa,\mathbf{m}}(\hat{\phi}_{\mathbf{y}}^{\chi,\mathbf{n}})=0$ for such $\chi,\mathbf{n},\mathbf{m}$. Multiplying out the product over $i\in\theta\cap\chi$ and eliminating all zero terms therefore gives us the identity
$\prod_{i\in\theta}(\mbox{Id}-I^{i})\mathcal{R}_{\mathbf{x}}^{\kappa,\mathbf{n}}(\hat{\phi}_{\mathbf{z}}^{\chi,\mathbf{n}})=(-1)^{\sharp\chi}\prod_{i\in\theta\setminus\chi}(\mbox{Id}-I^{i})\mathcal{R}^{\kappa,\mathbf{n}+{\bm{1}}_{\chi}}_{\mathbf{x}}(\hat{\phi}^{\chi,\mathbf{n}}_{\mathbf{z}})\,.$
Note that in the one-dimensional case, one has
$\sum_{y\in\Delta_{n}}F_{x}(\varphi^{n}_{y})\left\langle\varphi^{n}_{y},\varphi^{n}_{z}\right\rangle=F_{x}(\varphi^{n}_{z})=\sum_{y\in\Delta_{n+1}}F_{x}(\varphi^{n+1}_{y})\left\langle\varphi^{n+1}_{y},\varphi^{n}_{z}\right\rangle\,,$
implying in $d\geq 1$ that for all $i\in\kappa^{c}$, we have
$(\mbox{Id}-I^{i})\mathcal{R}^{\kappa,\mathbf{n}+{\bm{1}}_{\chi}}_{\mathbf{x}}(\hat{\phi}_{\mathbf{z}}^{\chi,\mathbf{n}})=0\,.$
Thus, we conclude that
$\prod_{i\in\theta\setminus\chi}(\mbox{Id}-I^{i})\mathcal{R}^{\kappa,\mathbf{n}+{\bm{1}}_{\chi}}_{\mathbf{x}}(\hat{\phi}_{\mathbf{z}}^{\chi,\mathbf{n}})=0$ (3.15)
unless $\theta\setminus\chi\subset\kappa$. Therefore we can rewrite the right-hand side of (3.14) into
$\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}}(\hat{\phi}^{\chi,\mathbf{n}}_{\mathbf{z}})=\sum_{\theta\setminus\chi\subset\tilde{\kappa}\subset\kappa}\leavevmode\nobreak\ \prod_{i\in\theta\setminus\chi}(\mbox{Id}-I^{i})(-1)^{\sharp{\tilde{\kappa}}+\sharp\chi}\mathcal{R}_{\mathbf{x}}^{\tilde{\kappa},\mathbf{n}+{\bm{1}}_{\chi}}(\hat{\phi}_{\mathbf{z}}^{\chi,\mathbf{n}})\,.$ (3.16)
We further analyze the product on the right-hand side by expanding $\mathcal{R}^{\tilde{\kappa},\mathbf{n}+{\bm{1}}_{\chi}}_{\mathbf{x}}$.
$\prod_{i\in\theta\setminus\chi}(\mbox{Id}-I^{i})\mathcal{R}_{\mathbf{x}}^{\tilde{\kappa},\mathbf{n}+{\bm{1}}_{\chi}}(\hat{\phi}_{\mathbf{z}}^{\chi,\mathbf{n}})$ $=\sum_{\tilde{\theta}\subset\theta\setminus\chi}(-1)^{\sharp{\tilde{\theta}}}\mathcal{R}^{\tilde{\kappa},\mathbf{n}+{\bm{1}}_{\tilde{\theta}}+{\bm{1}}_{\chi}}_{\mathbf{x}}(\hat{\phi}_{\mathbf{z}}^{\mathbf{n},\chi})$
$=\sum_{\tilde{\theta}\subset\theta\setminus\chi}\sum_{\mathbf{y}\in\Delta_{\mathbf{n}+{\bm{1}}_{\chi}+{\bm{1}}_{\tilde{\theta}}}}(-1)^{\sharp\tilde{\theta}}F_{\pi^{\tilde{\kappa}}_{\mathbf{z}+\mathbf{y}}\mathbf{x}}(\phi_{\mathbf{z}+\mathbf{y}}^{\mathbf{n}+{\bm{1}}_{\chi}+{\bm{1}}_{\tilde{\theta}}})\left\langle\phi^{\mathbf{n}+{\bm{1}}_{\chi}+{\bm{1}}_{\tilde{\theta}}}_{\mathbf{z}+\mathbf{y}},\hat{\phi}_{\mathbf{z}}^{\mathbf{n},\chi}\right\rangle\,.$
Using the identity (3.12) together with the expansion
$\phi_{\mathbf{z}+\mathbf{y}}^{\mathbf{n}+{\bm{1}}_{\chi}+{\bm{1}}_{\tilde{\theta}}}=\sum_{\mathbf{k}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\theta\setminus(\chi\cup\tilde{\theta})}}a^{\mathbf{n}}_{\mathbf{k},\theta\setminus(\chi\cup\tilde{\theta})}\phi_{\mathbf{z}+\mathbf{y}+\mathbf{k}}^{\mathbf{n}+{\bm{1}}_{\theta}}\,,$
which itself follows from (2.8), we get
$\prod_{i\in\theta\setminus\chi}(\mbox{Id}-I^{i})\mathcal{R}_{\mathbf{x}}^{\tilde{\kappa},\mathbf{n}+{\bm{1}}_{\chi}}(\hat{\phi}_{\mathbf{z}}^{\chi,\mathbf{n}})=\sum_{\tilde{\theta}\subset\theta\setminus\chi}\sum_{\mathbf{y}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\chi\cup\tilde{\theta}}}\sum_{\mathbf{k}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\theta\setminus(\chi\cup\tilde{\theta})}}(-1)^{\sharp\tilde{\theta}}a_{\mathbf{y}+\mathbf{k},\theta\setminus\chi}^{\mathbf{n}}b_{\mathbf{y},\chi}^{\mathbf{n}}F_{\pi^{\tilde{\kappa}}_{\mathbf{z}+\mathbf{y}}\mathbf{x}}(\phi_{\mathbf{z}+\mathbf{y}+\mathbf{k}}^{\mathbf{n}+{\bm{1}}_{\theta}})\,.$
Let $\mathbf{u}:=\pi_{\mathbf{y}+\mathbf{k}}^{\theta\setminus\chi}{\bm{0}},\mathbf{v}:=\pi_{\mathbf{y}}^{\chi}{\bm{0}}$, such that $\mathbf{y}+\mathbf{k}=\mathbf{u}+\mathbf{v}$ holds. Using these notations allows us to pull the sum over $\tilde{\theta}\subset\theta\setminus\chi$ into the other two sums, which leads to
$\prod_{i\in\theta\setminus\chi}(\mbox{Id}-I^{i})\mathcal{R}_{\mathbf{x}}^{\tilde{\kappa},\mathbf{n}+{\bm{1}}_{\chi}}(\hat{\phi}_{\mathbf{z}}^{\chi,\mathbf{n}})$ $=\sum_{\mathbf{u}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\theta\setminus\chi}}\sum_{\mathbf{v}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\chi}}a_{\mathbf{u},\theta\setminus\chi}^{\mathbf{n}}b_{\mathbf{v},\chi}^{\mathbf{n}}\sum_{\tilde{\theta}\subset\theta\setminus\chi}(-1)^{\sharp\tilde{\theta}}F_{\pi^{\tilde{\kappa}}_{\pi^{\tilde{\theta}}_{\mathbf{z}+\mathbf{u}+\mathbf{v}}(\mathbf{z}+\mathbf{v})}\mathbf{x}}(\phi_{\mathbf{z}+\mathbf{u}+\mathbf{v}}^{\mathbf{n}+{\bm{1}}_{\theta}})$
$=\sum_{\mathbf{u}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\theta\setminus\chi}}\sum_{\mathbf{v}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\chi}}a_{\mathbf{u},\theta\setminus\chi}^{\mathbf{n}}b_{\mathbf{v},\chi}^{\mathbf{n}}(-1)^{\sharp(\theta\setminus\chi)}\square_{\mathbf{z}+\mathbf{v},\mathbf{z}+\mathbf{v}+\mathbf{u}}^{\theta\setminus\chi}F_{\pi^{\tilde{\kappa}}_{\cdot}\mathbf{x}}(\phi_{\mathbf{z}+\mathbf{u}+\mathbf{v}}^{\mathbf{n}+{\bm{1}}_{\theta}})$
$=\sum_{\mathbf{u}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\theta\setminus\chi}}\sum_{\mathbf{v}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\chi}}a_{\mathbf{u},\theta\setminus\chi}^{\mathbf{n}}b_{\mathbf{v},\chi}^{\mathbf{n}}(-1)^{\sharp(\theta\setminus\chi)}\square_{\pi^{\tilde{\kappa}}_{\mathbf{z}+\mathbf{v}}\mathbf{x},\mathbf{z}+\mathbf{v}+\mathbf{u}}^{\theta\setminus\chi}F(\phi_{\mathbf{z}+\mathbf{u}+\mathbf{v}}^{\mathbf{n}+{\bm{1}}_{\theta}})$
where we use the fact that $\theta\setminus\chi\subset\tilde{\kappa}$ holds. By summing the last part over $(-1)^{\sharp\tilde{\kappa}}$ one has
$\sum_{\theta\setminus\chi\subset\tilde{\kappa}\subset\kappa}(-1)^{\sharp\tilde{\kappa}}\square_{\pi^{\tilde{\kappa}}_{\mathbf{z}+\mathbf{v}}\mathbf{x},\mathbf{z}+\mathbf{v}+\mathbf{u}}^{\theta\setminus\chi}F$ $=\sum_{\hat{\kappa}\subset\kappa\setminus(\theta\setminus\chi)}(-1)^{\sharp\hat{\kappa}+\sharp(\theta\setminus\chi)}\square^{\theta\setminus\chi}_{\pi^{\hat{\kappa}}_{\mathbf{z}+\mathbf{v}}\pi_{\mathbf{z}}^{\theta\setminus\chi}\mathbf{x},\mathbf{z}+\mathbf{v}+\mathbf{u}}F$
$=(-1)^{\sharp\kappa}\square^{\kappa\setminus(\theta\setminus\chi)}_{\pi^{\theta\setminus\chi}_{\mathbf{z}}\mathbf{x},\mathbf{z}+\mathbf{v}+\mathbf{u}}\square^{\theta\setminus\chi}_{\cdot,\mathbf{z}+\mathbf{u}+\mathbf{v}}F\,,$
where we use the fact that $u_{i}=0$ for all $i\in\kappa\setminus(\theta\setminus\chi)$ and $v_{i}=0$ for all $i\in\theta\setminus\chi$. Using Lemma 2.1, it follows that
$\sum_{\theta\setminus\chi\subset\tilde{\kappa}\subset\kappa}(-1)^{\sharp\tilde{\kappa}}\square_{\pi^{\tilde{\kappa}}_{\mathbf{z}+\mathbf{v}}\mathbf{x},\mathbf{z}+\mathbf{v}+\mathbf{u}}^{\theta\setminus\chi}F=(-1)^{\sharp\kappa}\square^{\kappa}_{\pi^{\theta\setminus\chi}_{\mathbf{z}}\mathbf{x},\mathbf{z}+\mathbf{v}+\mathbf{u}}F.$
Putting this into (3.16) one obtains
$\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}}(\hat{\phi}^{\chi,\mathbf{n}}_{\mathbf{z}})=(-1)^{\sharp\theta+\sharp\kappa}\sum_{\mathbf{u}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\theta\setminus\chi}}\sum_{\mathbf{v}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\chi}}a_{\mathbf{u},\theta\setminus\chi}^{\mathbf{n}}b_{\mathbf{v},\chi}^{\mathbf{n}}\square^{\kappa}_{\pi^{\theta\setminus\chi}_{\mathbf{z}}\mathbf{x},\mathbf{z}+\mathbf{v}+\mathbf{u}}F(\phi^{\mathbf{n}+{\bm{1}}_{\theta}}_{\mathbf{z}+\mathbf{u}+\mathbf{v}})\,.$
If $\mathbf{x}_{\kappa\setminus(\theta\setminus\chi)}\leq\mathbf{z}_{\kappa\setminus(\theta\setminus\chi)}$ holds, we get by the coherence property for $\eta\subset\kappa$:
$\left\|\mathbb{E}^{\eta}_{\pi^{\theta\setminus\chi}_{\mathbf{z}}\mathbf{x}}\square^{\kappa}_{\pi^{\theta\setminus\chi}_{\mathbf{z}}\mathbf{x},\mathbf{z}+\mathbf{v}+\mathbf{u}}F(\phi^{\mathbf{n}+\mathbf{1}_{\theta}}_{\mathbf{z}+\mathbf{u}+\mathbf{v}})\right\|_{m}\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}2^{-\mathbf{n}*{\bm{\alpha}}-\frac{\mathbf{n}}{2}}$ $(\pi_{{\bm{0}}}^{\theta\setminus\chi}\left|\mathbf{x}-\mathbf{z}\right|_{\text{c}}+2^{-\mathbf{n}})^{{\bm{\gamma}}-{\bm{\alpha}}}_{\kappa}$
$\times$ $(\pi_{{\bm{0}}}^{\theta\setminus\chi}\left|\mathbf{x}-\mathbf{z}\right|_{\text{c}}+2^{-\mathbf{n}})^{{\bm{\delta}}}_{\eta}\,,$
where we use the property that $a_{\mathbf{u},\theta\setminus\chi}^{\mathbf{n}}b_{\mathbf{v},\chi}^{\mathbf{n}}\neq 0$ implies that $\left|u_{i}+v_{i}\right|\lesssim 2^{-n_{i}}$ for all $i\in\theta$. The claim then follows from the fact that the sum
$\sum_{\mathbf{u}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\theta\setminus\chi}}\sum_{\mathbf{v}\in\Delta_{\mathbf{n}+{\bm{1}}_{\theta}}^{\chi}}a_{\mathbf{u},\theta\setminus\chi}^{\mathbf{n}}b_{\mathbf{v},\chi}^{\mathbf{n}}=\sum_{\mathbf{u}\in\Delta_{\theta}^{\theta\setminus\chi}}\sum_{\mathbf{v}\in\Delta_{\theta}^{\chi}}a_{\mathbf{u},\theta\setminus\chi}^{{\bm{0}}}b_{\mathbf{v},\chi}^{{\bm{0}}}$
is independent of $\mathbf{n}$. ∎
The next step is to test $\prod_{i\in\theta}(\mbox{Id}-I^{i})g^{\kappa,\mathbf{n}}_{\mathbf{x}}$ against a general test function $\psi_{\mathbf{z}}^{\bm{\lambda}}$. To compare ${\bm{\lambda}}$ with the dyadic scale we will write ${\bm{\lambda}}\approx 2^{-\mathbf{n}}$ if there exists a constant $\mathbf{c}>{\bm{0}}$, such that $2^{-n_{i}-1}\leq c_{i}\lambda_{i}\leq 2^{-n_{i}}$ for all $i\in[d]$. Whenever such a constant $\mathbf{c}>{\bm{0}}$ is chosen, we allow all constants in $\lesssim$ to depend on $\mathbf{c}$.
Lemma 3.17.
Let ${\bm{0}}<{\bm{\lambda}}<{\bm{1}}$, $\eta\subset\kappa\subset[d]$ and $\theta\subset[d]$. Choose $\mathbf{n}$ such that ${\bm{\lambda}}\approx 2^{\mathbf{n}}$ and $\mathbf{z},\mathbf{x}\in[0,T]^{d}$, such that $z_{i}=x_{i}$ for all $i\in\kappa$. Finally, choose a test function $\psi\in C^{rd}_{c}([1/4,3/4]^{\kappa}\times[-3/4,3/4]^{\kappa^{c}})$, and assume that $[\mathbf{z},\mathbf{z}+{\bm{\lambda}}]^{\kappa}\times[\mathbf{z}-{\bm{\lambda}},\mathbf{z}+{\bm{\lambda}}]^{\kappa^{c}}\subset[0,T]^{d}$. Then for any $\mathbf{k}\geq{\bm{0}}$ one has
$\left\|\mathbb{E}_{\mathbf{x}}^{\eta}\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}+\mathbf{k}}(\psi_{\mathbf{z}}^{{\bm{\lambda}}})\right\|_{m}\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}\left\|\psi\right\|_{C_{c}^{rd}}{\bm{\lambda}}^{{\bm{\alpha}}}_{\kappa^{c}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\kappa}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}\cdot 2^{-\mathbf{k}*{\bm{\alpha}}}_{\theta^{c}}2^{-\mathbf{k}*{\bm{\xi}}}_{\theta}\,,$
where ${\bm{\xi}}=({\bm{\alpha}}+r{\bm{1}})\wedge({\bm{\gamma}}+{\bm{\delta}})\wedge({\bm{\gamma}}+\frac{1}{2}{\bm{1}})>{\bm{0}}$.
Proof.
Using again the expressions (3.13), and (3.15) we have
$\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}+\mathbf{k}}(\psi_{\mathbf{z}}^{\bm{\lambda}})$ $=\sum_{\begin{subarray}{c}\chi\subset\theta\\ \theta\setminus\chi\subset\kappa\end{subarray}}\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}+\mathbf{k}}(\hat{P}_{\mathbf{n}+\mathbf{k}}^{\chi}\psi_{\mathbf{z}}^{\bm{\lambda}})$
$=\sum_{\begin{subarray}{c}\chi\subset\theta\\ \theta\setminus\chi\subset\kappa\end{subarray}}\sum_{\mathbf{y}\in\Delta_{\mathbf{n}+\mathbf{k}}}\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}+\mathbf{k}}(\hat{\phi}_{\mathbf{y}}^{\chi,\mathbf{n}+\mathbf{k}})\left\langle\hat{\phi}^{\chi,\mathbf{n}+\mathbf{k}}_{\mathbf{y}},\psi_{\mathbf{z}}^{\bm{\lambda}}\right\rangle\,.$
Let us fix an indexset $\chi$ and set $\tilde{\theta}:=\theta\setminus\chi$, $\tilde{\eta}:=\eta\cap\tilde{\theta}$. Note that the scalar product is only non-zero if $\left|y_{i}-z_{i}\right|\lesssim\lambda_{i}$ for all $i\in[d]$, which holds true for at most $\approx 2^{k_{i}}$ many $\mathbf{y}\in\Delta_{\mathbf{n}+\mathbf{k}}^{\{i\}}$ for each $i\in[d]$. For each $\mathbf{y}\in\Delta_{\mathbf{n}+\mathbf{k}}$, we we write $\mathbf{y}=\mathbf{y}_{\tilde{\theta}}+\mathbf{y}_{\tilde{\theta}^{c}}$ and we further split the sum over $\mathbf{y}\in\Delta_{\mathbf{n}+\mathbf{k}}$ to obtain
$\sum_{\begin{subarray}{c}\chi\subset\theta\\ \theta\setminus\chi\subset\kappa\end{subarray}}\sum_{\mathbf{y}_{\tilde{\theta}}\in\Delta_{\mathbf{n}+\mathbf{k}}^{\tilde{\theta}}}\sum_{\mathbf{y}_{\tilde{\theta}^{c}}\in\Delta_{\mathbf{n}+\mathbf{k}}^{\tilde{\theta}^{c}}}\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}+\mathbf{k}}(\hat{\phi}_{\mathbf{y}}^{\chi,\mathbf{n}+\mathbf{k}})\left\langle\hat{\phi}^{\chi,\mathbf{n}+\mathbf{k}}_{\mathbf{y}},\psi_{\mathbf{z}}^{\bm{\lambda}}\right\rangle\,$
Let us assume for the moment that we can apply the BDG-type inequality (2.13) to the quantity
$h(\mathbf{y}_{\tilde{\theta}}):=\sum_{\mathbf{y}_{\tilde{\theta}^{c}}\in\Delta^{\tilde{\theta}^{c}}_{\mathbf{n}+\mathbf{k}}}\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}+\mathbf{k}}(\hat{\phi}^{\chi,\mathbf{n}+\mathbf{k}}_{\mathbf{y}})\left\langle\hat{\phi}^{\chi,\mathbf{n}+\mathbf{k}}_{\mathbf{y}},\psi_{\mathbf{z}}^{{\bm{\lambda}}}\right\rangle\,,$
to estimate
$\left\|\sum_{\mathbf{y}_{\tilde{\theta}}\in\Delta_{\mathbf{n}+\mathbf{k}}^{\tilde{\theta}}}h(\mathbf{y}_{\tilde{\theta}})\right\|_{m}\,.$
Combining Lemma 3.16 and Lemma 2.8 one has
$\left\|\mathbb{E}_{\pi^{\tilde{\theta}}_{\mathbf{y}}\mathbf{x}}^{\eta}h(\mathbf{y}_{\tilde{\theta}})\right\|_{m}\lesssim$
$\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}\left\|\psi\right\|_{C_{c}^{rd}}2^{\mathbf{k}}_{{\tilde{\theta}^{c}}}2^{-(\mathbf{n}+\mathbf{k})*({\bm{1}}+{\bm{\alpha}}+r{\bm{1}}_{\chi})}{\bm{\lambda}}^{-({\bm{1}}+r{\bm{1}}_{\chi})}(\pi_{\bm{0}}^{\tilde{\theta}}{\bm{\lambda}}+2^{-(\mathbf{n}+\mathbf{k})})^{({\bm{\gamma}}-{\bm{\alpha}})}_{\kappa}(\pi_{\bm{0}}^{\tilde{\theta}}{\bm{\lambda}}+2^{-(\mathbf{n}+\mathbf{k})})^{{\bm{\delta}}}_{\eta}$
$\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}\left\|\psi\right\|_{C_{c}^{rd}}{\bm{\lambda}}^{{\bm{\alpha}}}_{{\kappa^{c}}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\kappa}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}\cdot 2^{-\mathbf{k}*({\bm{1}}_{\tilde{\theta}}+{\bm{\alpha}}_{\tilde{\theta}^{c}}+r{\bm{1}}_{\chi}+{\bm{\gamma}}_{\tilde{\theta}}+{\bm{\delta}}_{\tilde{\eta}})}\,,$
where we used that the support of $\psi$ gives us that $\mathbf{y}_{\kappa}\geq\mathbf{x}_{\kappa}$, as long as one chooses $\mathbf{c}$ small enough in ${\bm{\lambda}}\approx 2^{-\mathbf{n}}$. Starting from this estimate, we use (2.13) and get
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\prod_{i\in\theta}(\mbox{Id}-I^{i})g_{\mathbf{x}}^{\kappa,\mathbf{n}+\mathbf{k}}(\psi_{\mathbf{z}}^{{\bm{\lambda}}})\right\|_{m}=\left\|\sum_{\begin{subarray}{c}\chi\subset\theta\\ \theta\setminus\chi\subset\kappa\end{subarray}}\mathbb{E}^{\eta}_{\mathbf{x}}\sum_{\mathbf{y}_{\tilde{\theta}}\in\Delta_{\mathbf{n}+\mathbf{k}}^{\tilde{\theta}}}h(\mathbf{y}_{\tilde{\theta}})\right\|_{m}$
$\qquad\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}\left\|\psi\right\|_{C_{c}^{rd}}{\bm{\lambda}}^{{\bm{\alpha}}}_{{\kappa^{c}}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\kappa}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}\cdot 2^{-\mathbf{k}*{\bm{\alpha}},}_{\tilde{\theta}^{c}}2^{-r\mathbf{k}}_{\chi}$
$\qquad\qquad\sum_{\tilde{\theta}\setminus\tilde{\eta}=\theta_{1}\sqcup\theta_{2}}\underbrace{\left(\sum_{\mathbf{y}_{1}\in\Delta_{\mathbf{n}+\mathbf{k}}^{\theta_{1}\cup\tilde{\eta}}}2^{-\mathbf{k}*({\bm{1}}+{\bm{\gamma}}+{\bm{\delta}})}_{{\theta_{1}\cup\tilde{\eta}}}\right)}_{\lesssim 2^{-\mathbf{k}*({\bm{\gamma}}+{\bm{\delta}})}_{{\theta_{1}\cup\tilde{\eta}}}}\underbrace{\left(\sum_{\mathbf{y}_{2}\in\Delta_{\mathbf{n}+\mathbf{k}}^{\theta_{2}}}2^{-2\mathbf{k}*({\bm{1}}+{\bm{\gamma}})}_{{\theta_{2}}}\right)^{\frac{1}{2}}}_{\lesssim 2^{-\mathbf{k}*({\bm{\gamma}}+\frac{1}{2}{\bm{1}}),}_{{\theta_{2}}}}$
$\qquad\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}\left\|\psi\right\|_{C_{c}^{rd}}{\bm{\lambda}}^{{\bm{\alpha}}}_{{\kappa^{c}}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\kappa}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}\cdot 2^{-\mathbf{k}*({\bm{\alpha}}_{\theta^{c}}+({\bm{\alpha}}+r{\bm{1}})_{\chi}+(({\bm{\gamma}}+{\bm{\delta}})\wedge({\bm{\gamma}}+\frac{1}{2}{\bm{1}}))_{\tilde{\theta}})}$
$\qquad\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}\left\|\psi\right\|_{C_{c}^{rd}}{\bm{\lambda}}^{{\bm{\alpha}}}_{\kappa^{c}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\kappa}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}2^{-\mathbf{k}*{\bm{\alpha}}}_{{\theta^{c}}}2^{-\mathbf{k}*{\bm{\xi}}}_{\theta}\,,$
where we implicitly do not sum over all $\mathbf{y}$, but only the non-zero terms. This result is conditional on the possibility of applying Lemma 2.14. Unfortunately the step size of $\Delta_{\mathbf{n}+\mathbf{k}}$ is given by $\mathbf{u}=(2^{-n_{1}-k_{1}},\dots,2^{-n_{d}-k_{d}})$, so we would need $h(\mathbf{y}_{\tilde{\theta}})$ to be $\mathcal{F}^{\tilde{\theta}}_{\mathbf{y}+\mathbf{u}}$-measurable to apply Lemma 2.14. But since the support of $\hat{\phi}^{\chi}$ is in $[C,R]^{d}$ for some $0<C<R$ instead of $[0,1]^{d}$, $h(\mathbf{y}_{\tilde{\theta}})$ is $\mathcal{F}^{\tilde{\theta}}_{\mathbf{y}+2R\cdot\mathbf{u}}$-measurable. This can easily be solved by splitting the sum over $\mathbf{y}_{\tilde{\theta}}\in\Delta^{\tilde{\theta}}_{\mathbf{n}+\mathbf{k}}$ into finitely many sums over coarser grids:
$\sum_{\mathbf{y}_{\tilde{\theta}}\in\Delta_{\mathbf{n}+\mathbf{k}}^{\tilde{\theta}}}h(\mathbf{y}_{\tilde{\theta}})=\sum_{\mathbf{r}}\sum_{\mathbf{y}_{\tilde{\theta}}\in\Delta_{\mathbf{n}+\mathbf{k}}^{\tilde{\theta}}(\mathbf{r})}h(\mathbf{y}_{\tilde{\theta}})\,,$
see [Ker23], pages 19-20 for the details. This procedure only adds a constant depending on $R$ and the dimension $d$. ∎
With all these estimates, we can now tackle the proof of the reconstruction theorem.
Proof of Theorem 3.12.
Set $h_{\mathbf{k}}:=g^{\kappa,\mathbf{k}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{{\bm{\lambda}}})$ for a $\psi\in C_{c}^{rd}((0,1)^{\kappa}\times(-1,1)^{\kappa^{c}})$ and a $\mathbf{z}\in[0,T]^{d}$ such that $z_{i}=x_{i}$ for all $i\in\kappa$. Lemma 3.17 together with Lemma 3.15 immediately give us that $g^{\kappa,\mathbf{k}}_{\mathbf{x}}(\psi_{\mathbf{x}}^{{\bm{\lambda}}})$ is a Cauchy sequence in $L_{m}$. Thus, it is convergent. It follows by induction over $\sharp\kappa$, that all $\mathcal{R}^{\kappa,\mathbf{k}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}})$ are convergent sequences in $L_{m}$. Since $\mathcal{R}^{\kappa,\mathbf{k}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}})$ does not depend on $x_{i},i\in\kappa$ by construction, we can drop the assumption on $\mathbf{z}$ and find that $\mathcal{R}^{\kappa,\mathbf{k}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}})=\mathcal{R}^{\kappa,\mathbf{k}}_{\pi_{\mathbf{z}}^{\kappa}\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}})$ is convergent for all $\mathbf{z}\in[0,T]^{d}$ such that $\mbox{supp}(\psi_{\mathbf{z}}^{\bm{\lambda}})\subset(0,T)^{d}$. It follows that $\mathcal{R}^{\kappa,\mathbf{k}}_{\mathbf{x}}(\psi)$ is a convergent series for all $\kappa\subset[d]$, $\mathbf{x}\in[0,T]^{d}$ and $\psi$ with support in $(0,T)^{d}$.
We need to show that the limits $\mathcal{R}^{\eta}_{\mathbf{x}}(F)$ fulfill the properties in the statement. For this, let ${\bm{0}}<{\bm{\lambda}}<{\bm{1}}$ and $\mathbf{n}\in\mathbb{N}^{d}$ such that $2^{\mathbf{n}}\approx{\bm{\lambda}}$. Observe that again by Lemma 3.15, this time using $h_{\mathbf{k}}=\mathbb{E}^{\eta}_{\mathbf{x}}g^{\kappa,\mathbf{n}+\mathbf{k}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{{\bm{\lambda}}})$:
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}(g^{\kappa,\mathbf{n}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}})-g^{\kappa,\mathbf{n}+\mathbf{k}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}}))\right\|_{m}\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}\left\|\psi\right\|_{C_{c}^{rd}}{\bm{\lambda}}^{{\bm{\alpha}}}_{{\kappa^{c}}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\kappa}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}\,.$
So it suffices to show that
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}(g^{\kappa,\mathbf{n}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\lambda}))\right\|_{m}\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}\left\|\psi\right\|_{C_{c}^{rd}}{\bm{\lambda}}^{{\bm{\alpha}}}_{{\kappa^{c}}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\kappa}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}$ (3.17)
to show (3.10) by letting $\mathbf{k}\to\infty$. To see (3.17), we calculate
$g^{\kappa,\mathbf{n}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}})$ $=\sum_{\tilde{\kappa}\subset\kappa}(-1)^{\sharp{\tilde{\kappa}}}\mathcal{R}^{\tilde{\kappa},\mathbf{n}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}})$
$=\sum_{\mathbf{y}\in\Delta_{\mathbf{n}}}\sum_{\tilde{\kappa}\subset\kappa}(-1)^{\sharp{\tilde{\kappa}}}F_{\pi_{\mathbf{y}}^{\tilde{\kappa}}\mathbf{x}}(\phi_{\mathbf{y}}^{\mathbf{n}})\left\langle\phi_{\mathbf{y}}^{\mathbf{n}},\psi_{\mathbf{z}}^{{\bm{\lambda}}}\right\rangle$
$=\sum_{\mathbf{y}\in\Delta_{\mathbf{n}}}\square^{\kappa}_{\mathbf{x},\mathbf{y}}F(\phi_{\mathbf{y}}^{\mathbf{n}})\left\langle\phi_{\mathbf{y}}^{\mathbf{n}},\psi_{\mathbf{z}}^{{\bm{\lambda}}}\right\rangle\,.$
Using coherence and Lemma 2.8, we get that for all $\eta\subset\kappa$:
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}(g^{\kappa,\mathbf{n}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}}))\right\|_{m}\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}\left\|\psi\right\|_{C_{c}^{rd}}{\bm{\lambda}}^{{\bm{\alpha}}}_{{\kappa^{c}}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\kappa}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}\,.$
where we use that the order of non-zero terms is independent of $\mathbf{n}$ and that $\left|x_{i}-y_{i}\right|\lesssim\lambda_{i}$ for $i\in[d]$ and non-zero terms. It follows that for all $\mathbf{k}\geq{\bm{0}}$,
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}(g^{\kappa,\mathbf{n}+\mathbf{k}}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}}))\right\|_{m}\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}}\left\|\psi\right\|_{C_{c}^{rd}}{\bm{\lambda}}^{{\bm{\alpha}}}_{{\kappa^{c}}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\kappa}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}\,,$
which shows (3.10). We further note that all $\mathcal{R}^{\kappa}_{\mathbf{x}}(F)$ are linear maps from $C_{c}^{rd}\to\mathbb{R}$ and by the above inequality bounded in $\left\|\psi\right\|_{C_{c}^{rd}}$, so they are random distributions. By construction, $\mathcal{R}^{\kappa}_{\mathbf{x}}(F)$ is adapted and $\mathcal{R}_{\mathbf{x}}^{\kappa}(F)=\mathcal{R}_{\pi_{\mathbf{x}}^{\kappa}{\bm{0}}}(F)$. We finally show
$F_{\mathbf{x}}(\psi)=\lim_{\mathbf{n}\to\infty}\mathcal{R}^{\emptyset,\mathbf{n}}_{\mathbf{x}}(\psi)=\lim_{n\to\infty}F_{\mathbf{x}}(P_{\mathbf{n}}\psi)\,.$
We already know that $\mathcal{R}^{\emptyset,\mathbf{n}}_{\mathbf{x}}$ is a Cauchy-sequence and by [Ker23, Lemma 62], it follows that $F_{\mathbf{x}}=\lim_{n\to\infty}F_{\mathbf{x}}(P_{n{\bm{1}}}\psi)$, so the limit needs to be $F_{\mathbf{x}}$. This finishes the proof of convergence and to show the properties in Theorem 3.12.
It remains to show uniqueness. To do so, assume that $Z^{\eta}_{\mathbf{x}}$ has $Z^{\emptyset}_{\mathbf{x}}=0$ and fulfills properties 2-4. (Think $Z=\mathcal{R}-\tilde{\mathcal{R}}$ for two families $\mathcal{R},\tilde{\mathcal{R}}$ fulfilling 1-4). We show by induction over $\sharp\eta$, that $Z^{\eta}_{\mathbf{x}}=0$. Fix a $\theta\subset[d]$ and a $\psi_{\mathbf{z}}^{\bm{\lambda}}$ with $\mbox{supp}(\psi)\subset{[1/4,3/4]^{\theta}\times[-3/4,3/4]^{\theta^{c}}}$ and $\mathbf{z},{\bm{\lambda}}$ such that $[\mathbf{z}-{\bm{\lambda}}_{\theta^{c}},\mathbf{z}+{\bm{\lambda}}]\subset[0,T]^{d}$. Since $Z^{\theta}_{\mathbf{x}}$ does not depend on $\mathbf{x}_{\theta}$, we can without loss of generality assume that $z_{i}=x_{i}$ for all $i\in\theta$. Using (3.10) and the induction hypothesis, it follows that
$\left\|\mathbb{E}^{\eta}_{\mathbf{z}}Z^{\theta}_{\mathbf{x}}(\psi_{\mathbf{z}}^{\bm{\lambda}})\right\|_{m}\lesssim{\bm{\lambda}}^{{\bm{\alpha}}}_{{\theta^{c}}}{\bm{\lambda}}^{{\bm{\gamma}}}_{\theta}{\bm{\lambda}}^{{\bm{\delta}}}_{\eta}\,.$
The important point here is again, that since $Z^{\theta}_{\mathbf{x}}$ does not depend on $\mathbf{x}_{\theta}$, the above inequality holds for all $\mathbf{x},\mathbf{z}$, as long as the support of $\psi_{\mathbf{z}}^{\bm{\lambda}}$ lies in $(0,T)^{d}$. We can especially apply this to our wavelets to find
$\left\|\mathbb{E}^{\eta}_{\mathbf{y}}Z_{\mathbf{x}}^{\theta}(\hat{\phi}_{\mathbf{y}}^{\chi,\mathbf{n}})\right\|_{m}\lesssim 2^{-\mathbf{n}*(\frac{1}{2}{\bm{1}}+{\bm{\alpha}}_{\theta^{c}}+{\bm{\gamma}}_{\theta}+{\bm{\delta}}_{\eta})}\,,$
for all $\mathbf{y}\in\Delta_{\mathbf{n}}\cap[0,T)^{d}$, provided that $\mathbf{n}$ is big enough for the expression to be well-defined. Let us now pick a test function $\psi$ with support in $(0,T)^{d}$ and a high enough $\mathbf{N}\in\mathbb{N}^{d}$, such that $Z_{\mathbf{x}}^{\theta}(P_{\mathbf{N}}\psi)$ is well-defined. It then follows for all $\mathbf{n}\geq\mathbf{N}$:
$Z_{\mathbf{x}}^{\theta}(P_{\mathbf{n}}\psi)$ $=\sum_{\chi\subset\theta^{c}}\sum_{\mathbf{N}_{\chi}\leq\mathbf{m}_{\chi}\leq\mathbf{n}_{\chi}}Z_{\mathbf{x}}^{\theta}(\hat{P}^{\chi}_{\mathbf{n}_{\theta}+m_{\chi}}\psi)$
$=\sum_{\chi\subset\theta^{c}}\sum_{\mathbf{N}_{\chi}\leq\mathbf{m}_{\chi}\leq\mathbf{n}_{\chi}}\sum_{\mathbf{y}\in\Delta_{\mathbf{n}_{\theta}+\mathbf{m}_{\chi}}}Z_{\mathbf{x}}^{\theta}(\hat{\phi}^{\chi,\mathbf{n}_{\theta}+\mathbf{m}_{\chi}}_{\mathbf{y}})\left\langle\phi^{\chi,\mathbf{n}_{\theta}+\mathbf{m}_{\chi}}_{\mathbf{y}},\psi\right\rangle\,.$
Note that at most $\lesssim 2^{-n_{i}}$ terms are non-zero for each direction $i\in\theta$ and $2^{-m_{i}}$ for $i\in\chi$. We apply Lemma 2.14 to the sum over $\mathbf{y}_{\theta}$ and get
$\left\|\sum_{\mathbf{y}\in\Delta_{\mathbf{n}_{\theta}+\mathbf{m}_{\chi}}}Z_{\mathbf{x}}^{\theta}(\hat{\phi}^{\chi,\mathbf{n}_{\theta}+\mathbf{m}_{\chi}}_{\mathbf{y}})\left\langle\phi^{\chi,\mathbf{n}_{\theta}+\mathbf{m}_{\chi}}_{\mathbf{y}},\psi\right\rangle\right\|_{m}$ $\lesssim\sum_{\theta=\theta_{1}\sqcup\theta_{2}}2^{-(\mathbf{n}_{\theta}+\mathbf{m}_{\chi})*({\bm{\alpha}}_{\theta^{c}}+{\bm{\gamma}}_{\theta}+{\bm{\delta}}_{\theta_{1}}+\frac{1}{2}{\bm{1}}_{\theta_{2}}+r_{\chi})}$
$\lesssim 2^{-\mathbf{n}*({\bm{\gamma}}+{\bm{\delta}})}_{\theta}2^{-\mathbf{m}*({\bm{\alpha}}+\mathbf{r})}_{\chi}\,,$
where we used that $\delta_{i}\leq 1/2$ for $i\in[d]$ holds. Summing over $\mathbf{m}_{\chi}$, it gives us
$\left\|Z_{\mathbf{x}}^{\theta}(P_{\mathbf{n}}\psi)\right\|_{m}\lesssim 2^{-n*({\bm{\gamma}}+{\bm{\delta}})}_{\theta}\,,$
which implies that $Z^{\theta}_{\mathbf{x}}(\psi)=\lim_{n\to\infty}Z^{\theta}_{\mathbf{x}}(P_{n{\bm{1}}}\psi)=0$, where we used [Ker23, Lemma 62]. ∎
It remains to calculate the Hölder regularity of the reconstruction $\mathcal{R}(F)$, which we claim to be in $C^{{\bm{\alpha}},{\bm{0}}}L_{m}$. By the coherence property (3.8) with $\eta=\theta=\emptyset$, we have
$\left\|\mathcal{R}^{\emptyset}_{\mathbf{x}}(F)(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{m}=F_{\mathbf{x}}(\psi_{\mathbf{x}}^{\bm{\lambda}})\leq\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}L_{m}}{\bm{\lambda}}^{\bm{\alpha}}\,,$
for all $\psi\in B^{r}([1/4,3/4]^{d})$, $\mathbf{x}\in[0,T]^{d},{\bm{\lambda}}\in(0,1]^{d}$ such that $[\mathbf{x},\mathbf{x}+{\bm{\lambda}}]\subset[0,T]^{d}$. Using (3.10) together with the expansion
$\left\|\mathcal{R}^{\theta}_{\mathbf{x}}(F)(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{m}\leq\left\|\sum_{\hat{\theta}\subset\theta}(-1)^{\sharp{\hat{\theta}}}\mathcal{R}^{\hat{\theta}}_{\mathbf{x}}(F)(\psi^{{\bm{\lambda}}}_{\mathbf{x}})\right\|_{m}+\left\|\sum_{\hat{\theta}\subsetneq\theta}(-1)^{\sharp{\hat{\theta}}}\mathcal{R}^{\hat{\theta}}_{\mathbf{x}}(F)(\psi^{{\bm{\lambda}}}_{\mathbf{x}})\right\|_{m}$ (3.18)
inductively extends this argument to all $\mathcal{R}^{\theta}_{\mathbf{x}}(F),\theta\subset[d]$, as long as ${\bm{\gamma}}\geq{\bm{\alpha}}$. That is, for all $\theta\subset[d]$, one sees that
$\left\|\mathcal{R}^{\theta}_{\mathbf{x}}(F)(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{m}\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}L_{m}}{\bm{\lambda}}^{\bm{\alpha}}\,.$
Let us summarize this observation as a short corollary:
Corollary 3.18.
Under the condition of Theorem 3.12, the reconstruction $\mathcal{R}(F)$ is in $C^{{\bm{\alpha}}}L_{m}$ and
$\left\|\mathcal{R}(F)\right\|_{C^{\bm{\alpha}}L_{m}}\lesssim\left\|F\right\|_{G^{{\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}}L_{m}}\,,$ (3.19)
Proof.
This is a direct consequence of (3.18) together with (3.10).
Remark 3.19.
It is interesting to see that the stochastic parameter ${\bm{\delta}}$ is lost at the level of the reconstruction, i.e. $\mathcal{R}(F)$ will in general not be in $C^{\alpha,{\bm{\delta}}}L_{m}$. This is natural, as $\mathcal{R}(F)$, when evaluated against $\psi^{{\bm{\lambda}}}_{\mathbf{x}}$ locally behaves like $F_{\mathbf{x}}(\psi^{{\bm{\lambda}}}_{\mathbf{x}})$ which is only in $C^{\alpha}L_{m}$. However, the multiplication with a (stochastic) white noise $F=(Y\cdot\xi)$ as in (1.3) does result in $\mathcal{R}(F)\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ (in fact even $\mathcal{R}(F)\in C^{{\bm{\alpha}},\infty}L_{m}$), see Corollary 4.6 for details.
4 Application to a mixed SPDE
Using the results established in the previous section we will now show the existence and uniqueness of the equation (1.8). As explained in the introduction, we can write (1.8) into its mild formulation to obtain formally
$u(\mathbf{x})=\mathcal{I}(v)(\mathbf{x})+\int_{{\bm{0}}}^{\mathbf{x}}\sigma(u(\mathbf{y}))\xi(d\mathbf{y})+\int_{\bm{0}}^{\mathbf{x}}f(\mathbf{y})u(\mathbf{y})\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}Z(d\mathbf{y})\,,$
where the boundary function $I(v)\colon\mathbb{R}^{d}_{+}\to\mathbb{R}$ is explicitly given by
$\mathcal{I}(v)(\mathbf{x})=\sum_{\theta\subset\{1,\ldots d\},\;\theta\neq[d]}(-1)^{1+\sharp(\theta^{c})}v(\pi^{\theta}_{\mathbf{x}}{\bm{0}})\,.$ (4.1)
To derive it, it is sufficient to use the identity $\int_{\bm{0}}^{\mathbf{x}}\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}u(d\mathbf{y})=\square_{{\bm{0}},\mathbf{x}}^{[d]}u$ which holds at the level of smooth functions $u$ and then separate the term depending on the boundary function.
Both integrals in the above equation present formal products of distributions which cannot be defined point-wise but they will be interpreted rigorously via the action of reconstruction theorem and a proper integration map. More precisely, we can introduce the germs
$(\sigma(u)\cdot\xi)_{\mathbf{x}}(\psi):=\sigma(u(\mathbf{x}))\xi(\psi)\,,\quad((fu)\cdot\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}Z)_{\mathbf{x}}(\psi):=f(\mathbf{x})u(\mathbf{x})\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}Z(\psi)\,.$
The reconstruction of these two germs can be easily linked with Walsh integration [Wal86] and Young product [CZ20]. Then, as already explained in the introduction we will construct an integration map mapping a distribution onto its primitive function for all distributions $f\in C^{\bm{\beta}}L_{m}$ which are ${\bm{\beta}}>-{\bm{1}}$-Hölder. Putting all together, we are interested in solving the equation in $C^{\alpha,{\bm{\delta}}}L_{m}$
$u(\mathbf{x})=\mathcal{I}(v)(\mathbf{x})+\int_{{\bm{0}}}^{\mathbf{x}}\mathcal{R}(\sigma(u)\cdot\xi+(fu)\cdot\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}Z)(d\mathbf{y})\,.$ (4.2)
which will be constructed as a fixed point in $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ for some ${\bm{\alpha}},{\bm{\delta}}>{\bm{0}}$.
4.1 Composition and scalar multiplication with functions
We recall that for any given smooth function $g\colon\mathbb{R}\to\mathbb{R}$ and any $u\in C^{{\bm{\alpha}}}$, $g(u)\in C^{\bm{\alpha}}$, i.e. composition with smooth functions preserves the regularity (For reference, see [QST07, Lemma 3.1] or [BHR22, Lemma 15]). This, however, is no longer the case in the stochastic setting. This is due to the fact, that one would like to apply the Taylor formula to $g(u)$ and use $\square^{\{1\}}_{\mathbf{x},\mathbf{y}}u\square^{\{2\}}_{\mathbf{x},\mathbf{y}}u$ to bound $\square^{(1,2)}_{\mathbf{x},\mathbf{y}}g(u)$. While for deterministic $u$ this gives the correct bounds for $\left|\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}g(u)\right|$, in the stochastic setting $\square^{\{1\}}_{\mathbf{x},\mathbf{y}}u\square^{\{2\}}_{\mathbf{x},\mathbf{y}}u$ is no longer in $L_{m}$, but only in $L_{m/2}$. Due to this lack of integrability, we can only hope for bounds in $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m/d}$. To keep the same index $m$ invariant we will need to lower the indices ${\bm{\alpha}}$ and ${\bm{\delta}}$. In what follows for $\epsilon\in(0,1]$ we will use the notation $\mathcal{C}^{\epsilon}$ to denote the set of all $\epsilon$-Holder functions $g\colon\mathbb{R}\to\mathbb{R}$ with the convention of taking Lipschitz functions when $\epsilon=1$.
Proposition 4.1.
Let $u\in C^{\bm{\alpha}}L_{m}$ and $g\in\mathcal{C}^{\epsilon}$ for some $\epsilon\in(0,1]$ and ${\bm{\alpha}}\in(0,1]^{d}$. Then $g(u)\in C^{\frac{{\bm{\alpha}}\epsilon}{d}}L_{m}$ and one has
$\left\|g(u)\right\|_{C^{\frac{{\bm{\alpha}}\epsilon}{d}}L_{m}}\lesssim\left|g(0)\right|+\left\|g\right\|_{\mathcal{C}^{\epsilon}}\left\|u\right\|_{C^{\bm{\alpha}}L_{m}}\,,$
where the constant in $\lesssim$ is allowed to depend on $T$.
Proof.
The proof is based on the observation that for any $i\in[d]$ and $\mathbf{x},\mathbf{y}\in[0,T]^{d}$ one has
$\left\|\square^{\{i\}}_{\mathbf{x},\mathbf{y}}g(u)\right\|_{m}\leq\left\|g\right\|_{\mathcal{C}^{\epsilon}}\left\|\left|\square^{\{i\}}_{\mathbf{x},\mathbf{y}}u\right|^{\epsilon}\right\|_{m}\leq\left\|g\right\|_{\mathcal{C}^{\epsilon}}\left\|u\right\|^{\epsilon}_{C^{\bm{\alpha}}L_{m}}\left|x_{i}-y_{i}\right|^{\alpha_{i}\epsilon}\,,$
where we used Jensen’s inequality in the last step. If $\theta\subset[d]$ is an indexset and $i\in\theta$, we can use (2.5) to extend this estimate to $\square^{\theta}_{\mathbf{x},\mathbf{y}}g(u)$.
$\left\|\square^{\theta}_{\mathbf{x},\mathbf{y}}g(u)\right\|_{m}$ $\leq\sum_{\theta^{\prime}\subset\theta\setminus\{i\}}\left\|\square^{\{i\}}_{\pi^{\theta^{\prime}}_{\mathbf{y}}\mathbf{x},\mathbf{y}}g(u)\right\|_{m}$
$\lesssim\left\|g\right\|_{\mathcal{C}^{\epsilon}}\left\|u\right\|^{\epsilon}_{C^{\bm{\alpha}}L_{m}}\left|x_{i}-y_{i}\right|^{\alpha_{i}\epsilon}\,.$
Interpolation over $i\in\theta$ gives us for any non-empty $\theta\subset[d]$
$\left\|\square^{\theta}_{\mathbf{x},\mathbf{y}}g(u)\right\|_{m}$ $=\prod_{i\in\theta}\left\|\square^{\theta}_{\mathbf{x},\mathbf{y}}g(u)\right\|_{m}^{\frac{1}{\sharp\theta}}$
$\lesssim\left\|g\right\|_{\mathcal{C}^{\epsilon}}\left\|u\right\|^{\epsilon}_{C^{\bm{\alpha}}L_{m}}\prod_{i\in\theta}\left|x_{i}-y_{i}\right|^{\frac{\alpha_{i}\epsilon}{\sharp\theta}}\,.$
The claim now follows from the observation that
$\left\|g(u)\right\|_{m}\leq\left|g(0)\right|+\left\|g\right\|_{\mathcal{C}^{\epsilon}}\left\|u\right\|_{m}^{\epsilon}$
holds. ∎
In the case of the scalar multiplication of $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$, the product follows the same structure of the usual Hölder inequality.
Proposition 4.2.
Let $f\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{n}$ and $u\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ with $0\leq{\bm{\delta}}\leq{\bm{\alpha}}$. Then the scalar multiplication fulfills $fu\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{k}$ where $\frac{1}{k}=\frac{1}{m}+\frac{1}{n}$. Furthermore, we have the inequality
$\left\|fu\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{k}}\lesssim\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{n}}\left\|u\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}$
Proof.
The proof is based on the decomposition (2.4). Note that by the adaptedness property of $f$ and $u$, we have that $\square_{\mathbf{x},\mathbf{y}}^{\theta}f,\square_{\mathbf{x},\mathbf{y}}^{\theta}u$ are $\mathcal{F}_{\pi_{\mathbf{x}}^{\theta^{c}}\mathbf{y}}$-measurable for any $\mathbf{x}\leq\mathbf{y}$ and $\theta\subset[d]$.
Let $\mathbf{x}\leq\mathbf{y}$ and $\eta\subset\theta\subset[d]$ with $\theta\neq\emptyset$. Whenever we are given two sets $\theta_{1},\theta_{2}\subset\theta$ with $\theta=\theta_{1}\cup\theta_{2}$, we decompose $\eta$ into three disjoint sets
$\eta_{1}=\eta\cap(\theta_{1}\setminus\theta_{2}),\qquad\eta_{2}=\eta\cap(\theta_{2}\setminus\theta_{1}),\qquad\eta_{3}=\eta\cap\theta_{1}\cap\theta_{2}\,.$
One now calculates, using the adaptedness of $\square^{\theta_{1}}_{\mathbf{x},\mathbf{y}}f,\square^{\theta_{2}}_{\mathbf{x},\mathbf{y}}u$ and the Hölder inequality:
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square^{\theta}_{\mathbf{x},\mathbf{y}}(fu)\right\|_{k}$ $\leq\sum_{\theta=\theta_{1}\cup\theta_{2}}\left\|\mathbb{E}^{\eta}_{\mathbf{x}}(\square^{\theta_{1}}_{\mathbf{x},\mathbf{y}}f\square^{\theta_{2}}_{\mathbf{x},\mathbf{y}}u)\right\|_{k}$
$=\sum_{\theta=\theta_{1}\cup\theta_{2}}\left\|\mathbb{E}^{\eta_{3}}_{\mathbf{x}}\left[(\mathbb{E}^{\eta_{1}}_{\mathbf{x}}\square^{\theta_{1}}_{\mathbf{x},\mathbf{y}}f)(\mathbb{E}^{\eta_{2}}_{\mathbf{x}}\square^{\theta_{2}}_{\mathbf{x},\mathbf{y}}u)\right]\right\|_{k}$
$\leq\sum_{\theta=\theta_{1}\cup\theta_{2}}\left\|\mathbb{E}^{\eta_{1}}_{\mathbf{x}}\square^{\theta_{1}}_{\mathbf{x},\mathbf{y}}f\right\|_{n}\left\|\mathbb{E}^{\eta_{2}}_{\mathbf{x}}\square^{\theta_{2}}_{\mathbf{x},\mathbf{y}}u\right\|_{m}$
$\leq\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{n}}\left\|u\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\sum_{\theta=\theta_{1}\cup\theta_{2}}\left|\mathbf{y}-\mathbf{x}\right|_{\text{c},\theta}^{{\bm{\alpha}}_{\theta}+{\bm{\delta}}_{\eta_{1}\cup\eta_{2}}+{\bm{\alpha}}_{\eta_{3}}}\,.$
Using also the standard Hölder inequality
$\left\|f(\mathbf{x})u(\mathbf{x})\right\|_{k}\leq\left\|f(\mathbf{x})\right\|_{n}\left\|u(\mathbf{x})\right\|_{m}$
for all $\mathbf{x}\in[0,T]^{d}$, one obtains the claim. ∎
4.2 Products of processes and distributions
This section discusses how multiplying a random field with a random distribution results in a stochastic coherent germ. More precisely, we want to understand under which conditions we can reconstruct the germ
$(u\cdot\zeta)_{\mathbf{x}}(\psi):=u(\mathbf{x})\zeta(\psi)$ (4.3)
when $u\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ and $\zeta\in C^{{\bm{\beta}},{\bm{\delta}}^{\prime}}L_{n}$. Instead of a generic analysis, we will consider two different cases: The case in which $\zeta$ is deterministic, so we will need to use the stochastic properties of $u$, and the case in which $\zeta=\xi$ is the white noise we will obtain a coherent germ by taking simply $u\in C^{{\bm{\alpha}}}L_{m}$.
Deterministic distributions
In the case in which $\zeta\in C^{\bm{\beta}}$ for some ${\bm{\beta}}<{\bm{0}}$, any stochastic improvement needs to come from the process $u\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$, so we will assume ${\bm{\delta}}>{\bm{0}}$. Note that we do not expect the reconstruction to have stochastic properties, as it behaves locally like an integrable random field times an element of $C^{\bm{\beta}}$, see also Remark 3.19.
Proposition 4.3.
Let $u\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ and $\zeta\in C^{{\bm{\beta}}}$ for ${\bm{\alpha}},{\bm{\delta}}>0$ and ${\bm{\beta}}<0$. Then $u\cdot\zeta\in G^{{\bm{\beta}},{\bm{\alpha}}+{\bm{\beta}},{\bm{\delta}}}L_{m}$ and
$\left\|(u\cdot\zeta)\right\|_{G^{{\bm{\beta}},{\bm{\alpha}}+{\bm{\beta}},{\bm{\delta}}}L_{m}}\leq\left\|u\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\left\|\zeta\right\|_{C^{\bm{\beta}}}\,.$ (4.4)
It follows that $u\cdot\zeta$ can be reconstructed, as long as ${\bm{\alpha}}+{\bm{\beta}}>-\frac{1}{2}{\bm{1}}$, ${\bm{\alpha}}+{\bm{\beta}}+{\bm{\delta}}>{\bm{0}}$ and $m\geq 2$. In this case, we have $\mathcal{R}(u\cdot\zeta)\in C^{{\bm{\beta}}}L_{m}$ and the estimate
$\left\|\mathcal{R}(u\cdot\zeta)\right\|_{C^{\bm{\beta}}L_{m}}\lesssim\left\|u\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\left\|\zeta\right\|_{C^{\bm{\beta}}}\,.$ (4.5)
Proof.
$u\cdot\zeta\in G^{{\bm{\beta}},{\bm{\alpha}}+{\bm{\beta}},{\bm{\delta}}}L_{m}$ as well as (4.4) follow immediately from the identity $\square^{\theta}_{\mathbf{x},\mathbf{y}}(u\cdot\zeta)(\psi)=(\square_{\mathbf{x},\mathbf{y}}^{\theta}u)\zeta(\psi)$ and definition of stochastic coherent germ. (4.5) is a direct consequence of Corollary 3.18. ∎
Product with white noise and stochastic integration
We now consider the case of the product (4.3) when $\zeta=\xi$ is white noise. Since $\xi\in C^{-\frac{1}{2}{\bm{1}},\infty}L_{m}$ regularises with “infinite” regularity the conditional expectation we allow ourselves to consider integrands $u\in C^{{\bm{\alpha}}}L_{m}$. In this case, the reconstruction coincides with the stochastic integral with respect to the Brownian sheet, see [Ker23, Section 7], [CW75] so we expect the reconstruction $f$ to have more properties than the ones given by the reconstruction theorem.
Proposition 4.4.
Let $u\in C^{{\bm{\alpha}}}L_{m}$, ${\bm{\alpha}}>{\bm{0}}$ and $\xi$ be a white noise over +$\mathbb{R}^{d}_{+}$. Then the product germ $u\cdot\xi\in G^{-\frac{1}{2}{\bm{1}},-\frac{1}{2}{\bm{1}}+{\bm{\alpha}},\infty}L_{m}$ and for all test functions $\psi\in C^{r}_{c}((0,T)^{d})$, we have the identity
$\mathcal{R}(u\cdot\xi)(\psi)=\int_{{\bm{0}}}^{+\infty}\psi(\mathbf{y})u(\mathbf{y})\xi(d\mathbf{y})$
where the right-hand side is the Walsh integral associated with the white noise. Moreover, we have for $m\geq 2$
$\left\|\mathcal{R}(u\cdot\xi)(\psi)\right\|_{m}\lesssim\sup_{\mathbf{x}\in[0,T]^{d}}\left\|u(\mathbf{x})\right\|_{m}\left\|\psi\right\|_{L_{2}([0,T]^{d})}\,.$ (4.6)
Furthermore, if $(\mathcal{F}_{\mathbf{x}})_{\mathbf{x}\in[0,T]^{d}}$ is the filtration generated by $\xi$, for any $\mathbf{x}\in[0,T]^{d}$, $i\in[d]$, and any test function $\psi$ with $\mbox{supp}(\psi)\subset(x_{1},T]\times\dots\times(x_{d},T]$ we have
$\mathbb{E}^{i}_{\mathbf{x}}\mathcal{R}(u\cdot\xi)(\psi)=0\,.$ (4.7)
Remark 4.5.
For the classical Walsh integral, (4.6) is a direct consequence of the Itô-isometry. Recall that in the case $m=2$ and for an adapted process $u$ the Walsh integral fulfills
[…]
$\leq\sup_{\mathbf{x}\in[0,T]^{d}}\left\|u(x)\right\|_{2}\left\|\psi\right\|_{L_{2}([0,T]^{d})}\,.$
The general case with $m\geq 2$ generic was treated in [Cai70] but since we are in a distributional setting we present a proof of (4.6) which derives this directly from the reconstruction operator.
Proof.
We start by showing $u\cdot\xi\in G^{-\frac{1}{2}{\bm{1}},-\frac{1}{2}{\bm{1}}+{\bm{\alpha}},\infty}L_{m}$. Observe that for all $\mathbf{x},\mathbf{y}\in[0,T]^{d}$ and $\theta\subset[d]$, we have
$\square_{\mathbf{x},\mathbf{y}}^{\theta}(u\cdot\xi)(\psi)=(\square^{\theta}_{\mathbf{x},\mathbf{y}}u)\xi(\psi)\,.$
For a $\psi$ with support in $[y_{1},T]\times\dots\times[y_{d},T]$ and $x_{i}\leq y_{i}$ for at least one $i\in[d]$, we get by the adaptedness of $u$ and the fact that $\xi(\psi)$ is independent of $\mathcal{F}^{\{i\}}_{y_{i}}$:
$\left\|\square_{\mathbf{x},\mathbf{y}}^{\theta}(u\cdot\xi)(\psi)\right\|_{m}=\left\|\square^{\theta}_{\mathbf{x},\mathbf{y}}u\right\|_{m}\left\|\xi(\psi)\right\|_{m}$
as well as
$\mathbb{E}^{\eta}_{\mathbf{x}}\square^{\theta}_{\mathbf{x},\mathbf{y}}(u\cdot\xi)(\psi)=\mathbb{E}^{\eta}_{\mathbf{x}}[\square_{\mathbf{x},\mathbf{y}}^{\theta}u\leavevmode\nobreak\ \mathbb{E}^{i}_{\mathbf{y}}\xi(\psi)]=0\,,$
for any $\eta\subset[d],i\in\eta$ such that $x_{i}\leq y_{i}$ and $\psi$ as above. It follows easily that $u\cdot\xi\in G^{-\frac{1}{2}{\bm{1}},-\frac{1}{2}{\bm{1}}+{\bm{\alpha}},\infty}L_{m}$. To see (4.6), recall that
$\mathcal{R}(u\cdot\xi)(\psi)=\lim_{\mathbf{n}\to\infty}\sum_{\mathbf{x}\in\Delta_{\mathbf{n}}}(u\cdot\xi)_{\mathbf{x}}(\phi_{\mathbf{x}}^{\mathbf{n}})\left\langle\phi^{\mathbf{n}}_{\mathbf{x}},\psi\right\rangle\,.$
We can use Lemma 2.14 together with $\mathbb{E}^{i}_{\mathbf{x}}((u\cdot\xi)_{\mathbf{x}}(\phi_{\mathbf{x}}^{n}))=0$ for all $i\in[d]$, $\mathbf{x}\in\Delta_{\mathbf{n}}$ and $\mathbf{n}\in\mathbb{N}^{d}$, to show
$\left\|\sum_{\mathbf{x}\in\Delta_{\mathbf{n}}}(u\cdot\xi)_{\mathbf{x}}(\phi_{\mathbf{x}}^{\mathbf{n}})\left\langle\phi^{\mathbf{n}}_{\mathbf{x}},\psi\right\rangle\right\|_{m}$ $\lesssim\left(\sum_{\mathbf{x}\in\Delta_{\mathbf{n}}}\left\|u(\mathbf{x})\right\|_{m}^{2}\left\|\xi(\phi_{\mathbf{x}}^{\mathbf{n}})\right\|_{m}^{2}\left|\left\langle\phi^{\mathbf{n}}_{\mathbf{x}},\psi\right\rangle\right|^{2}\right)^{\frac{1}{2}}$
$\leq\sup_{\mathbf{x}\in[0,T]^{d}}\left\|u(\mathbf{x})\right\|_{m}\left(\sum_{\mathbf{x}\in\Delta_{\mathbf{n}}}\left\|\phi_{\mathbf{x}}^{\mathbf{n}}\right\|_{L_{2}(\mathbb{R}^{d})}^{2}\left|\left\langle\phi^{\mathbf{n}}_{\mathbf{x}},\psi\right\rangle\right|^{2}\right)^{\frac{1}{2}}$
$=\sup_{\mathbf{x}\in[0,T]^{d}}\left\|u(\mathbf{x})\right\|_{m}\left\|P_{\mathbf{n}}\psi\right\|_{L_{2}(\mathbb{R}^{d})}\,.$
(4.6) then follows from $\left\|\psi\right\|_{L_{2}(\mathbb{R}^{d})}=\lim_{\mathbf{n}\to\infty}\left\|P_{\mathbf{n}}\psi\right\|_{L_{2}(\mathbb{R}^{d})}$ To see (4.7), we fix $\psi$ with support in $(y_{1},T]\times\dots\times(y_{d},T]$. Thus, there exists a $\mathbf{z}>\mathbf{y}$ such that
$\mbox{supp}(\psi)\subset[z_{1},T]\times\dots[z_{d},T]\subset(y_{1},T]\times\dots\times(y_{d},T]\,.$
Assume $\mathbf{x},\mathbf{n}$ are chosen in such a way, that $\left\langle\phi_{\mathbf{x}}^{\mathbf{n}},\psi\right\rangle\neq 0$. Then the support of the functions must overlap, and by construction of our wavelets (recall $\mbox{supp}(\phi)\subset[C,R]^{d}$), we can conclude that
$\mbox{supp}(\phi_{\mathbf{x}}^{\mathbf{n}})\subset[z_{1}-2^{-n_{1}}(R-C),T+2^{-n_{1}}(R-C)]\times\dots\times[z_{d}-2^{-n_{d}}(R-C),T+2^{-n_{d}}(R-C)]\,.$
Thus, provided that $\mathbf{n}$ is large enough such that $z_{i}-2^{-n_{i}}(R-C)>y_{i}$ for $i\in[d]$ and denoting $\delta_{i}:=2^{-n_{i}}(R-C)$, we have for all $\mathbf{x}\in\Delta_{\mathbf{n}}$ that at least one of the following properties holds:
$\mbox{supp}(\phi_{\mathbf{x}}^{\mathbf{n}})\subset(y_{1},T+\delta_{1}]\times\dots\times(y_{d},T+\delta_{d}]\quad\text{or}\quad\left\langle\phi^{\mathbf{n}}_{\mathbf{x}},\psi\right\rangle=0\,.$
Since both cases imply that
$\mathbb{E}^{i}_{\mathbf{y}}(u\cdot\xi)_{\mathbf{x}}(\phi_{\mathbf{x}}^{\mathbf{n}})\left\langle\phi_{\mathbf{x}}^{\mathbf{n}},\psi\right\rangle=0$
for any $i\in[d]$, we conclude that
$\mathbb{E}^{i}_{\mathbf{y}}\mathcal{R}(u\cdot\xi)(\psi)=\lim_{\mathbf{n}\to\infty}\sum_{\mathbf{x}\in\Delta_{\mathbf{n}}}\mathbb{E}^{i}_{\mathbf{x}}(u\cdot\xi)_{\mathbf{x}}(\phi_{\mathbf{x}}^{\mathbf{n}})\left\langle\phi_{\mathbf{x}}^{\mathbf{n}},\psi\right\rangle=0\,.$
Gathering all the above inequalities, one has simpler estimates on $\mathcal{R}(u\cdot\xi)$.
Corollary 4.6.
Let $u\in C^{{\bm{\alpha}}}L_{m}$, ${\bm{\alpha}}>{\bm{0}}$ and $\xi$ be white noise over $\mathbb{R}_{+}^{d}$. Then $\mathcal{R}(u\cdot\xi)\in C^{-\frac{1}{2}{\bm{1}},\infty}L_{m}$ and we have
$\left\|\mathcal{R}(u\cdot\xi)\right\|_{C^{-\frac{1}{2}{\bm{1}},\infty}L_{m}}\lesssim\left\|u\right\|_{C^{\bm{\alpha}}L_{m}}\,.$ (4.8)
If $g\in\mathcal{C}^{1}$ is Lipschitz with Lipschitz constant $\left\|g\right\|_{\mathcal{C}^{1}}$ and $u,v\in C^{\bm{\alpha}}L_{m}$, we have
$\left\|\mathcal{R}(g(u)\cdot\xi)-\mathcal{R}(g(v)\cdot\xi)\right\|_{C^{-\frac{1}{2}{\bm{1}},\infty}L_{m}}\lesssim\left\|g\right\|_{\mathcal{C}^{1}}\left\|u-v\right\|_{C^{\bm{\alpha}}L_{m}}$ (4.9)
Proof.
Observe that for all $\mathbf{x}\in[0,T]^{d},{\bm{\lambda}}\in(0,1]^{d}$ we have
$\sup_{\mathbf{x}\in[0,T]^{d}}\left\|u(\mathbf{x})\right\|_{m}\left\|\psi_{\mathbf{x}}^{\bm{\lambda}}\right\|_{L_{2}([0,T]^{d})}\leq\left\|u\right\|_{C^{{\bm{\alpha}}}L_{m}}{\bm{\lambda}}^{-\frac{1}{2}{\bm{1}}}\left\|\psi\right\|_{L_{2}([0,T]^{d})}\,.$
Putting this in (4.6), together with (4.7) we obtain (4.8). To see (4.9), we use (4.6) to calculate
$\left\|(\mathcal{R}(g(u)\cdot\xi)-\mathcal{R}(g(v)\cdot\xi))(\psi)\right\|_{m}$ $=\left\|(\mathcal{R}(g(u)-g(v))\cdot\xi)(\psi)\right\|_{m}$
$\lesssim\sup_{\mathbf{x}\in[0,T]^{d}}\left\|g(u(\mathbf{x}))-g(v(\mathbf{x}))\right\|_{m}\left\|\psi\right\|_{L_{2}(\mathbb{R}^{d})}$
$\lesssim\left\|g\right\|_{\mathcal{C}^{1}}\sup_{\mathbf{x}\in[0,T]^{d}}\left\|u(\mathbf{x})-v(\mathbf{x})\right\|_{m}\left\|\psi\right\|_{L_{2}(\mathbb{R}^{d})}\,.$
The relation (4.9) then follows with the same argument as (4.8). ∎
4.3 Primitives and integration
We finally provide the last step to formulate (4.2) rigorously. That is an integration map
$f\mapsto\int_{{\bm{0}}}^{\mathbf{x}}f(d\mathbf{y})$
to recover a function after obtaining a distribution. Following the same approach of [Bra19, Lemma 3.10] we show that our stochastic spaces $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ behave as one expects under integration. To do so, we fix a wavelet basis $\Phi=\{\hat{\phi}^{\mathbf{n},\zeta}_{\mathbf{x}}\colon\mathbf{n}\in\mathbb{N}^{d},\mathbf{x}\in\Delta_{\mathbf{n}},\zeta\subset[d]\}$ like in Definition 2.7 and we denote by $P_{\mathbf{n}}$ the projection onto the respective spaces.
Let us address a problem arising from the wavelet approximation applied to $1_{[\mathbf{s},\mathbf{t}]}$ for any $0\leq\mathbf{s}\leq\mathbf{t}\leq T{\bm{1}}$. Since the approximation enlarges the support, for any fixed $\mathbf{n}\in\mathbb{N}^{d}$, $P_{\mathbf{n}}1_{[\mathbf{s},\mathbf{t}]}$ might be supported outside of $[0,T]^{d}$, causing undefined terms. We further will struggle to condition the term onto $\mathcal{F}_{\mathbf{s}}$, due to the same fattening effect. The first problem can be solved by extending our distribution $f$ to $\mathbb{R}^{d}$ by setting $\tilde{f}(\psi)=f(\psi|_{[0,T]^{d}})$, but this does not solve the second problem. Because of this, we use another approach and only consider $\mathbf{s},\mathbf{t}$ which have a certain distance ${\bm{\epsilon}}$ to the boundary. We further condition these points on a $\hat{\mathbf{s}}<\mathbf{s}$, also depending on the chosen ${\bm{\epsilon}}$. Under these conditions, we can use Brault’s construction and we will call the result the interior primitive. We can then use a dyadic argument to pass to the whole space $\mathbf{s},\mathbf{t}\in[0,T]^{d}$ and condition on $\mathcal{F}_{\mathbf{s}}$ sharply.
Proposition 4.7.
Let $f\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ with $-{\bm{1}}<{\bm{\alpha}}<{\bm{0}}$ Then $f(P_{\mathbf{n}}1_{(\mathbf{s},\mathbf{t}]})$ converges in $L_{m}$ for all $0<\mathbf{s}\leq\mathbf{t}<{\bm{1}}T$ as $\mathbf{n}\to\infty$. We denote its limit by
$\oint_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y})\,.$
and we call it the interior primitive of $f$. The resulting limit satisfies the a.s. identity
$\oint_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y})=\oint_{\mathbf{s}}^{\pi^{i}_{\mathbf{u}}\mathbf{t}}f(d\mathbf{y})+\oint_{\pi^{i}_{\mathbf{u}}\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y})$ (4.10)
for any $\mathbf{s}<\mathbf{u}<\mathbf{t}\in(0,T)^{2}$ and $i\in[d]$. Furthermore, for each $\emptyset\neq\eta\subset[d]$, ${\bm{\epsilon}}>{\bm{0}}$ we have the estimate
$\left\|\oint_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y})\right\|_{m}$ $\lesssim{\bm{\epsilon}}^{{\bm{\alpha}}}\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}^{{\bm{\alpha}}+{\bm{1}}}$ (4.11)
$\left\|\mathbb{E}^{\eta}_{\hat{\mathbf{s}}}\oint_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y})\right\|_{m}$ $\lesssim{\bm{\epsilon}}^{{\bm{\alpha}}}\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}^{{\bm{\alpha}}+{\bm{1}}}\left|\mathbf{t}-\mathbf{s}\right|^{{\bm{\delta}}}_{\text{c},\eta}\,,$ (4.12)
where $\mathbf{s},\mathbf{t}\in[0,T]^{d}$ satisfy $\hat{\mathbf{s}}:=\mathbf{s}-\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}*{\bm{\epsilon}}\in[0,T]^{d}$ and $\hat{\mathbf{t}}:=\mathbf{t}+\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}*{\bm{\epsilon}}\in[0,T]^{d}$, and the constant in $\lesssim$ is allowed to depend on $\lambda$.
Remark 4.8.
Note that the distance between $\hat{s}$ and $s$, as well as $\hat{t}$ and $t$, depends on $\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}$. This choice allows us to later use a dyadic argument.
Remark 4.9.
Property (4.10) extends the usual additive property of a two-parameter function $A\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$
$A(s,t)=A(s,u)+A(u,t)$
at the level of $d$-parameter functions see [BHK23, Definition 4]. From this condition one identifies for any $\mathbf{s}\in(0,T)^{d}$ a unique function
$\mathbf{y}\mapsto\oint_{\mathbf{s}}^{\mathbf{y}}f(d\mathbf{z})=:I_{\mathbf{s},\mathbf{x}}(f)$
such that $I_{\mathbf{s},\mathbf{y}}(f)=0$ for any $\mathbf{y}$ such that $\pi^{i}\mathbf{y}=\pi^{i}\mathbf{s}$ for some $i\in[d]$ and
$\oint_{\mathbf{x}}^{\mathbf{y}}f(d\mathbf{z})=\square_{\mathbf{x},\mathbf{y}}^{[d]}I_{\mathbf{s},\cdot}(f)$
for any $\mathbf{s}\leq\mathbf{x}\leq\mathbf{y}\in(0,T)^{d}$. Since we cannot choose $\mathbf{s}={\bm{0}}$ at this stage, we prefer to think of the interior primitive as a $2d$-parameter object and will only go the $d$-parameter setting with the primitive itself.
Proof.
Given $\mathbf{s}$ and $\mathbf{t}$ we choose $\mathbf{n}$ such that $2^{-n_{i}}\leq\frac{\epsilon_{i}}{R}\left|t_{i}-s_{i}\right|\leq 2^{-n_{i}+1}$ for $i\in[d]$, where $R$ comes from the wavelet $\phi$ (recall from Section 2 that $\mbox{supp}(\phi)\subset[C,R]^{d}$ for some $0<C<R$). Then for any $\mathbf{k}\geq\mathbf{n}$ we write
$P_{\mathbf{k}}1_{(\mathbf{s},\mathbf{t}]}=\sum_{\chi\subset[d]}\leavevmode\nobreak\ \sum_{\begin{subarray}{c}\mathbf{n}_{\chi}\leq\mathbf{m}_{\chi}\leq\mathbf{k}_{\chi}\\ \mathbf{m}_{\chi^{c}}=\mathbf{n}_{\chi^{c}}\end{subarray}}\hat{P}^{\chi}_{\mathbf{m}}1_{(\mathbf{s},\mathbf{t}]}\,,$ (4.13)
In order to prove the convergence of $P_{\mathbf{k}}1_{(\mathbf{s},\mathbf{t}]}$ it suffices to find a proper bound on $f(\hat{P}^{\chi}_{\mathbf{m}}1_{[\mathbf{s},\mathbf{t}]})$ , which we can estimate as
$\left\|f(\hat{P}_{\mathbf{m}}^{\chi}1_{[\mathbf{s},\mathbf{t}]})\right\|_{m}\leq\sum_{\mathbf{x}\in\Delta_{\mathbf{m}}}\left\|f(\hat{\phi}_{\mathbf{x}}^{\chi,\mathbf{m}})\right\|_{m}\left|\left\langle\hat{\phi}_{\mathbf{x}}^{\chi,\mathbf{m}},1_{(\mathbf{s},\mathbf{t}]}\right\rangle\right|\,.$
By our choice of $\mathbf{n}$, all non-zero terms $\mathbf{x}\in\Delta_{\mathbf{m}}$ satisfy $\mathbf{x}\in[\hat{\mathbf{s}},\mathbf{t}]$ and the support of $\hat{\phi}^{\chi,\mathbf{m}}_{\mathbf{x}}$ is in $[\hat{\mathbf{s}},\hat{\mathbf{t}}]\subset[0,T]^{d}$. For any fixed subset $\chi\subset[d]$ we note that in each direction $i\notin\chi$, the number of non-zero terms is of order $2^{-m_{i}}\left|t_{i}-s_{i}\right|$. Moreover, note that for the one-dimensional wavelet $\hat{\varphi}$ we have
$\left\langle\hat{\varphi},1_{(s,t]}\right\rangle=0\,,$
unless the support of $\hat{\varphi}$ contains one of the points $s,t\in[0,T]$. It follows that for all $i\in\chi$, the number of $x_{i}\in\Delta_{m_{i}}$ such that $\left\langle\hat{\varphi}^{m_{i}}_{x_{i}},1_{(s_{i},t_{i}]}\right\rangle$ is non-zero is a number independent of $m_{i}$ or $\left|t_{i}-s_{i}\right|$. Thus, by counting the non-zero terms and using Lemma 2.8 with $r=0$, we get
$\begin{split}\left\|f(\hat{P}_{\mathbf{m}}^{\chi}1_{(\mathbf{s},\mathbf{t}]})\right\|_{m}&\lesssim\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}2^{-\mathbf{m}*{\bm{\alpha}}}2^{-\mathbf{m}}_{\chi}\left|\mathbf{t}-\mathbf{s}\right|^{{\bm{1}}}_{\text{c},\chi^{c}}\\ &=\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}2^{-\mathbf{m}*{\bm{\alpha}}}_{\chi^{c}}2^{-\mathbf{m}*({\bm{\alpha}}+{\bm{1}})}_{\chi}\left|\mathbf{t}-\mathbf{s}\right|^{{\bm{1}}}_{\text{c},\chi^{c}}\,.\end{split}$ (4.14)
Note that the support of our wavelet $\phi$ is in general not in $[1/4,3/4]^{d}$. But we can always chose $l\in\mathbb{N}$ and $\mathbf{y}\in\Delta_{l{\bm{1}}}$ such that $\phi^{l{\bm{1}}}_{\mathbf{y}}$ has support in $[1/4,3/4]^{d}$, allowing us to use (3.6) on $f(\phi^{\mathbf{m}}_{\mathbf{x}})=f(\phi^{l{\bm{1}}_{\mathbf{y}}})^{\mathbf{m}-l{\bm{1}}}_{\mathbf{x}-2^{l{\bm{1}}-\mathbf{m}}\mathbf{y}}$ for any $\mathbf{m}\geq l{\bm{1}}$. This only adds a constant depending on the wavelet to our calculations.
[…]
$\left\|f(P_{\mathbf{k}}1_{(\mathbf{s},\mathbf{t}]})\right\|_{m}\lesssim\sum_{\chi\subset[d]}\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}2^{-\mathbf{n}*{\bm{\alpha}}}_{\chi^{c}}2^{-\mathbf{n}({\bm{\alpha}}+{\bm{1}})}_{\chi}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c},\chi^{c}}^{\bm{1}}\,.$
Letting $\mathbf{k}\to\infty$, (4.11) then follows from $2^{-n_{i}}\approx\epsilon_{i}\left|t_{i}-s_{i}\right|$. The estimate (4.12) follows analogously. To see the identity (4.10), we simply observe that $P_{\mathbf{n}}1_{(\mathbf{s},\mathbf{t}]}$ is additive for each $\mathbf{n}\in\mathbb{N}^{d}$. ∎
Due to the presence of the variable ${\bm{\epsilon}}$ in Proposition 4.7, the interior primitive of $f$ does not suffice to define a proper integration over the whole cube $[0,T]^{d}$ when $\mathbf{s}$ and $\mathbf{t}$ approach the boundary. If only $\mathbf{s}$ or $\mathbf{t}$ approaches the boundary while the other one has some fixed distance to it, we can absorb this singularity via a dyadic expansion. Thanks to the additive property (4.10), this is an extension to the interior primitive.
Lemma 4.10.
Let $f\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m},-{\bm{1}}<{\bm{\alpha}}<{\bm{0}}$ and ${\bm{0}}\leq\mathbf{s}\leq\mathbf{t}<T{\bm{1}}$ such that $\mathbf{t}+\frac{1}{2}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}\in[0,T]^{d}$. For any $\mathbf{k}\in\mathbb{N}^{d}$, we define $\mathbf{s}_{\mathbf{k}}$ by $(\mathbf{s}_{\mathbf{k}})_{i}:=s_{i}+2^{-k_{i}}\left|t_{i}-s_{i}\right|,i\in[d]$. Then, the limit
$\int_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y}):=\sum_{\mathbf{k}\geq{\bm{1}}}\oint_{\mathbf{s}_{\mathbf{k}}}^{\mathbf{s}_{\mathbf{k}-{\bm{1}}}}f(d\mathbf{y})$
exists in $L_{m}$. It satisfies the additivity property (4.10) and is $\mathcal{F}_{\mathbf{t}}$-measurable. Furthermore, we have
$\left\|\int_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y})\right\|_{m}$ $\lesssim\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}^{{\bm{\alpha}}+{\bm{1}}}$ (4.15)
$\left\|\mathbb{E}^{\eta}_{\mathbf{s}}\int_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y})\right\|_{m}$ $\lesssim\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}^{{\bm{\alpha}}+{\bm{1}}}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c},\eta}^{{\bm{\delta}}}\,.$ (4.16)
for all $\eta\subset[d]$.
Proof.
By construction of $\mathbf{s}_{\mathbf{k}}$, we have for all $\mathbf{k}\geq{\bm{1}}$ and $i\in[d]$
$\left|(\mathbf{s}_{\mathbf{k}-{\bm{1}}})_{i}-(\mathbf{s}_{\mathbf{k}})_{i}\right|=2^{-k_{i}}\left|t_{i}-s_{i}\right|=\left|(\mathbf{s}_{\mathbf{k}})_{i}-s_{i}\right|\,.$
Together with the assumption on $\mathbf{t}$, this allows us to apply Proposition 4.7 to $\oint_{\mathbf{s}_{\mathbf{k}}}^{\mathbf{s}_{\mathbf{k}-{\bm{1}}}}f(d\mathbf{y})$ to get
$\left\|\oint_{\mathbf{s}_{\mathbf{k}}}^{\mathbf{s}_{\mathbf{k}-{\bm{1}}}}f(d\mathbf{y})\right\|\lesssim\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}2^{-\mathbf{k}*({\bm{\alpha}}+{\bm{1}})}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}^{{\bm{\alpha}}+{\bm{1}}}\,.$
This is clearly summable in $\mathbf{k}$ and summing over $\mathbf{k}$ gives (4.15). (4.16) follows analogously. ∎
Remark 4.11.
With the same proof, we can also tackle the case in which $\mathbf{s}-\frac{1}{2}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}\in[0,T]$ and $\mathbf{s}\leq\mathbf{t}\leq T{\bm{1}}$. More generally, the same argument allows us to construct
$\int_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y})$
for any ${\bm{0}}\leq\mathbf{s}\leq\mathbf{t}\leq T{\bm{1}}$, such that for any $i\in[d]$, we have
$0\leq s_{i}-\frac{1}{2}\left|t_{i}-s_{i}\right|\qquad\text{or}\qquad t_{i}+\frac{1}{2}\left|t_{i}-s_{i}\right|\leq T\,.$
If we are given any $\mathbf{s}\leq\mathbf{t}\in[0,T]^{d}$, we can extend the primitive by separating the rectangle spanned between $\mathbf{s}$ and $\mathbf{t}$ into $2^{d}$ smaller rectangles at the middle point $u=\frac{1}{2}(\mathbf{s}+\mathbf{t})$. The additivity property (4.10) as well as (4.15) and (4.16) immediately extend to this case. Let us summarize this in a proposition.
Proposition 4.12.
Let $f\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ with $-{\bm{1}}<{\bm{\alpha}}<{\bm{0}}$. For any $\mathbf{s}\leq\mathbf{t}\in[0,T]^{d}$, let $\mathbf{u}=\frac{1}{2}(\mathbf{s}+\mathbf{t})$. Define
$\int_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y}):=\sum_{\theta\subset[d]}\int_{\pi^{\theta}_{\mathbf{s}}\mathbf{u}}^{\pi^{\theta}_{\mathbf{u}}\mathbf{t}}f(d\mathbf{y})\,.$
Then the above is well-defined, fulfills the additivity property (4.10), and is $\mathcal{F}_{\mathbf{t}}$-measurable. Furthermore, we have for all $\eta\subset[d]$
$\left\|\int_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y})\right\|_{m}$ $\lesssim\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}^{{\bm{\alpha}}+{\bm{1}}}$ (4.17)
$\left\|\mathbb{E}^{\eta}_{\mathbf{s}}\int_{\mathbf{s}}^{\mathbf{t}}f(d\mathbf{y})\right\|_{m}$ $\lesssim\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}^{{\bm{\alpha}}+{\bm{1}}}\left|\mathbf{t}-\mathbf{s}\right|_{\text{c},\eta}^{{\bm{\delta}}}\,.$ (4.18)
Similar to Remark 4.9, due to its additivity we can think of the primitive of $f$ as a one-parameter process $Y_{\mathbf{t}}=\int_{\bm{0}}^{\mathbf{t}}f(d\mathbf{y})$. (4.17) and (4.18) immediately give us the Hölder regularity of $Y$.
Corollary 4.13.
Due to the additivity of $\int f(d\mathbf{y})$, there is a unique process $Y$, which is 00 on the boundary $\bigcup_{i\in[d]}\pi^{\{i\}^{c}}_{\bm{0}}[0,T]^{d}$ such that
$\square^{[d]}_{\mathbf{x},\mathbf{y}}Y=\int_{\mathbf{x}}^{\mathbf{y}}f(d\mathbf{y})\,.$
This implies for any $\theta\subset[d]$, we have
$\square^{\theta}_{\mathbf{x},\mathbf{y}}Y=\int_{\pi_{\mathbf{x}}^{\theta}{\bm{0}}}^{\pi^{\theta}_{\mathbf{y}}\mathbf{x}}f(d\mathbf{y})\,.$
It follows that $Y\in C^{{\bm{\alpha}}+{\bm{1}},{\bm{\delta}}}L_{m}$. We call $Y$ the primitive, and we have
$\left\|Y\right\|_{C^{{\bm{\alpha}}+{\bm{1}},{\bm{\delta}}}L_{m}}\lesssim\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\,.$
4.4 Existence and uniqueness
We are now ready to state the main theorem for the existence and uniqueness of a mild solution (4.2) on any finite time interval $[0,T]^{d}$ and an $m\geq 2$.
Theorem 4.14.
Let $\xi$ be white noise over +$\mathbb{R}^{d}_{+}$ and $Z\in C^{{\bm{\beta}}}([0,T]^{d})$ for some $\frac{1}{2}{\bm{1}}<{\bm{\beta}}<{\bm{1}}$. Let also be ${\bm{\delta}}={\bm{\beta}}-\frac{1}{2}{\bm{1}}$ and ${\bm{\alpha}}$ satisfying ${\bm{\delta}}\leq{\bm{\alpha}}<\frac{1}{2}{\bm{1}}$ and ${\bm{\delta}}+{\bm{\beta}}+{\bm{\alpha}}>{\bm{1}}$. Under the assumption $\sigma\in\mathcal{C}^{1}(\mathbb{R})$ and $f\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{\infty}([0,T]^{d})$ for any initial datum $v\colon\partial_{{\bm{0}}}[0,T]^{d}\to\mathbb{R}$ such that $\mathcal{I}(v)\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,T]^{d})$ there exists a unique solution $u\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,T]^{d})$ to the equation (4.2)
Proof.
We will use the shorthand notation
$\zeta=\frac{\partial^{d}Z}{\partial x_{1}\ldots\partial x_{d}}\in C^{{\bm{\beta}}-{\bm{1}}}\,,\quad v_{0}=\mathcal{I}(v)\,.$
We start by showing local existence and uniqueness, that is we first show that for any initial condition $v_{0}\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ there exists a time horizon $\tau>0$ where the equation (4.2) admits a unique solution. The global result will then be constructed by patching together local solutions over the whole time $T>0$. For any $0<\tau\leq T$ we consider the affine space
$C_{\tau,v_{0}}=\{u\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,\tau]^{d})\colon\pi^{i}_{{\bm{0}}}u=\pi^{i}_{{\bm{0}}}v_{0}\quad\text{for any}\;i\in[d]\}\,.$
Similarly, we define the solution map
$\mathcal{S}_{\tau}(u)=\left\{\mathbf{x}\mapsto v_{0}(\mathbf{x})+\int_{\bm{0}}^{\mathbf{x}}\mathcal{R}(\sigma(u)\cdot\xi+(fu)\cdot\zeta)(d\mathbf{y})\colon\mathbf{x}\in[0,\tau]^{d}\right\}$
We claim that $\mathcal{S}_{\tau}\colon C_{\tau,v_{0}}\to C_{\tau,v_{0}}$ is a well-posed application which becomes a contraction for $\tau>0$ small enough, by endowing $C_{\tau,v_{0}}$ with the distance $d(u,v)=\left\|u-v\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}([0,\tau]^{d})}$. Let us first check that $\mathcal{S}_{\tau}(u)\in C_{\tau,v_{0}}$. Our choice of parameters ensures ${\bm{\beta}}-{\bm{1}}+{\bm{\alpha}}>-\frac{1}{2}{\bm{1}}$ as well as ${\bm{\beta}}-{\bm{1}}+{\bm{\alpha}}+{\bm{\delta}}>{\bm{0}}$. Thus, by Proposition 4.2 together with Proposition 4.3, we have
$\mathcal{R}((fu)\cdot\zeta))\in C^{{\bm{\beta}}-{\bm{1}}}L_{m}$
Similarly, from Proposition 4.1 and Corollary 4.6
$\mathcal{R}(\sigma(u)\cdot\xi)\in C^{-\frac{1}{2}{\bm{1}},+\infty}L_{m}$
Summing the two distributions, we have from Lemma 3.6 that
$\mathcal{R}(\sigma(u)\cdot\xi+(fu)\cdot\zeta)\in C^{-\frac{1}{2}{\bm{1}},{\bm{\delta}}}L_{m}\,,$
where we used ${\bm{\delta}}={\bm{\beta}}-\frac{1}{2}{\bm{1}}$. After the integration and the trivial inclusion $C^{\frac{1}{2}{\bm{1}},{\bm{\delta}}}L_{m}\subset C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ we can conclude that $\mathcal{S}_{\tau}(u)\in C_{\tau,v_{0}}$. Passing to the contraction property, we fix $u,v\in C_{\tau,v_{0}}$. Using Corollary 4.6 we have the estimate
$\left\|\mathcal{R}(\sigma(u)\cdot\xi)-\mathcal{R}(\sigma(v)\cdot\xi)\right\|_{C^{-\frac{1}{2}{\bm{1}},\infty}L_{m}}$ $\lesssim\left\|\sigma\right\|_{\mathcal{C}^{1}}\left\|u-v\right\|_{C^{{\bm{\alpha}}}L_{m}}$ (4.19)
$\lesssim\left\|\sigma\right\|_{\mathcal{C}^{1}}\left\|u-v\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\,,$
where we used the Hölder embedding given in Proposition 3.3 in the second line and we simply bound the elements with $\tau$ by $T$. Let us now analyze the term $\mathcal{R}((fu)\cdot\zeta-(fv)\cdot\zeta)$. Propositions 4.3 and 4.2 come also with the estimate
$\begin{split}&\left\|\mathcal{R}((fu)\cdot\zeta)-\mathcal{R}((fv)\cdot\zeta)\right\|_{C^{{\bm{\beta}}-{\bm{1}}}L_{m}}=\left\|\mathcal{R}((f(u-v))\cdot\zeta)\right\|_{C^{{\bm{\beta}}-{\bm{1}}}L_{m}}\\ &\lesssim\left\|f(u-v)\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\left\|\zeta\right\|_{C^{\bm{\beta}}}\lesssim\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{\infty}}\left\|u-v\right\|_{C^{\alpha,{\bm{\delta}}}L_{m}}\left\|\zeta\right\|_{C^{\bm{\beta}}}\,.\end{split}$ (4.20)
By Lemma 3.6, we have
$\left\|\mathcal{R}(\sigma(u)\cdot\xi)+\mathcal{R}((fu)\cdot\zeta)-\mathcal{R}(\sigma(v)\cdot\xi)-\mathcal{R}((fv)\cdot\zeta)\right\|_{C^{-\frac{1}{2}{\bm{1}},{\bm{\delta}}}L_{m}}$
$\qquad\qquad\qquad\qquad\lesssim\left(\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{\infty}}\left\|\zeta\right\|_{C^{\bm{\beta}}}+\left\|\sigma\right\|_{\mathcal{C}^{1}}\right)\left\|u-v\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\,.$
Using the Corollary (4.13), it follows that
$\left\|\int_{\bm{0}}^{\cdot}\left(\mathcal{R}(\sigma(u)\cdot\xi)+\mathcal{R}((fu)\cdot\zeta)-\mathcal{R}(\sigma(v)\cdot\xi)-\mathcal{R}((fv)\cdot\zeta)\right)(d\mathbf{y})\right\|_{C^{\frac{1}{2}{\bm{1}},{\bm{\delta}}}L_{m}}$
$\qquad\qquad\qquad\qquad\lesssim\left(\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{\infty}}\left\|\zeta\right\|_{C^{\bm{\beta}}}+\left\|\sigma\right\|_{\mathcal{C}^{1}}\right)\left\|u-v\right\|_{C^{{\bm{\alpha}},{\bm{\beta}}}L_{m}}$
Finally, we use the Hölder embedding result shown in Proposition 3.2 to compute
$\left\|\mathcal{S}_{\tau}(u)-\mathcal{S}_{\tau}(v)\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\lesssim\sum_{\theta\subset[d]}(\tau{\bm{1}})^{\frac{1}{2}{\bm{1}}-{\bm{\alpha}}_{\theta}}\left\|\mathcal{S}_{\tau}(u)-\mathcal{S}_{\tau}(v)\right\|_{C^{\frac{1}{2}{\bm{1}},{\bm{\delta}}}L_{m}}$
$\sum_{\theta\subset[d]}(\tau{\bm{1}})^{\frac{1}{2}{\bm{1}}-{\bm{\alpha}}_{\theta}}\left\|\int_{\bm{0}}^{\cdot}\left(\mathcal{R}(\sigma(u)\cdot\xi)+\mathcal{R}((fu)\cdot\zeta)-\mathcal{R}(\sigma(v)\cdot\xi)-\mathcal{R}((fv)\cdot\zeta)\right)(d\mathbf{y})\right\|_{C^{\frac{1}{2}{\bm{1}},{\bm{\delta}}}L_{m}}$
$\lesssim\sum_{\theta\subset[d]}(\tau{\bm{1}})^{\frac{1}{2}{\bm{1}}-{\bm{\alpha}}_{\theta}}\left(\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{\infty}}\left\|\zeta\right\|_{C^{\bm{\beta}}}+\left\|\sigma\right\|_{\mathcal{C}^{1}}\right)\left\|u-v\right\|_{C^{{\bm{\alpha}},{\bm{\beta}}}L_{m}}\,,$
which becomes a contraction for some parameter $\tau>0$ small enough, which we denote by $T^{*}$. By the uniqueness of Banach’s fixed point theorem, we immediately have the uniqueness of the local (and thus global) solution. To see the global existence, we will explicitly show how to construct the solution on $[0,T]^{2}$, for the sake of clarity leaving the general construction in general dimension $d$ as a trivial extension. Note that the value $T^{*}$ constructed above does not depend on $v_{0}$, but only on $\left\|f\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{\infty}}$, $\left\|\zeta\right\|_{C^{\bm{\beta}}}$, $\left\|\sigma\right\|_{\mathcal{C}^{1}}$. Thus, by choosing $N$ large enough, such that $T/N\leq T^{*}$, the above proof shows the existence and uniqueness of the solution over every small patch
$A^{\mathbf{k}}=A^{(k_{1},k_{2})}=[k_{1}T/N,(k_{1}+1)T/N]\times[k_{2}T/N,(k_{2}+1)T/N]\,,$
where $k_{1},k_{2}=0,\dots,N-1$ and the initial condition for each patch is given by the equation solved on the previous patches in lexicographic order. We call the solutions on those patches $u^{\mathbf{k}}$ and set for any $\mathbf{x}\in A^{\mathbf{k}}$ $u(\mathbf{x})=u^{\mathbf{k}}(\mathbf{x})$ This gives us a unique solution to the equation (4.2), as long as $u\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$. To see this, we use the following notation: For any $x_{i}\in[0,T],i=1,2$ and $k\in[N-1]$, set
$x_{i}^{(k)}:=\begin{cases}x_{i}\qquad&x_{i}\in[kT/N,(k+1)T/N]\\ kT/N\qquad&x_{i}<kT/N\\ (k+1)T/N\qquad&x_{i}>(k+1)T/N\,,\end{cases}$
and for any patch $A^{\mathbf{k}}$, $\mathbf{x}\in[0,T]^{2}$ we set $\mathbf{x}^{(\mathbf{k})}:=(x_{1}^{(k_{1})},x_{2}^{(k_{2})})$. Then, the additivity of the rectangular increment immediately gives us for any $\mathbf{x}\leq\mathbf{y}$ in $[0,T]^{2}$ and $\eta\subset\theta\subset[d]$:
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square^{\theta}_{\mathbf{x},\mathbf{y}}u\right\|_{m}\leq\sum_{\mathbf{k}\in[N-1]^{2}}\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square_{\mathbf{x}^{(k)},\mathbf{y}^{(k)}}^{\theta}u\right\|_{m}\,.$
Note that $\square^{\theta}_{\mathbf{x}^{(k)},\mathbf{y}^{(k)}}u=0$ for all $\mathbf{k}$ such that $x_{i}\leq y_{i}\leq k_{i}T/N$ for some $i\in\theta$, so we get no problems with measurability. Furthermore, we have $\left|x^{(k)}-y^{(k)}\right|\leq\left|x-y\right|$ for all $x,y\in[0,T]$ and $k\in[N-1]$. One easily calculates now that
$\left\|u\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}\leq\sum_{\mathbf{k}\in[N-1]^{2}}\left\|u^{\mathbf{k}}\right\|_{C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}}<\infty.$
5 Connection to stochastic multiparameter sewing
In dimension $d=1$, sewing can be seen as a special case of the reconstruction theorem [BZ22]. This observation also extends to the stochastic case [Ker23], but in general no longer holds in dimension $d>1$. Indeed, Harang’s multiparameter sewing lemma [Har21] and the reconstruction theorem described in [CZ20] are not connected in similarly. We claim that our formulation using rectangular increment is the missing link between the two results. More precisely, this section shows that Theorem 3.12 extends the stochastic multiparameter sewing lemma from [BHK23]. Let us start by recalling the sewing lemma in question. Throughout this section, we fix a commuting filtration $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in[0,T]^{d}}$.
A germ in the world of sewing is a $2d$-parameter process given by a map $\Xi:\Delta_{T}\times\Omega\to\mathbb{R}$, where
$\Delta_{T}=\{(\mathbf{s},\mathbf{t})\colon{\bm{0}}\leq\mathbf{s}\leq\mathbf{t}\leq T{\bm{1}}\}$
is the simplex over $[0,T]^{d}$. We say that $\Xi$ is $\mathcal{F}$-adapted, if for all $(\mathbf{s},\mathbf{t})\in\Delta_{T}$, $\Xi_{\mathbf{s},\mathbf{t}}$ is $\mathcal{F}_{\mathbf{t}}$ measurable. One aims to construct a Riemann-type sum
$\mathcal{I}\Xi_{\mathbf{s},\mathbf{t}}:=\lim_{\left|\mathcal{P}\right|\to 0}\sum_{[\mathbf{u},\mathbf{v}]\in\mathcal{P}}\Xi_{\mathbf{u},\mathbf{v}}\,,$ (5.1)
where the limit goes over grid-like partitions with vanishing mesh size. A grid-like partition of the interval $[s_{1},t_{1}]\times\dots\times[s_{d},t_{d}]$ is a set of the form
$\mathcal{P}=\mathcal{P}_{1}\times\dots\times\mathcal{P}_{d}\,$
where $\mathcal{P}_{i}$ is a partition of $[s_{i},t_{i}],i\in[d]$ and its mesh size is the supremum
$\left|\mathcal{P}\right|=\sup_{[u_{i},v_{i}]\in\mathcal{P}_{i},i\in[d]}\left|u_{i}-v_{i}\right|\,.$
For each indexset $\theta\subset[d]$, we define
$\mathcal{P}^{\theta}=\mathcal{Q}_{1}\times\dots\mathcal{Q}_{d}\,$
$\mathcal{Q}_{i}=\mathcal{P}_{i}$ for $i\in\theta$ and $\mathcal{Q}_{i}=\{[s_{i},t_{i}]\}$ for $i\notin\theta$. Similar to one-dimensional sewing, the multiparameter sewing lemma is based on the $\delta$-operator, which in $d=1$ is given by
$\delta_{u}\Xi_{s,t}=\Xi_{s,t}-(\Xi_{s,u}+\Xi_{u,t})\,,$
for $s\leq u\leq t$. It can be generalized to a multiparameter setting by applying a $\delta$ operator in each direction. More precisely, in $d\geq 1$ we set for all $i\in[d]$ and $\mathbf{s}\leq\mathbf{u}\leq\mathbf{t}$
$\delta^{i}_{\mathbf{u}}\Xi_{\mathbf{s},\mathbf{t}}=\Xi_{\mathbf{s},\mathbf{t}}-(\Xi_{\mathbf{s},\pi^{i}_{\mathbf{u}}\mathbf{t}}+\Xi_{\pi^{i}_{\mathbf{u}}\mathbf{s},\mathbf{t}})\,.$
For indexsets $\eta\subset[d]$, this can be extended as usual by
$\delta_{\mathbf{u}}^{\eta}:=\prod_{i\in\eta}\delta^{i}_{\mathbf{u}}\,.$
With these definitions, the coherence property for $2d$-parameter objects reads as follows.
Definition 5.1.
Let $\Xi:\Delta_{T}\times\Omega\to\mathbb{R}$ be an $\mathcal{F}$-adapted $2d$-parameter process. Further let ${\bm{\alpha}},{\bm{\beta}},{\bm{\gamma}}\geq{\bm{0}}$ and $1\leq m<\infty$. Then we call $\Xi$ stochastic $({\bm{\alpha}},{\bm{\beta}},{\bm{\gamma}})$-coherent in the sewing sense if there is a constant $C>0$ such that for all $(\mathbf{s},\mathbf{t})\in\Delta_{T}$
$\left\|\Xi_{\mathbf{s},\mathbf{t}}\right\|_{m}\leq C\left|\mathbf{t}-\mathbf{s}\right|_{\text{c}}^{{\bm{\alpha}}}\,,$ (5.2)
holds and for all $\eta\subset\theta\subset[d]$ with $\theta\neq\emptyset$ and $\mathbf{s}\leq\mathbf{u}\leq\mathbf{t}$,
$\left\|\mathbb{E}^{\eta}_{\mathbf{s}}\delta^{\theta}_{\mathbf{u}}\Xi_{\mathbf{s},\mathbf{t}}\right\|_{m}\leq C\left|\mathbf{t}-\mathbf{s}\right|^{{\bm{\alpha}}}_{\text{c},\theta^{c}}\left|\mathbf{t}-\mathbf{s}\right|^{{\bm{\beta}}}_{\text{c},\theta}\left|\mathbf{t}-\mathbf{s}\right|^{{\bm{\gamma}}}_{\text{c},\eta}\,.$ (5.3)
We write $\Xi\in C_{2}^{{\bm{\alpha}},{\bm{\beta}},{\bm{\gamma}}}L_{m}$ and denote by $\left\|\Xi\right\|_{C_{2}^{{\bm{\alpha}},{\bm{\beta}},{\bm{\gamma}}}L_{m}}$ the smallest possible constant $C\geq 0$ such that (5.3), (5.2) hold.
This coherence property suffices to show the following.
Theorem 5.2 (Multiparameter stochastic sewing, [BHK23]).
Let $\Xi\in C_{2}^{{\bm{\alpha}},{\bm{\beta}},{\bm{\gamma}}}L_{m}$. Assume $2\leq m<\infty,{\bm{\alpha}}\in[0,1]^{d}$, ${\bm{\gamma}}\in[0,\frac{1}{2})^{d},{\bm{\beta}}\in(\frac{1}{2},\infty)^{d}$ with ${\bm{\beta}}+{\bm{\gamma}}>{\bm{1}}$. Then, for $\theta\subset[d],(\mathbf{s},\mathbf{t})\in\Delta_{T}$ the following limit exists:
$\mathcal{I}^{\theta}\Xi_{\mathbf{s},\mathbf{t}}=L_{m}-\lim_{\left|\mathcal{P}\right|\to 0}\sum_{[\mathbf{u},\mathbf{v}]\in\mathcal{P}^{\theta}}\Xi_{\mathbf{u},\mathbf{v}}\,.$
Furthermore, the limits $(\mathcal{I}^{\theta}\Xi_{\mathbf{s},\mathbf{t}})_{\theta\subset[d],(\mathbf{s},\mathbf{t})\in\Delta_{T}}$ are the unique family of $2d$-parameter processes such that $\mathcal{I}^{\theta}\Xi_{\mathbf{s},\mathbf{t}}\in L_{m}(\Omega)$ holds for all $\theta\subset[d]$ and $(\mathbf{s},\mathbf{t})\in\Delta_{T}$ and the following four properties hold.
  • i)
    $\mathcal{I}^{\emptyset}\Xi=\Xi$.
  • ii)
    $\mathcal{I}^{\theta}\Xi_{\mathbf{s},\mathbf{t}}$ is $\mathcal{F}_{\mathbf{t}}$-measurable for all $\theta\subset[d],\mathbf{s}\leq\mathbf{t}$.
  • iii)
    For all $\theta\subset[d]$, $\mathcal{I}^{\theta}\Xi$ is additive in $\mathbb{R}^{\theta}$. I.e. for all $i\in\theta$, we have that $\delta^{i}_{\mathbf{u}}\mathcal{I}^{\theta}\Xi_{\mathbf{s},\mathbf{t}}=0$ for all $\mathbf{s}\leq\mathbf{u}\leq\mathbf{t}$.
  • iv)
    For all $\eta\subset\theta\subset[d]$ with $\theta\neq\emptyset$ and $\mathbf{s}\leq\mathbf{t}$, we have:
    $\left\|\mathbb{E}^{\eta}_{\mathbf{s}}\sum_{\hat{\theta}\subset\theta}(-1)^{\left|\hat{\theta}\right|}\mathcal{I}^{\hat{\theta}}\Xi_{\mathbf{s},\mathbf{t}}\right\|_{m}\lesssim\left\|\Xi\right\|_{C_{2}^{{\bm{\alpha}},{\bm{\beta}},{\bm{\gamma}}}L_{m}}\left|\mathbf{t}-\mathbf{s}\right|^{{\bm{\alpha}}}_{\text{c},\theta^{c}}\left|\mathbf{t}-\mathbf{s}\right|^{{\bm{\beta}}}_{\text{c},\theta}\left|\mathbf{t}-\mathbf{s}\right|^{{\bm{\gamma}}}_{\text{c},\eta}\,.$
Remark 5.3.
Comparing the coherence properties for sewing and reconstruction, we see a slight difference in the choice of $\mathbf{x},\mathbf{y}\in[0,T]^{d}$ compared to $(\mathbf{s},\mathbf{t})\in\Delta_{T}$: Reconstruction allows all $\mathbf{x},\mathbf{y}$ such that $\mathbf{x}_{\theta}\leq\mathbf{y}_{\theta}$ holds in equation (3.8), whereas for sewing, (5.3) only holds for $\mathbf{s}\leq\mathbf{t}$. The property $\mathbf{x}_{\theta}\leq\mathbf{y}_{\theta}$ is necessary to construct the reconstructions $f_{\mathbf{x}}^{\theta}(\psi)$ for test functions with support in $(0,T)^{d}$. If we restrict (3.8) to only hold for $\mathbf{x}\leq\mathbf{y}$, the limit $f_{\mathbf{x}}^{\theta}(\psi)$ only exists for test functions $\psi$ with support in $(x_{1},T)\times\dots(x_{d},T)$. Apart from this, Theorem 3.12 still holds in this setting. This restriction on $\mbox{supp}(\psi)$ mirrors the fact that $\mathcal{I}^{\theta}\Xi_{\mathbf{s},\mathbf{t}}$ only exists for $\mathbf{s}\leq\mathbf{t}$.
We claim that under the assumption of Theorem 5.2, the distributional derivative of $\Xi$ given by
$\mathcal{F}_{\mathbf{x}}(\psi):=(-1)^{d}\int_{[0,T]^{d}}\Xi_{\mathbf{x},\mathbf{z}}\partial^{\bm{1}}\psi(\mathbf{z})d\mathbf{z}=\left\langle\partial^{\bm{1}}\Xi_{\mathbf{x},\cdot},\psi\right\rangle$
is stochastic $(-{\bm{1}},{\bm{\beta}}-{\bm{1}},{\bm{\gamma}})$-coherent and can thus be reconstructed (up to the restriction described in Remark 5.3). Here, $\partial^{\bm{1}}\psi:=\frac{\partial^{d}}{\partial x_{1}\dots\partial x_{d}}\psi$. Furthermore, comparing the four properties uniquely characterizing $\mathcal{I}^{\theta}\Xi$ and $f^{\theta}$, one quickly sees that $f^{\theta}$ is given by the distributional derivative of $\mathcal{I}^{\theta}\Xi$. Let us formalize this observation.
Proposition 5.4.
Let $\Xi\in C_{2}^{{\bm{\alpha}},{\bm{\beta}},{\bm{\gamma}}}L_{m}$ with ${\bm{\alpha}},{\bm{\beta}},{\bm{\gamma}},m$ as in Theorem 5.2 and we set $r=2$. Further assume that for all $\mathbf{s}\in[0,T]^{d}$, $\mathbf{t}\mapsto\Xi_{\mathbf{s},\mathbf{t}}$ is almost surely in $L^{1}([s_{1},T]\times\dots\times[s_{d},T])$. Then the germ $F_{\mathbf{x}}:=\partial^{\bm{1}}\Xi_{\mathbf{x},\cdot}$ is extended stochastic $(-{\bm{1}},{\bm{\beta}}-{\bm{1}},{\bm{\gamma}})$-coherent. It follows that it can be reconstructed, and the reconstruction $f^{\theta}$ is connected to the sewings $\mathcal{I}^{\theta}\Xi$ via
$f^{\theta}_{\mathbf{x}}(\psi)=(-1)^{d}\int_{[0,T]^{d}}\mathcal{I}^{\theta}\Xi_{\mathbf{x},\mathbf{z}}\partial^{\bm{1}}\psi(\mathbf{z})d\mathbf{z}=\left\langle\partial^{\bm{1}}\mathcal{I}^{\theta}_{\mathbf{x},\cdot},\psi\right\rangle\,.$
for all $\theta\subset[d],\mathbf{x}\in[0,T]^{d}$ and $\psi\in C^{rd}_{c}((x_{1},T)\times\dots\times(x_{d},T))$.
As mentioned in Remark 5.3, $f^{\theta}(\psi)$ only exists for $\psi\in C^{rd}_{c}((x_{1},T)\times\dots\times(x_{d},T))$.
Proof.
We define the projections $\pi^{i,1},\pi^{i,2}$ by
$\pi^{i,1}_{\mathbf{u}}(\mathbf{s},\mathbf{t})=(\pi_{\mathbf{u}}^{i}\mathbf{s},\mathbf{t})\qquad\pi^{i,2}_{\mathbf{u}}(\mathbf{s},\mathbf{t})=(\mathbf{s},\pi^{i}_{\mathbf{u}}\mathbf{t})\,,$
for all $i\in[d]$. Since we only consider test functions with compact support, Fubini gives us for all $\psi\in C_{c}^{rd}([0,T]^{d})$:
$\int_{[0,T]^{d}}\pi^{i,2}_{\mathbf{y}}\Xi_{\mathbf{x},\mathbf{z}}\partial^{\bm{1}}\psi(\mathbf{z})d\mathbf{z}=0$
for all $i\in[d]$ and $\mathbf{x}\leq\mathbf{y}$. Thus, the following holds for $\mathbf{x}\leq\mathbf{y}$ and $\psi\in C_{c}^{rd}((y_{1},T)\times\dots\times(y_{d},T))$.
$\square^{\theta}_{\mathbf{x},\mathbf{y}}F(\psi)$ $=(-1)^{d}\prod_{i\in\theta}\left\langle(\pi_{\mathbf{y}}^{i,1}-\mbox{Id})\Xi_{\mathbf{x},\cdot},\partial^{\bm{1}}\psi\right\rangle$
$=(-1)^{d}\prod_{i\in\theta}\left\langle(\pi_{\mathbf{y}}^{i,1}+\pi_{\mathbf{y}}^{i,2}-\mbox{Id})\Xi_{\mathbf{x},\cdot},\partial^{\bm{1}}\psi\right\rangle$
$=(-1)^{d+\sharp{\theta}}\left\langle\delta^{\theta}_{\mathbf{y}}\Xi_{\mathbf{x},\cdot},\partial^{\bm{1}}\psi\right\rangle\,,$
Let us conclude that $F$ is stochastic coherent. For $\mathbf{x},\mathbf{y},\psi$ and ${\bm{\lambda}}$ as in Definition 3.8 and $\eta\subset\theta\subset[d]$, we have
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square^{\theta}_{\mathbf{x},\mathbf{y}}F(\psi_{\mathbf{y}}^{\bm{\lambda}})\right\|_{m}$ $\leq\sup_{\mathbf{z}\in\mbox{supp}(\psi_{\mathbf{y}}^{\bm{\lambda}})}\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\delta^{\theta}_{\mathbf{y}}\Xi_{\mathbf{x},\mathbf{z}}\right\|_{m}\left\|\partial^{\bm{1}}\psi_{\mathbf{y}}^{\bm{\lambda}}\right\|_{L_{1}(\mathbb{R}^{d})}$
$\leq\left\|\Xi\right\|_{C_{2}^{{\bm{\alpha}},{\bm{\beta}},{\bm{\gamma}}}L_{m}}{\bm{\lambda}}^{-{\bm{1}}}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\beta}}_{\theta}+{\bm{\gamma}}_{\eta}+{\bm{\alpha}}_{\theta^{c}}}$ (5.4)
$\lesssim\left\|\Xi\right\|_{C_{2}^{{\bm{\alpha}},{\bm{\beta}},{\bm{\gamma}}}L_{m}}{\bm{\lambda}}^{-{\bm{1}}}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{{\bm{\beta}}_{\theta}+{\bm{\gamma}}_{\eta}}\,.$
By assumption, ${\bm{\beta}}-{\bm{1}}>-\frac{1}{2}{\bm{1}}$ and ${\bm{\beta}}+{\bm{\gamma}}-{\bm{1}}>0$. Thus, we can apply Theorem 3.12 to get reconstructions $f^{\theta}$, $\theta\subset[d]$. We claim that $f^{\theta}_{\mathbf{x}}=\partial^{\bm{1}}\mathcal{I}^{\theta}\Xi_{\mathbf{x},\cdot}$ for all $\theta\subset[d],\mathbf{x}\in[0,T]^{d}$. To prove this, we check the four properties uniquely characterizing $f^{\theta}$ in Theorem 3.12:
  • $\partial^{\bm{1}}\mathcal{I}^{\emptyset}\Xi_{\mathbf{x},\cdot}=\partial^{\bm{1}}\Xi_{\mathbf{x},\cdot}=F_{\mathbf{x}}$.
  • $\left\langle\partial^{\bm{1}}\mathcal{I}^{\theta}\Xi_{\mathbf{x},\cdot},\psi\right\rangle$ is $\mathcal{F}_{\mathbf{x}\vee\mathbf{y}}$-measurable for all $\psi$ with $\mbox{supp}(\psi)\subset(-\infty,\mathbf{y}]$.
  • As before, $\left\langle\partial^{\bm{1}}\mathcal{I}^{\theta}\Xi_{\mathbf{x},\pi^{i}_{\tilde{\mathbf{x}}}\cdot},\psi\right\rangle=0$ for all $\mathbf{x}\leq\tilde{\mathbf{x}}$ and $i\in[d]$ due to $\psi$ having a compact support. Together with the additivity of $\mathcal{I}^{\theta}\Xi$, this implies that
    $\left\langle\partial^{\bm{1}}\mathcal{I}^{\theta}\Xi_{\mathbf{x},\cdot},\psi\right\rangle=\left\langle\partial^{\bm{1}}\mathcal{I}^{\theta}\Xi_{\pi^{i}_{\tilde{\mathbf{x}}}\mathbf{x},\cdot}+\partial^{\bm{1}}\mathcal{I}^{\theta}\Xi_{\mathbf{x},\pi^{i}_{\tilde{\mathbf{x}}}\cdot},\psi\right\rangle=\left\langle\partial^{\bm{1}}\mathcal{I}^{\theta}\Xi_{\pi^{i}_{\tilde{\mathbf{x}}}\mathbf{x},\cdot},\psi\right\rangle\,,$
    for all $i\in\theta$. It follows that $\partial^{\bm{1}}\mathcal{I}^{\theta}\Xi_{\mathbf{x},\cdot}$ only depends on $x_{i}$ for $i\notin\theta$.
  • Let $\psi,\mathbf{x}\leq\mathbf{z},{\bm{\lambda}},\eta,\theta$ be as in (3.10) of Theorem 3.12. Then:
    $\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\sum_{\hat{\theta}\subset\theta}(-1)^{\sharp{\hat{\theta}}}\left\langle\partial^{\bm{1}}\mathcal{I}^{\hat{\theta}}\Xi_{\mathbf{x},\cdot},\psi_{\mathbf{z}}^{\bm{\lambda}}\right\rangle\right\|_{m}$ $\leq\sup_{\mathbf{y}\in\mbox{supp}(\psi_{\mathbf{z}}^{\bm{\lambda}})}\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\sum_{\hat{\theta}\subset\theta}(-1)^{\sharp{\hat{\theta}}}\mathcal{I}^{\hat{\theta}}\Xi_{\mathbf{x},\mathbf{y}}\right\|_{m}\left\|\partial^{\bm{1}}\psi_{\mathbf{z}}^{\bm{\lambda}}\right\|_{L^{1}(\mathbb{R}^{d})}$
    $\lesssim{\bm{\lambda}}^{-{\bm{1}}}{\bm{\lambda}}^{{\bm{\beta}}}_{\theta}{\bm{\lambda}}^{{\bm{\gamma}}}_{\eta}\,,$
    where we used $\left|y_{i}-x_{i}\right|=\left|y_{i}-z_{i}\right|\lesssim\lambda_{i}$ for all $\mathbf{y}\in\mbox{supp}(\psi_{\mathbf{z}}^{\bm{\lambda}})$ and $i\in\theta$.
This shows that $\partial^{\bm{1}}\mathcal{I}^{\theta}\Xi_{\mathbf{x},\cdot}=f_{\mathbf{x}}^{\theta}$, which finishes the proof. ∎
Remark 5.5.
Surprisingly, the parameter ${\bm{\alpha}}$ from the sewing lemma does not turn up at all on the reconstruction side. It is rather replaced by the regularity of the distributional derivative $-{\bm{1}}$. If one wants to keep track of the Hölder regularity ${\bm{\alpha}}$, one would need to exchange (3.8) in the coherence property with a term of the form
$\left\|\mathbb{E}^{\eta}_{\mathbf{x}}\square_{\mathbf{x},\mathbf{y}}^{\theta}F(\psi_{\mathbf{y}}^{\bm{\lambda}})\right\|_{p}\lesssim{\bm{\lambda}}^{{\bm{\alpha}}}(\left|\mathbf{x}-\mathbf{y}\right|_{\text{c}}+{\bm{\lambda}})^{({\bm{\gamma}}-{\bm{\alpha}})_{\theta}+{\bm{\delta}}_{\eta}+\tilde{\bm{\alpha}}_{\theta^{c}}}\,,$
where ${\bm{\alpha}}$ would be the typical parameter encountered in reconstruction and $\tilde{\bm{\alpha}}$ is the Hölder continuity inherited from sewing.